A Complete Bibliography of Publications in Biometrika for the decade 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.03

Title word cross-reference

#4315 [Har79].

0 · 1 [dVW66]. 2 × 2 [BH60a, BH61a, OI61a, OI61b, Pla64]. 50 [Bar66d]. _{k} [Bur60b]. A [Hol66, Mar63]. A + B [Mar63]. b_{2} [Pea65b]. β [Mos62]. β_{2} [JNAP63, JNAP65]. χ [JP69, JP70]. χ^{2} [Hit62, MA65, Put64, Tik65b, Wis63a, Wis63b, You62]. χ_{r}^{2} [Sen67]. d [dVW66]. E_{k}/M/1 [Bur60c]. ϵ [RK69a]. exp(−a) + ka = 1 [BDM60]. exp(b) − b/(1 − p) = 1 [BDM63]. F [Ati62, BDO60, CB63, CB66, GS62b, PB61, Pri64a, SZ60, Tik65b, Tik66]. GI/G/1 [Kin62]. GI/M/1 [Fin60a]. k [DG68a, Kor69, Maa66]. M [Har69a, Pea69, Cro61, Lin60]. M/M/1 [GS65]. X^{2} [Wis63a, Wis63b]. N [Gil65, Hai65b, ST66, Arc62, Arc64, Hof63]. p [Rao68b, Rao69]. Q [Gow66a]. R^{3} [Mil65a]. ρ [Sno63]. S [Bur60a, Bar66d]. s = 1 [Bar66d]. s_{max}^{2}/s_{0}^{2} [Cha67a]. S_{U} [Joh65b]. σ_{B}^{2}/σ^{2} [Sis68]. \sqrt{b_{1}} [Pea65b]. \sqrt{β_{1}} [JNAP63].

∑_{0}^{m} f(Y_{t}) [Con65]. √β_{1} [JNAP65]. t [Amo64, FKM67, GJ68, Haj61, HPW61, Joh61, MSA66, Owe65, SA62a, Sis64]. T_{0}^{2} [IS64]. U^{2} [Maa66]. U_{M,N}^{2} [Ste65b]. U_{N}^{2} [PS62b, Ste63a, Ste64, Tik65a]. V_{1}^{(s)} [Bar66d]. V_{N} [Ste65a]. W_{N}^{2} [PS62b, Tik65a, SM68]. X^{2} [Wis64].

-Dependent [Rao69, Rao68b]. -Distribution [GJ68, HPW61, Joh61, Mos62, Owe65]. -Distributions [Pri64a]. -Ratio [PB61]. -Ratios [CB63]. -sample [Maa66]. -Technique [Gow66a]. -Test [Sen67, Ati62, CB66, Haj61, Sis64, You62]. -tests [MA65, SA62a].

/M/1 [Bur60b].

1 [Fin61].

2 [KLTW65].

30 [Har79].

51 [Har79]. 54 [MT72]. 55 [DS76]. 58 [Ano64d].

Aaboe [Ano64e]. Absolute [Nab61]. Absorbing [BW65a]. Abstracts [Bar60c, Yat62]. Abt [Ano63h]. Abuse [Bar63g]. Acceptance [Cam68]. Accident [Bha67, Hai65b]. Accidents [Hai65b]. Accuracy [Pea69]. AChE [BLN62]. Acheson [Wet60]. Acknowledgement [Kam67]. Action [AS64, AS66]. Acton [Buc60]. Adams [Dav62c]. Adding [Gow68]. Additional [Amo63, Bra64, Bra54]. Adjudging [Gea66]. Admitting [Rao60]. Advanced [Cox67b, Dav62f]. after [GM62, Saw60]. Against [Dav64h]. Age [Bat63, Goo67, Wau68]. Age-Dependent [Goo67, Wau68]. Agriculture [Dav64a, Hea62]. Aitken [Gon61a, Dav64h, Gon61b]. Alanen [Yat62]. Albert [Dav61i]. Alder [Dav62b]. Algebra [Dav60j, Dav61c]. Aliasing [HG61]. Alison [Ano64d, Pea66b]. Allgemeine [Joh60b, Joh62a]. Allgemeinen [Joh60c]. Allied [Gob67, Pea66c]. Allocation [Sti69, Zin62]. Alternative [CC67, KK66, Pea66a]. Alternatives [Cro68, Cro61, You62]. Amendments [DS76, Har79, LMH71, Tar90, Wis90, MT72]. Among [Dow66a, Mos63, Mos62]. Amount [Swa62, Swa63]. Amtlichen [Bar63f]. analogous [Hai61]. Analogs [Hal62]. Analogue [Kud63, San64, Maa66]. Analogues [Bre63]. Analysing [HG68, Pol66b, Bar62c]. Analysis [Ano63h, ASB60, Ati64, Bar61a, Bar64b, Bar67b, BDF63, BKM67, Bra54, Bra64, Bur62, Dur69, Ewe61, Gab68, Gar64, Gea66, Goo64, Gow66b, Gow68, HR67, HL65, HSD66, Ito62, Ito63, IS64, Joh65a, Ken61c, Ksh62, Ksh66, Lar69, McG67, Pha65, Pil64, Pil65b, Pil65a, Pil67, PB64, PR64, Pri64b, Rad64, Rad66, Rad67, Rao65a, Ree69, Rei61, Rei62, Sch53, Sch69b, Stu64, Tan62, TZ64, TT65, TT66, Tia66, TB67, TD68, Wal60, Wal67, Wal70, WL64, Wat66, Wet61, Whi64b, WG68, Bar65b, DG66, Mik65, MA65, SW65, Ano63f, Bar62d, Buc61, Cha68a, Dav60a, Saw61e, Buc60]. Analytical [Sed66]. Andrew [Yat62]. Angle [Hol67]. Angle-Count [Hol67]. Animal [Man69b, Wat66]. Animals [Cor64b]. Announcement [Ano66a]. ANOVA [Cox92]. Anthropology [Ano64n]. Any [Ver61]. Apart [Ver61]. Apelt [Bro62]. Application [Gar63, Gar64, GZ67, Les62, Pat62, Rao65a, Sed66, SW62, Tia66, Wis62, HF69, Ste66]. Applications [Bar65d, Buc65, BO59, BO61, Dav64a, Joh60a, JY60, KZ63, Loy67a, Mar67a, Mil63, Tar68, You67b, Blo60b, Ait66a, Bar62e, Buc65, Cox67a, Har63b, Mil63]. Applied [BT64a, Ken64, Sis64, Ryb66, Saw61b]. Approach [BT64a, BT68, Bro67, DS68, Kru67, Mar68c, DS76, RK67, Rut67]. Approaches [Bro64]. Appropriate [Hea64b]. Approximants [Sil69]. Approximate [Lin60, McG62, MB65, PT65, Rad67, Ste62b, Ste62c, Whi63, McG60, Yao65]. Approximating [Hoa68, Pea63b, Tik64, Tik66]. Approximation [Bai68a, Bha60, Cha67b, GJ68, Joh60a, Lea64, Pat65, Pat69, Pin62, Tar68, Tra69, Wis64, SZ60]. Approximations [DR63, Ewe63, GCP66, JY60, Kal66, KLTW65, Les60, Man69a, Ord68, Pea69, Rub64, San63, Tik65a]. Arbitrarily [Geh65b]. Archaeological [HSD66]. Arguments [Ano60g]. Arising [Cro64, Pea63b, Ram60, Ram61]. Arithmetic [Dav60j]. Armitage [Dav62e]. Aroian [Ano60f]. Arrangements [Ano65d, KZ63]. Ascending [Kri67]. Asger [Ano64e]. Aspects [Ano64h, Rei60b, Sok60]. Assay [AS66, Bar66f]. Assays [ASB60]. Assembly [Gre67]. Assessment [Men65]. Associated [Bar69, BG62, DB69, DB70, Har63c]. Assumption [Ewe61]. Assumptions [BT64a]. Assurance [MS69]. Astronomy [Dav63e]. Asymptotic [BG62, Bri69, Cha67b, Cha64, Far63, Gho66, Gle68, Gow61, KG67, Mil69, Pea66c, PP64, Rao69, RK69a, Sen67, Stu67, Tra69, Whi61, Rao68b]. Asymptotically [BN65]. Attente [Haw64]. attribute [BC61]. Aty [BD61]. Author [Ano69d]. Auto [Hea64a]. Auto-Primer [Hea64a]. Autocorrelated [TT66]. Autocorrelation [Dan62a, EK60]. Autoregressions [Whi63]. Autoregressive [Cha61, Cha64, Fin60b, Han69a, Han69b, Wal62]. Autoregressive-Moving [Han69b]. Average [Bri65, Fin60b, Gea66, Han69a, Han69b, Sha69, Stu67, Wal61a, Wal61c, Wal61b, Wal62]. Axiomatic [Joh60c]. Axiomatischen [Joh60c]. Aylmer [Ken63a].

B [Ano60i, Ano60j, Ano60c, Ano63j, Ano63i, Ano64h, Bar63d, Bar63b, Cor64a, Dav60c, Dav60b, Dav62b, Dav63d, Dav64c, Joh64, Lin63, Rei60b]. Back [Ano60a, Ano60b, Ano61a, Ano61b, Ano62a, Ano62b, Ano63a, Ano63b, Ano64a, Ano64b, Ano65a, Ano65b, Ano66b, Ano66c, Ano67a, Ano67b, Ano68a, Ano68b, Ano68c, Ano69a, Ano69b, Ano69c]. Bacterial [Gan62]. Bahadur [Sto68]. Bailey [Bar65d, Dav60g, Wal63]. Balance [Pea60, Pea61]. Balanced [Ati61, Ati62, Ber68, KLTW65, Mul65, Mul66, Pre66, Pre67, Pre69, Rag62, Sis68, SBJ68, TD68]. Balking [Hai60, GS65]. Ball [Smi60].

BAN [Dos62]. Bargmann [Joh66]. Barnett [Pea66c]. Barnicot [Dav65a]. Bartlett [Bar62a, Bar62f, Cox67a, Har69a]. Barton [Fel63, Gob67]. Based [Cro66, Cro67, Dem67, Dos62, Dur69, Gho63, Hol65b, Oka63, PS62b, PT65, PJ67, PJ68]. Basic [Ano64c, Loy66, Dra64, Wal64]. Basis [Moo61]. Batchelor [Bar60b]. Batches [Cam68]. Batschelet [Wat66]. Bayes [BT63, CK70, Mar68b, Boh66, BT62, Bro67, CK68, Kru67, LK69, Mar66, Mar67d, Mar68c, Mar69b, MK69, Pra66, RK67, Rut67, RK69a, RK69b, TZ64]. Bayesian [Bar66c, Bar65a, BB66, BT64a, BD65b, BT68, DG66, DG68a, DG68b, ST66, Sto69a, TT65, TT66, Tia66, TB67, TD68, TF69, Wet61]. Bayesians [Sto69a]. be [Kha78, RG66, Dav61i]. Beament [Kal61c]. Beckenbach [Ryb66]. Begin [Ken60b]. Behaviour [KD69, Kin61, MGR60, MGR61, Dav62c, Mal61b]. Behrens [Pat65, Pat69, dVW66, Yao65]. Bell [Sok60]. Benes [Bar65c]. Benjamin [Cor64a]. Berge [Har63a, Har63b]. Bergstrom [Tay64]. Berkeley [Dav62d, Dav63e]. Bernoulli [Cha68b, Eul61, Ken61a]. Bessel [Bar66e]. Best [CJ64, Zin61]. Beta [Amo63, Pea64, Seb63, Wis60, Dav60f]. Beta-Function [Wis60]. Beta-Variables [Dav60f]. Between [Ait66b, Atk69a, Bro64, CC67, DG67, EK60, Fai64, LG60, McG62, MB65, Mor67a, Pea68b, Rao66, RGL68, Sil64, Wis62, You67b, AH66, Ano60g, Ber61, PT65, Pea68c]. Bhattacharya [San64]. Bias [CO65, GZ67, RW66, Saw61a, SW62, Sti69, Swi68]. Biases [LB63]. Bibliography [Ano64d, Ano64c, Dav64b, Ken66b, Pea66b]. Binary [CC67]. Binomial [All66, All68, BH60b, Cha69, Cra62, Gon61a, Gon61b, Gon66, HG68, KG62, KM66, Lac69, Mor63a, Pat62, PW66a, Sha66, SW62, ST66, Tar79, Tar90, VW60, Vid64, Wis68, Wis90, LB65, MY61]. Bioassay [Bro61, GZ67]. Biological [ASB60, AS66, Bar66f, Les60, RA68, Sok60, Wat66, Cur66]. Biologistes [Ano64g]. Biology [Dav62d, Dav65a, Dav60g, Dav62a, Kal61c]. Biometrical [Smi62a]. Biometrie [Hol61b]. Biometrika [DS76, Har79, MT72, Ano69d]. biometry [Pea65a]. Birth [Bai68b, Blo60a, Blo60b, Dav64i, Goo67, Ken60a, Ker64, Pur66, Wil65a, Wil69, Wis62]. Birth-And-Death [Pur66, Dav64i, Goo67, Ken60a]. Birth-Death [Wil65a, Wil69, Wis62]. Bivariate [AH65, AH68, Ano60j, Ano60h, BJ69, Far60, Hol64, Hol66, Joh62c, Mar67c, Mar68a, Mar69a, Mor69, MM64, RGL68, Sno63, Sri60, Ste68a, TF69, AH66, Owe65]. Blanch [Ano60h]. Block [Bra54, Bra64, FK60, Joh65a, Mul65, Pre66, Pre67, Pre69, Rag62, Sis68, TD68, CB66, KZ63]. Blocking [FK60]. Blom [Dav60f]. Blyth [Cor62b]. Bolza [Ano63e]. Book [Ait66a, Ano60i, Ano60j, Ano60h, Ano60g, Ano60f, Ano60d, Ano60c, Ano60e, Ano63j, Ano63i, Ano63d, Ano63g, Ano63e, Ano63f, Ano63c, Ano63k, Ano63h, Ano64n, Ano64m, Ano64l, Ano64k, Ano64i, Ano64d, Ano64c, Ano64f, Ano64j, Ano64g, Ano64e, Ano64h, Ano66d, Ano67c, Arc62, Arc64, Arc61, Bar60b, Bar64a, Bar61b, Bar62b, Bar63a, Bar66b, Bar60c, Bar61c, Bar62e, Bar62f, Bar62c, Bar62d, Bar63d, Bar63b, Bar63c, Bar63g, Bar63f, Bar63e, Bar65d, Bar65e, Bar65c, Bar66c, Bar66g, Bar66f, Bar66d, Bar66e, Bea61, Bee63, Bro62, Buc60, Buc61, Buc62a, Buc62d, Buc62b, Buc62c, Buc64a, Buc64b, Buc65, Bur62, Cha68a, Clu63, Cor62b, Cor62a, Cor64a, Cox61, Cox67a, Cox67b, Cro60, Cur66, Dav60i, Dav60f, Dav60h, Dav60j, Dav60d, Dav60a, Dav60c, Dav60b]. Book [Dav60e, Dav60g, Dav61d, Dav61a, Dav61b, Dav61g, Dav61j, Dav61e, Dav61f, Dav61i, Dav61h, Dav61c, Dav62f, Dav62d, Dav62c, Dav62e, Dav62b, Dav62a, Dav63e, Dav63b, Dav63c, Dav63d, Dav63f, Dav63a, Dav64a, Dav64b, Dav64e, Dav64c, Dav64h, Dav64d, Dav64f, Dav64g, Dav65a, Dav65b, Dra64, Edw61, Ela59, Fel63, Fin62, Gob67, Gru61, Hal63, Har63b, Har63a, Haw64, Haw65a, Hea62, Hea64a, Hol61a, Hol61b, Jen64, Joh60c, Joh60b, Joh62f, Joh62e, Joh62c, Joh62d, Joh62a, Joh62b, Joh64, Joh66, Kal61b, Kal61c, Ken61b, Ken64, Ken66a, Ken66b, Kes61, Lin63, Lin69, Llo60, Loy66, Loy67a, Loy67b, Mal61a, Mal61b, Men65, Mil61, Mil63, Moo61, Moo63, Moo66, Mor62, Noe66, Pea64, Pea66b, Pea66c, Pea67a, Rad66, Rei60a, Rei60b, Ryb66, Saw61b, Saw61e, Saw61d, Saw61c, Saw62a, Smi60, Smi62b]. Book [Smi62a, Smi63, Sok60, Sto62, Stu64, Tay63, Tay64, Wal63, Wal64, Wat61a, Wat66, Wel62, Wet60, Wet62, Yat62]. Books [Ano60n, Ano60o, Ano61f, Ano61g, Ano62e, Ano62f, Ano63n, Ano63o, Ano64r, Ano64s, Ano65c, Ano65g, Ano66h]. Boole [Dav61h]. Borel [HB60, Hai61, Tan61]. Borne [Dow68]. Both [Bar67a, Jol65, Jol63]. Bound [San64]. Boundaries [MSA66]. Bounded [Ben63, Ben65, Ben68, Ste67a, Ste68b]. Bounds [BM65, Ben65, Lea64, RG57, RG61, SW60]. Bowker [Buc62b]. Box [Pea69]. Brailey [Rei60b]. Branching [Bar65e, Gan62, Jag67, Pol68, Wau68, Whi64a]. Brinley [Cor64a]. British [Tay63]. Brownien [Ken66a]. Brownlee [Wet62]. Bruce [BD66]. Brunk [Dav61a]. Buckland [Ano64c, Men65, Bar60c].

C [Ait66a, Ano63d, Ano63c, Bar63e, Bro62, Buc65, Cor62b, Cur66, Dav63a, Dav64h, Har63a, Joh62e, Llo60, Loy67b, Tay63, Wal64, Yat62]. Cajori [Smi60]. Calcoli [Dav61e]. Calculation [Gow66a]. Calculations [Dav60e]. Calculus [Ano63e, Dav61h, Dav64g, KZ63, Ano63d]. Calder [Cor64a]. Calibration [Tal69]. Canonical [Gow66a, Ksh62, Lan66, Whi63]. Capacity [Tan67]. Caps [Gil65]. Capture [Dar61, Jol65]. Capture-Recapture [Dar61, Jol65]. Carcinogenesis [Ken60a]. Care [Rei60a]. Carl [Ano64o]. Carl-Erik [Ano64o]. Carlo [CB66, HK68, Pea66a, Rob67]. Carrier [Dow68]. Carrier-Borne [Dow68]. Carslaw [Smi60]. Cartesian [Joh62c]. Case [BP63, Dav64h, Han68, Kal62c, Spr69, DH67b, Owe65]. cases [Put64]. Categorical [HL65, MA65]. Cell [Bar69, You62]. Censored [Ano64o, Dos62, Geh65a, Geh65b, HM66, HM67, HM69, Saw61a, Saw61f, Saw62b, Saw66b, Swa62, Tik67]. Censoring [GT69]. Census [Dar61, Seb62, Seb65]. Censuses [Dav61f]. Central [BDO60, BD61, JP69, Pri64a, San63, Seb63, Tik65b, Tik66, Amo64, HPW61, Owe65, SZ60]. Certain [Bar60a, Ghi68, KG62, LL60b, LB63, MGR60, MGR61, Mil69, Mot62, Pol66b, PP64, Wat66, Sis64]. Chain [BW63, BDF62, Con65, Goo64].

Chains [BW65a, BG60]. Chance [Ano64m, Fel63]. Change [BT65, DH69]. Change-Over [DH69]. Changeover [Ber68]. Characteristic [AG63, CO65, Chi67, Pil65b, Tur60, AG65, Men65, Edw61]. Characteristics [Bar60a, DG68a, Man69a]. Characterization [Pat63]. Characterizing [Kot66]. Chart [Pag62]. Charts [Moo61]. Chernoff [Mal61a]. Chi [Goo69, Goo70, GCP66, Har64a, Kha78, Oka63, Pea64, Pea67a, San63, Seb63, Sla66, Tik65a, Har65, Har79, SZ60]. Chi-Square [Har64a, Oka63, Pea64, Pea67a, San63, Sla66, Tik65a, Har65, Har79, SZ60]. Chi-Squared [Goo69, Goo70, Kha78, Seb63]. Children [Rei60b]. Choice [BD63, Dav69, EEG67, Mal61b, Ber61, BD65a]. Chorafas [Buc62d]. Chords [DF64]. Chromosome [BDF63]. Chronic [ZF69]. Chung [Bar63d]. Circle [BDF63, Bre63, DF64, Ste63b, Ste69a, Wat61b, Wat62, WW64]. Circles [Fai64, MD62]. Circular [Ajn68, Alw62, Ber69, BJ69, Gil65, Mor61, Sch69a, Wat67]. Clark [Cor64a]. Class [Kam65, Kam66a, Sub69, Sub70, WL67, Put64]. Classes [HK63]. Classical [BC60]. Classification [Ati64, BM65, BD68, Coo63, EEG67, Har69b, Har70, KW63, Lar69, Low64]. Classifications [Gho64b, KLTW65, MT67, Mos67a, MT72, Ano63h]. Claude [Har63b]. Closed [All66, All68, Hoe68, Pri64a, SA62a, SA62b, SA69, MSA66]. Closely [Hol65a]. Cluster [RS68, Sam62, Rob67]. Clustering [Nau65]. Clusters [Sam62]. Co [Bar63e, KT69]. Co-Ordinate [KT69]. Coefficient [Cox66a, Cox67c, Far63, Gho66, GW61, IMH68, JW65, KOR64, Whi61, Hof63, PS62a]. Coefficients [AH68, Dow66b, Far60, FP61, Gle68, Joh62d, Moh68, Mor63b, Pea68b, Sri60, PS62a, Joh62f, Joh62e]. coincidences [Ken68b]. Coleman [Dav62c]. Colin [Cor64a]. collected [Cox92]. College [Dav61c]. Collinearity [Ksh62]. Colloquium [Bee63]. Colony [Gan62]. Combinatorial [Ryb66, Fel63]. Comment [JW65]. Comments [Pea69, Pea68c]. Commerce [Ano60g]. Committee [Rei60a]. Common [Ano66d, BD65b]. Communication [Bar61b]. Comparative [Cha61, Sla66]. Compared [Wis64]. Comparing [DP60, Gar69, Geh65b]. Comparison [Bar65a, BB66, BGL60, BT64a, Cox66b, Day69a, Gle60, Ito62, Ito63, Pea63a, RT66, Saw66b, SS69, SD61a, SD61b, Tia66, TL67, Wan67, Mik65, dC69]. Comparisons [Alw62, Bra54, Bra64, DB69, DB70, KLTW65, MT67, PJ67, PJ68, Rei61, Rei62, Sla61, Ste69a, MT72]. Competing [Les62]. Competition [Arc61, McG67]. Complementi [Dav61e]. Complete [Wis64, Bai63, SW65]. Complex [Mor63b, Tur60, Clu63]. Component [Ber68, Ksh61, TB67, Wan67]. Components [Cun69, Day69b, Har69b, Har70, KW63, Ksh66, Low64, Mos67a, RT69, TT65, TD68, Wil62a]. Composite [Bha60, Rao66]. Compound [Mar68c, Mos63, Mos62]. Compounds [Kas60]. Computation [Har69a]. Computations [Bro62]. Computer [Hea64a, Joh66, Yat62]. Computers [Pea66c]. Computing [Imh61, Imh62]. Comrie [Dav63a]. Concentric [Fai64]. Concept [Bee63]. Concepts [Cor62a]. Concerning [Sha68]. Concomitant [Atk69b]. Concurrence [Saw66a]. Condition [RG66, Wau68]. Conditional [Bri66, Bro64]. Conditions [Goo69, Goo70, Kha78]. Confidence [BH60b, BH61b, Har64b, KOR64, Lac69, RG57, RG61, ST66, Wil62a]. Confounded [HG61]. Confounding [Mul66]. Congestion [Buc62c]. Connected [Cro68, LL60a, Ste66]. Consecutive [Wal61a, Wal61c, BC61]. Consequences [Wal60]. Conservative [Saw66a]. Considering [Mor65]. Consistency [Fra65]. Consistent [McN66]. Constants [Cun69]. Constructing [Dur61, Dur66, Mul65]. Construction [Rad64]. Contagious [HG68]. Containing [GS62a]. Contingency [BH61a, IK68, Mar67c, Mar68a, Mar69a, Ste68a, BH60a]. Contingency-Type [Mar67c, Mar68a, Mar69a, Ste68a]. Continuation [Pra66]. continuity [Pea67b, Pla64]. Continuous [EK60, Kam65, Mar67d, Mar68b, Pri63, Sil69, Ver61]. Contrasts [Sch53, Sch69b]. Contributions [Blo63, Dav62d, Mar67c, Mar68a, RG57, RG61, Bar63c, Buc62a, Dav64c]. Control [Moo61, Pag62, Sti69, Wet60]. Convergence [Bri62]. Conversion [Ano60d]. Convex [Efr65, LS68]. Convolutions [Tay64]. Correction [Ait71, All68, Bar62a, BGKO71, CK70, Cox67c, Cro67, Dan70, Dur66, Fin61, HM67, Hin70, JP70, Pre69, Sch67, Ste68b, Pla64]. Corrections [Ati69, DB70, DS76, Goo70, HM69, Har70, LMH71, Mar68a, Mar68b, MT72, PG70, Rao69, Sch69b, SA69, Sub70, Tar90, Wal70, Wis90, Har79]. Correlated [GS62b, GPS64, GP65, pH68, Hin69b, Hin70, KLTW66, Lar69, Mar67b, Mit68, MR68, SH68, Spr65]. Correlation [AH65, Ait66b, AH66, AH68, ANR68, Ait69, Ait71, Cox66a, Cox67c, Dav60a, Dur69, Far60, Far63, FP61, Gho66, GW61, Gle68, Hil69, KLTW66, Mar67b, McG62, MB65, MG69b, Mor67a, Mor67b, Mor63b, RGL68, SH68, Sno63, Whi61, Whi62, Hof63, McG60, PS62a, AH65, AH66]. Correlations [Lan66, Mos63, Mos62]. correspondence [Pea68c]. Corrigenda [AG65, Ano60k, Bar65b, BH61a, BT63, BD65a, BW65b, BO61, Gon61a, Har61a, Har65, Imh62, Ito63, JNAP65, Kal62a, LEH61, MGR61, OI61a, Pea61, Ram61, Rei62, RG61, Rub65, She59, She61, SD61a, Wal61a, WGH63a, Wis63a]. Corrigendum [Bra64, Bur60b, Ver64]. Count [Hol67]. Countercurrent [Kas60]. Counting [Hai65a]. Counts [Mos63]. Course [Ano63c, Arc64]. Covariance [Ati64, Bri65, Fin60b, Han68, IS64, Kor68, Kor69, Mar63, PJ68, RT69, Sha69, Whi62]. Covariances [Gla62, Mor63b]. Coverage [Mil69, Mor61]. Covering [Gil65]. Cox [Rao69, Buc64a, Rao68b]. Craig [Sha66]. Cramér [Ano63g]. Craters [Mar67a]. Criteria [PJ67, PJ68]. Criterion [Bar62a, BT64b, PG69, PG70, Sch66, Sch67]. Critical [Gea66, GW61, Nic61, Stu67, Ver63, Ver64]. Criticism [JW65]. Cross [Mor63b]. Cross-Correlation [Mor63b]. Crossed [Ano63h, Mos67a]. Crosses [HK63]. Crossover [Gar69]. Cruon [Mil63]. cuboidal’ [DL66, DL65]. Cumulative [Dav64i, Tar68]. Curvature [Rao60]. Curve [Gon61a, Gon61b, Mil61, Mor66, PR64, SC66, Wil65b, Pat60]. Curves [JNAP63, LM64, PT65, Pea66c, Rao65a, Wis60, JNAP65, Joh65b, WCV66]. Customer [Fin61, Fin60a]. Cut [KLTW65]. Cybernetic [Bar63a]. Cycle [RA68]. Cyclic [DH69].

D [Ano60i, Ano60j, Ano63j, Ano64l, Bar63e, Bar66c, Bar66f, Buc62d, Buc64a, Buc65, Dav60h, Dav61d, Dav61a, Dav61g, Dav63d, Fel63, Gob67, Hea62, Joh64, Lin63, Lin69, Pea66c, Smi62b, Stu64, Yat62]. Daniel [Ano64g, Bee63, Ken61a]. Daniels [Far63]. Darwin [Sok60]. Data [Cox66b, Geh65a, HL65, HG68, HSD66, Jol65, Kal66, Les62, Lit68, McN68, Ohl64, Saw62b, Sto69b, WG68, Jol63, MA65, Ano63i, Bar62c, Buc60]. d’Attente [Mil63]. David [Bar61b, Fel63, Gob67, Wel62, Saw60]. Death [Bai68b, Blo60a, Blo60b, Dav64i, Goo67, Jol65, Ken60a, Mar67a, Pur66, Wil65a, Wil69, Wis62, Jol63]. Decimal [Ano60g]. Decimalisation [Dav64h]. Decision [Kru67, Mal61a, Mar68c, Saw61b, Sto62, LB65]. Decomposition [Kab64]. Defares [Dav62a]. degrees [Yao65]. Delays [Tan62, Yeo64]. Densities [Hoa68]. Density [Dan62a, Hoa68, Mil65b, Whi63]. Department [Ano60g]. Departures [Ewe61]. Dependence [BDF62]. Dependent [Goo67, Pha65, Rao69, Saw62b, Wau68, You67b, Put64, Rao68b]. Derivation [Ber69, KK66, Ksh60, LM64, Sil69, Tan61]. Derivations [KLTW65]. Derivative [Joh62d]. Derived [Pri63]. Deriving [Rad64]. Design [Boh66, BD63, CB63, Dav69, DH66, DH67a, Pat68, SBJ68, TB67, TD68, BD65a, CB66, DH67b, Saw62a]. Designs [Ati61, Ber68, Bra54, Bra64, DH69, Day69a, DL65, FK60, Gar69, Hex63, KZ63, LS68, Mos67a, Mul65, Pea68a, Pre66, Pre67, Pre69, Rag62, Ree69, Sis68, DL66]. Detection [Zin61]. Determinants [Fin60b]. Determination [Hea64b]. Determine [HM60]. Deterministic [Fin60a, Fin61, Gan62, Jol63]. Development [FMB68, Pea67b, Sea67]. Deviates [KLTW66, Pea66c, Dav63f]. Deviation [Kam65, Kam66a, Kam66b, Mea66, PS64, Tik67, You67a]. Deviations [Ben63, PT65]. Dexter [Joh62e]. Diagrams [Gow68]. Diallel [HK63]. Dialysis [Kas60]. Dichotomized [HT65]. Dichotomous [EEG67]. Dictionnaire [Cor62a]. Difference [Dav68, Har67a, Wat61a]. Differences [BP63, Ram60, Ram61, Ano63d, Dav61h, Joh62d, Tay63]. Different [DP60, Dow66a, McG67, Roa63, Sil64]. Differential [Bar63e, Dav61j, Gan65, Wat61a]. Diffusion [Ewe63, Ken66a, Whi62]. Digest [Mor62]. dilution [Jol63]. Dimensional [Bar64b, Pri64b, RS68, Bar65b, Rob67]. Dimensions [Dow66a, Nau65, Arc62, Arc64]. Direct [KZ63, Mor63a, Pol66b]. Direction [Ksh62]. Directions [Ste62b, Ste62c, Ste69b, Ste66]. Discarding [BKM67]. Discontinuous [Pri64b]. Discrepancies [Eri67]. discrepant [Ber61]. Discrete [Bar60a, Kam66a, Les60, Mar66, Ord67, Pri63]. Discrete-Time [Les60]. Discriminant [Rad67, Rao66]. Discriminating [Atk69a]. Discrimination [BP63, CC67, CRR60, GD68, Han68, LMH67, LMH71]. Disease [Mor65]. Diseases [ZF69]. Dispersion [Rao69, Sel65, Ste67b, Ury66, Rao68b]. dissertation [Eul61]. Distance [Gow66b, Hol65a, Hol65b]. Distances [Fai64, PT65]. Distinction [Bro64].

Distributed [Ait63, Bai68b, Kha78, Mor67b, Stu62]. Distribution [AH65, Ait66b, AH68, Ajn68, AC64, Alw62, Amo63, Ano60h, Ano60f, BDO60, BN68, Ber69, Bha60, BJ69, BGKO66, BGKO71, Bre63, Bur60a, Cam68, CB63, Con65, Cox66a, Cox67c, Cro66, Cro67, Far60, Fin64, GJ68, Gho66, Goo69, Goo70, Gow61, GPS64, GG66, HJ63, HT65, HM60, Har67a, Har64a, Har67b, Has69a, Has69b, Hea64b, Hol64, Imh61, Imh62, Joh60a, JY60, JW65, Joh66, Kab64, KK66, Kor68, Kot65, Kot66, Kri67, Ksh66, LEH60, LEH61, MS69, Mar68b, Mar69b, Mar63, McG62, MB65, Moh68, Mos63, MM64, Nab61, Nic61, Ord68, Pat62, Pat63, Pat69, Pea64, Pea69, Pil64, Pil65b, PG69, PG70, Pin62, PW66b, Rao60, RGL68, Roa63, San63, SW62, Sil69, Sno63, Sri60, SO62, Ste62a, Ste63a, Ste64, Ste67b, Ste69b, Ste69c, Sto67, Tan61]. Distribution [Tar79, Tar90, Tho69, TT65, TF69, Tik66, Ver61, Vid64, Wat67, WW64, WGH62, WGH63a, WGH63b, Wil65a, Wil69, Wis64, You67a, AH66, HB60, Hai61, Har65, Har79, HPW61, Hol66, Joh61, KK65, Mos62, Owe65, RK67, SC65, Ste66, PS62a, Ste65a, Ano60j, Ano63j, Pea67a]. Distribution-Free [WW64]. Distributions [Ait63, Bar60a, Bar69, Bar63d, BGKO66, BGKO71, Cha67b, Cha69, Day69b, Dow66a, DG67, DH67a, Hai65a, HG67, HG68, Hol69, Kam65, Kam66a, Kam66b, KG62, KOR64, Lac69, Mar67c, Mar68a, Mar69a, Mar66, Mar67d, Mor69, OS65, Ord67, Pat62, Pea63b, Pol66a, PW66a, Pri64a, Ram60, Ram61, Sch69a, Sch66, Sch67, Seb63, Sel64, Ste68a, SS65b, Sub69, Sub70, Tho69, TL67, Tik64, Tik65a, Tik65b, Wat65, Wis63a, Wis63b, vHG66, Amo64, DH67b, Hof63, MY61, Pat65, Ano63g, Pea64]. Divisible [HK63]. Do [Dav65b]. Dodge [Ano60e]. Doig [Ano64d, Pea66b]. Donald [Dav61j]. Doris [Hea64a]. Dose [Wis60]. Dose-Response [Wis60]. Double [Ano60e, CJ64, KLTW65, SS65a]. Doubly [Ait66b, Geh65a, HM66, HM67, HM69]. Drawings [Cra62]. Drawn [Rao60, Roa63]. Drug [Col63, Kin64]. Drugs [AS64]. Due [Kal64]. Duncan [Mal61b, Wet60]. Durbin [Wal61a, Wal61c]. Durfee [Dav61c]. Dwight [Ano63i].

Each [Roa63]. Eagle [Dav61i]. Earl [Ano64l, Dav61g]. Early [Ano64e, She68, Pea65a, Pea68c]. Ecology [Bar62f]. Economic [Buc65, Cor62b, Tay63, Wis62]. Economical [VW60]. Économique [Cor62a]. Edgeworth [Ken68a]. Editorial [Ano61c, Ano65d]. Edmond [Cor62a]. Edmund [Bar62d]. Edward [Wat66]. Effect [BN64, Hai65b, TT65, TT66, MA65, TT66]. Effectiveness [Gle60]. Effects [Ber68, Lar69, Mos67b, Pol66a, DG66]. Efficiencies [Ano64h, Cha61, Rao69, Rao68b]. Efficiency [Cha68b, Dos62, Far63, FK60, Gea66, KG62, Pea68a, Sen67, Sto68, Stu67]. Efficient [Saw62b]. Ekambaram [Moo61]. Elderton [Pea62]. Elementary [Bar63b, Pil65a, Mal61a, Saw61c, Wel62]. Elements [Bar65d, Buc65, Mil63, Dav63b, Smi63]. Elizabeth [Joh62e]. Elliptic [Dav61i]. Emanuel [Arc62]. Emphasis [GPS64]. Empirical [BGL60, CK68, CK70, CB63, Kru67, LK69, Mar66, Mar67d, Mar68b, Mar68c, Mar69b, MK69, RK69a, RK69b, RK67, Rut67, AH68]. Employing [HK68]. Enders [Bar63a, Fin62]. Engineering [Bro62, Moo66, Buc62d, Wet62, Buc62b]. Ensemble [Rob69]. Entwistle [Hea64a]. Epidemic [Bai68a, Dow67, Gan65, HS69, Ohl64, Sev69, Sis65, Wil65b, Bai63]. Epidemics [Dow68]. Epidemiological [Rei60b]. Epidemiology [Bar62f]. Episodes [Ano64e]. Equal [IS64, MGR60, MGR61, Sla66, You62]. Equality [Gle68, GM62, Kor69, MG69b, PJ68]. Equally [Spr65]. Equation [Bar66a, BDM60, BDM63, Gan65, Kal61a, Kal62a]. Equations [Kal62c, Phi66, Dav61j, Wat61a]. Equatorial [Wat65]. Equi [Pea68a]. Equi-Replicate [Pea68a]. Equicorrelated [SO62, Ste62a]. Equiprobable [Cro68]. Ergodic [Kin61, Par62]. Erik [Ano64o]. Erlang [EF62]. Ernest [Dav62c]. Error [Lea64, Ver61, Wal60, Saw60]. Error-Scale [Ver61]. Errors [Han63, TT66, Ver61]. Especially [PR64]. Estermann [Clu63]. Estimated [Cur60]. Estimates [AC64, BMS67, Bri69, Cor64b, Dav60f, Dos62, Dow66b, Jol63, Jol65, Low64, RT69, Saw61a]. Estimating [Bat63, Cha61, Cha69, Day69b, HG67, Man69b, McN66, Mea66, Ohl64, Tik67, MG69a, RK67, Rut67]. Estimation [AD68, Ati62, Ati68, Ati69, Bar61d, Bar63c, Big61, BC60, BD65b, Cun69, DB69, DB70, DH66, DH67a, Gil60, GG66, HT65, Han67, Han69a, HM66, HM67, HM69, HR67, HR68, Har69b, Har70, Has69a, Has69b, Hex63, Hit62, Hol64, Kal62b, KG62, Lit68, Mar66, Mar67d, Mar68b, Mar69b, Mor65, MM64, Mos67a, Pat62, Rad66, Rao62, Rao63, Rao65b, RW66, RW64, RK69b, Saw61f, Sub65, Swa62, Swa63, TF69, Ver61, Wal61b, Wal62, WD67, WL67, WGH62, WGH63a, WGH63b, Wil63, DH67b, WCV66]. Estimator [Bar66a, Bro61, Kal66, Rao67]. Estimators [BG62, CK68, CK70, GZ67, Gla62, Joh69, MK69, RK69a, Saw62b, She58, She59, She61, SW62, Swi68, Wan67]. Eugene [Hol61b]. Evaluation [Bar66a, MS69, Joh61]. Event [WD67]. Events [BDF62, Goo64, TR66]. Evolution [Dav65a]. Evolutionary [Bar66g]. Exact [BH61a, Dur61, Dur66, Gar69, Har64b, Ksh60, PG69, PG70, RT69, Sch66, Sch67, Ste62b, Ste62c, Sto68, BH60a]. Example [Cox66b, Eri67]. Exercises [Wel62]. existence [HF69]. Expansions [Cha64, Dan62b, Gho66, HM60, Lea63, McF60, Whi61]. Expectation [Ben63, Bri66]. Expectations [Ano63h, Mor63b, Sil64]. Expected [GR69, Har61a, Har61b, SC66, Sla66, Tra69, Saw60]. Experimental [FMB68, Kal61c, KD69, Les62, Saw62a, Sto69a]. Experiments [Big61, Dav69, DH66, DH67a, HG61, HSD66, Joh65a, McG67, Mul66, SS69, SD61a, SD61b, DH67b]. Explicit [Jol65]. Exponential [BDM60, BDM63, Hoe68, LM64, Mit68, MR68, Pat63, Pea66c, Pat60]. Exponential-Type [Pat63]. Exponentiality [PP64]. Exponentially [Mor67b]. Expressed [JNAP63, Pri64a, JNAP65]. Expressing [Mar63]. Expressions [OS65]. Extended [Ver63, Ver64]. Extension [Cha67a, DB69, DB70]. Extensions [Fai68]. Extinction [Gan62, Goo67, Wau68]. Extrema [BO59, BO61]. Extremal [Bri66, HJ63]. Extreme [HT68, Has69a, Sin67, Sto67, Sto68, TW63]. Extremes [Gre67, Dav60i]. Ezekiel [Dav60a].

F [Ano60e, Bar62d, Dav61f, Dav63b, Dav64f, Dav65b, Fel63, Gob67, Kes61, Loy67b, Ryb66, Smi60, Wel62]. facilitate [Joh65b]. Factor [Rad66, Spr69, Stu64]. Factorial [HG61, KZ63, Mul66, Pat68, Cox92]. Factorization [Cla66, Kel62, Rad67, Whi63]. Factors [Ksh62, Joh64]. Falconer [Smi62b]. Families [Jac68, Wal67, Wal70]. Family [SA62b, SA69]. Fate [Ken61b]. Feller [Loy67a]. Fiducial [Fra61, Fra64, Fra65]. Field [Bro62]. Fieller [Cox67d]. Figures [Smi60]. Finding [Ric67]. Finite [Ano63d, BD61, Bri65, Con65, Dav61h, Dem67, Rao68a, Sto69a, Wat61a]. Finite-Difference [Wat61a]. Finite-State [Con65]. Finney [Bar66f, Hea62]. First [Ano63c, Ber68, Cox66a, Cox67c, Hoa68, KG67, RT69, Sha69, SW60, You67a]. First- [RT69]. Fisher [Ken63a, Pat65, Pat69, Pea68c, Ste69b, Yao65]. Fisz [Bar64a, Ela59]. Fit [Cha64, Cro68, Ksh61, PS62b, Sla66, Ste63a, Ste69a, Tar79, Tik65a, Wat61b, Wat62, Put64, Ste64, Ste65a, Tar90]. Fitting [Ber60, Cro64, Cun69, Gon61a, Gon61b, Har67b, LM64, Pea66c, Spr65, Whi63, Wis60, Joh65b, Pat60, Mil61]. Five [Ano60g]. Fixed [DS68, Lar69, DS76, DG66]. Fixed-Effects [Lar69, DG66]. Fletcher [Dav63a]. Flow [Haw65a]. Fluctuating [Pol66a]. Follows [Kot65]. Food [Cor64a]. foregoing [Eul61]. Form [Chi67, Goo69, Goo70, Kha78, Pri64a, Ver61]. Forman [Buc60]. formation [Ber61]. Forms [BO59, BO61, Coo63, Imh61, Imh62, Sha68, Tik65b, Tur60]. Formula [Rao60, Wal61a, Wal61c]. Formulae [BD61, Sis64]. Forsythe [Wat61a]. Foundation [Rad66]. Foundations [Ano64k, Joh60c]. Four [PJ67, PJ68, SW60, Sno63]. Fourth [Dav62d, Dav63e]. Fox [Ano64c, Dav60a]. Fraction [Rub64]. Francis [Ken68a]. Frank [Haw65a]. Fraser [Dav60h, Lin69]. Fred [Yat62]. Free [Cro66, Cro67, Tho69, WW64]. freedom [Yao65]. Freeman [Ano64f, MY61]. Frequencies [Oka63, Sla66, You62]. Frequency [Gon61a, Gon61b, Gon66, Hol69, PT65, Roa63, Joh65b]. Frequentist [Bar65a, BB66]. Friedman [Sen67]. Front [Ano60l, Ano60m, Ano61d, Ano61e, Ano62c, Ano62d, Ano63l, Ano63m, Ano64p, Ano64q, Ano65e, Ano65f, Ano66e, Ano66f, Ano67d, Ano67e, Ano68d, Ano68e, Ano68f, Ano69e, Ano69f, Ano69g]. Fulfilment [You61]. Function [Ano60h, Ano60f, BH61a, Ber60, Cro66, Cro67, Dan62a, Gob67, Hit62, Hit66, LS68, Lea64, Pea64, Pea67a, Rao66, Sil69, Tur60, WD67, Wis60, BH60a]. Functional [Dav64i]. Functions [CO65, Dav61i, Gon61a, Gon61b, GPS64, GP65, HM60, LL60b, Mil65b, Mos67a, Ord68, Pil65a, Ric67, Whi62, Bai63, SZ60, Ano60h, Clu63, Dav61g, Edw61, Bar66e]. Fundamentals [Dav61c]. Furst [Bar63f]. Further [BT62, BT63, MR68, Rao63, RG57, RG61, SM68, Pat60].

G [Ano60e, Ano63j, Ano64d, Arc64, Bar63c, Buc62b, Buc65, Cox67b, Dav60b, Dav61d, Dav62f, Dav62a, Dav64c, Dav64d, Dav64g, Dav65a, Dra64, Gob67, Men65, Mil61, Pea66b, Rad66, Saw61c]. Gall [Haw64]. Galton [Pol68]. Games [Arc61]. Gamma [Ait63, Kot65, Mor69, Pea64, Pea67a, Pin62, RG66, Stu62, WGH62, WGH63a, WGH63b]. Gamma-Distributed [Ait63, Stu62]. Gamma-Function [Pea64]. Gantmacher [Kes61]. Gaps [Haw65b]. Garvin [Bea61]. Gary [Yat62]. Gass [Buc65]. Gauss [Sea67]. Gaussian [Mil65b]. Geary [Buc65]. Gebiete [Ano63k]. Gene [Pol66a]. General [AS64, Dow67, Far60, Joh60c, LS68, PB64, Saw64, Sis65, Swi68, Tik64, Whi64b, Bar65c]. Generalization [Cox67d, Fra66, Kam66a]. generalizations [Hol66]. Generalized [EF62, Far63, Geh65a, Geh65b, HG67, Pat62, Pat69, PR64, Ram60, Ram61, Wal61a, Wal61c]. Generating [Dan62b, Dav64i]. Genetic [BW63, Ewe63, Gan61, Wal63, Wil62b]. Genetical [Gru61]. Genetics [Bat63, Ken66a, Smi62b, Smi62a]. Genevieve [Joh62c]. Geometrical [Phi66, Dav64d]. Geometry [Arc62, Arc64, CRR60, Rao62, Rao63]. George [Dav61h, Wat61a]. Gerhard [Bar63f]. Gershenson [Moo63]. Gertrude [Ano60h]. Ghosh [Joh66]. Ghouila [Har63a]. Ghouila-Houri [Har63a]. GI [Fin61]. GI/M/1 [Fin61]. Gini [Dav68]. Girdle [Sel64]. Given [IK68, JNAP63, Saw61a, Saw61f, SW60, JNAP65]. Gmedenko [Bar63d]. Gnedenko [Bar63b, Lin63]. Goldberg [Saw61d]. Goodness [Cha64, Cro68, Ksh61, PS62b, Sla66, Ste63a, Ste64, Ste69a, Tar79, Tar90, Tik65a, Wat61b, Wat62, Put64, Ste65a]. Goodness-Of-Fit [Ste69a, Tik65a, Wat61b, Cro68, Ksh61, PS62b, Sla66, Ste63a, Wat62, Ste64, Put64, Ste65a]. Gosset [Pea68c]. Graig [Dav61b]. Graphs [Har63b]. Graybill [Dav64f]. Greater [Rei60a]. Greenberg [Dav64c]. Greenspan [Dav61j]. Greenwood [Dav63c]. Grenander [Dav60d]. Grid [Joh62c]. Group [Fra65, HK63, DG66, HF69]. Grouped [Bar63c, Kal66, McN66, McN68, Swa62, Swa63]. Grouping [Kal64]. Growth [Goo68, Pea66c, PR64, Rao65a, Cor64a, Dav65a]. Grundlagen [Joh60c, Bar62d]. Guest [Mil61]. Guide [Dav63c]. Gumbel [Dav60i]. Gunnar [Dav60f]. Gupta [AC64]. Guttman [Moo66].

H [Ano60d, Ano60e, Ano63i, Ano63g, Ano64n, Ano64f, Bar60b, Buc62b, Dav60j, Dav60e, Dav61d, Dav61a, Dav61c, Dav62b, Dav63c, Joh62f, Llo60, Pea64, Pea67a, Smi60]. Hacking [Bar66b]. Haight [Haw65a]. Half [Har67a]. Half-Square [Har67a]. Handbook [Dav63d, Dav64e, Noe66]. Hardy [Rei60b]. Harold [Buc62a, Tay64]. Harriet [Rei60b]. Harris [Bar65e]. Harrison [Dav65a]. Harry [Moo63]. Harter [Pea64]. Hartley [Bar62a, Dav63c]. Hausa [DR63]. having [BC61]. Hayden [Bee63]. Hazard [WL64]. Health [Hol61a, Rei60a, Rei60a]. Henri [Kal61b]. Henry [Ken66a]. Herbert [Bur62, Dav62c, Joh62e, Joh62c, Joh62d]. Herdan [Dav64g]. Herman [Ken66b, Mal61a]. Hermite [KK66, KK65]. Hermitian [Tur60].

Heterogeneity [LB66, MGR60, MGR61, Pea66a]. Hierarchical [Blo60a, Blo60b, Gho64b, TB67]. High [Mor68]. High-Speed [Mor68]. Higher [LL60b]. Hiorns [Pea66c]. historical [Sea67]. History [Ano64e, Has67, Ken60b, Ken61a, Ken61b, Ken63b, Rab69, She68, Ken68b, Ken68a, Pea65a, Pea67b, Pea68c, Sea67, Ken63b]. Hodges [BJ69]. Hoel [Saw61c]. Hogg [Dav61b]. Homogeneity [Bar62a, pH68, PW66a, PW66b, Wis68, Wis90]. Homogeneous [Con65, Pur66, Wil69, Wil64]. Hotelling [Buc62a, Lan66]. Houri [Har63a]. Household [Ohl64]. Households [Mor65]. Howard [Saw61b]. Hull [Efr65]. Human [Bat63, Dav65a, Dav65a]. Hypergeometric [Ano63j, Ord68]. Hypotheses [Bat63, BN65, Bha60, Hit66, Jac68, Mar68c, McN66, Mos67b, PJ67, Rao66, Seb64, Wal67, Wal70]. Hypothesis [DP60, GW61, MGR60, MGR61, PB64, Saw64, Thi67, Mik65, You62, Cro60]. Hypothetical [Ksh61].

Ian [Bar66b]. Ideas [Dra64]. Identification [Han69b]. if [Bat63]. II [Bra64, Dav62f, Dav63e, AH65, BB66, Bha60, Blo60b, Bra54, FP61, Hai60, MB65, PW66b, Saw61a, Saw62b, Ste62c, Ste64, TT66, Wat62, Loy67a, Noe66]. III [Dav63e, AH66, PS62a]. Immigration [Jol65, Mar67a]. Immigration-Stochastic [Jol65]. Impatience [Fin61, Fin60a]. Importance [BT64a]. Improvement [Wis64]. inadequacies [DL65]. Inanimate [Wil69]. incidents [Pea65a]. Incomplete [Amo63, Bra54, Bra64, FK60, Joh65a, Mul65, Nab61, Pea64, Pea67a, Pre66, Pre67, Pre69, Sis68, TD68, Wis60]. Incompletely [LB63]. Inconsistencies [Sla61]. Incorrect [Sis68]. Independence [GW61, KT69, OI61a, Thi67, OI61b]. Independent [Ait63, Bar63d, Ben63, Ben65, Ben68, Kot65, SS69, Stu62, WD67]. Independnece [Mar69a]. Index [Ano69d, Dav63a, Sel65, Wil62b, Wil63]. Indian [Ano64h]. Individual [Mal61b]. Individuals [Whi64a]. Induced [Seb64]. induction [Ber61]. Industrial [Wet60]. Inequalities [AG63, Ben65, Har67b, Ste67a, AG65, Kin62]. Inequality [Ben68, Dan62a, Mal68]. Infectives [HS69]. Inference [Ait66a, BW63, BW65a, BT64b, Dav62f, Fra64, Hin69a, Mor69, Ano64k, Ano64f, Bar66b, Lin69, Smi63]. Inferences [Bar65a, BB66, Dem67, Eri67, LEH60, LEH61]. Infinitely [Fin62]. Information [Ano60p, Ano61h, Ano62g, Ano63p, Ano64t, Ano65h, Ano66i, Ano67h, Ano68j, Ano69k, Kal64, Swa62, Swa63, MG69a, Ano64o]. Informative [Eri67]. Initial [HS69]. Input [EF62]. Inqualities [Ste68b]. Inspection [Ano60e, EK60]. Insurance [Rei60a]. Integral [Chi67, HM60, McF60, OS65, Pea67a, Phi66, Pin62, Pur66, SW60, Son61, Joh61, SC65, Ano60i]. Integrals [Ano63i, Jag67]. Integration [Joh62e]. Interaction [KW63, LG60]. Interactive [AS66]. interest [DL65]. Interference [Blo63]. Intermediate [Thi67]. International [Mor62, Bar60c]. Interpolation [Joh62f, Joh62c, Pea68b]. Intersection [Hin69a, Tan62, Tan67]. Intersections [DF64]. Interval [AD68, Bri65, Cox60, Wil62a]. Intervals [BH60b, BH61b, KOR64, LM63, LMH67, LMH71]. Intraclass [OI61a, OI61b]. Introduction [Bar61b, Bar63b, Cox67a, Dav61a, Dav61b, Dav62a, Dav64a, Dav65a, Fin62, Hea62, Loy67a, Mil63, Saw62a, Ano64l, Cur66, Dav60h, Saw61d, Ano64f, Bar66c, Bea61, Buc62c, Dav60b, Dav62b, Dav64f, Hol61a, Loy67b, Smi62b, Wal63]. Introductory [Moo66]. Invariance [Fra61]. Inventories [Loy66]. Inventory [Gho64a]. Inverse [Ait63, GR69, Mor63a, Pat64, Sha69, Sil69, Tar68]. Inversely [Mos63]. Investigation [CB63, HK68, Sri60, MA65]. Involving [Cox66b, Mal68, Pat69, SS65b]. Irrational [Rub64]. Isaac [Ken63b]. Isotropic [Ksh61]. Item [Bur62]. Items [Cam68]. Iteration [Kal61a, Kal62a, Kal62c]. Iterative [HM66, HM67, HM69]. Ito [Ken66a]. IV [Dav62d, AH68, SC65].

J [Ano63j, Ano64k, Ano64i, Bar63g, Bar65d, Bar66f, Bro62, Buc61, Buc62b, Dav60i, Dav60g, Dav61d, Dav62a, Dav63b, Dav63c, Dav63a, Dav65a, Hea62, Jen64, Joh60c, Kal61c, Llo60, Loy67b, Mor62, Smi60, Wal63, Wet60, Yat62]. Jackson [Dav60c]. James [Bar60b, Dav62c]. Janet [Rei60b]. Je [Hol61b]. Jerzy [Dav62d, Dav63e]. Jeux [Har63a]. Johann [Joh60b, Joh62a]. John [Ano64m, Dav64e, Noe66, Cox92]. Johnson [Dav60c, Kam66a, Saw60]. Joint [AS64, AS66, Bro64, Kri67, Moh68, Ver61]. Jordan [Ano63d]. Joreskog [Rad66]. Journal [Bar60c, Ano60c, Ken64]. Judging [Sch53, Sch69b].

Kapur [Llo60]. Karl [Pea68c]. Kaufmann [Mil63]. Kempthorne [Smi62a]. Kendall [Ano64d, Cox67b, Dav62f, Dav64d, Gob67, Pea66b, Arc64, Buc65, Bur60a, Dra64, Hof63, Men65, PS62a, Sno62, Sno63]. Khamis [Pea67a]. Khatri [Sha68]. Khinchin [Bar63b]. Khinchine [Cox61]. Kimbro [Joh62c]. Kimura [Ken66a]. Kingman [Loy67b]. Kiyosi [Ken66a]. Knowledge [TZ64]. Known [Bar67a, Hoa68, Ver61, Bai63, Zin61]. Kolmogoroff [Ste68b, Ste67a]. Kolmogoroff-Type [Ste68b, Ste67a]. Kolmogorov [Bar63d, Lan66]. Koopmans [JW65]. Kopal [Ano63f]. Kozelka [Smi63]. Krishnaiah [Cha68a]. Kulldorff [Bar63c].

L [Ano63k, Ano64k, Ano64j, Bar63d, Cro60, Dav62b, Dav63a, Kal61c, Mil63, Pea64, Saw61e]. Lagged [Han67]. Lagrangian [BN65, Pea68b]. Laguerre [Tik65b]. Lanczos [Bar63e]. Landau [Bar62d]. Lane [Rei60b, Yeo64]. L’Anglais [Cor62a]. Language [Joh66, Joh66]. Languages [Ano64h]. Laplace [Ano64l]. Large [Bai68a, Bar62a, Ben63, Cam68, Hil69, She68, Ury66, Wal61b, Wal62, Wil65b]. Large-Sample [Ury66, Wal61b, Wal62]. Largest [Cha67a, Pil64, Pil65b, Pil67]. Latent [Gow66b, Mal61c, Mot62, TF69]. Latin [HF69]. Lattice [Hol65a]. Law [HT62, Kot66, She68, Whi62]. Lawley [Stu64, Mik65]. Laws [Har64b]. Layout [Zin62]. Least [Ati62, Dur69, Kel62, Rao65a, Swi68]. Least-Squares [Ati62, Dur69, Kel62, Swi68]. Lectures [Ano63e]. Left [Hoa68]. Lehmann [Cro60]. Lehrbuch [Joh62b]. Length [Mor66]. Leo [Ano60f]. Leon [Pea64]. Leone [Yat62]. Lerner [Bee63]. Level [BT65]. Levy [Ken66a]. Lieberman [Ano63j, Buc62b]. Life [MS69, Men65, RA68]. Lifetimes [Whi64a]. Likelihood [Bar66a, Ber60, HM66, HM67, HM69, HR67, Kal61a, Kal62a, Kal62c, Kal66, Ken61a, Pat62, PB64, Rao60, RT66, Saw61a, Sch66, Sch67, She58, She59, She61, WGH63b, dC69]. Likelihood-Ratio [PB64, Sch67]. likely [Ber61]. Limit [BD61, Bar63d, Tay64]. Limited [Ghi68]. Limiting [Wal61a, Wal61c]. Limits [Joh64, Lac69, ST66]. Lincoln [Mal61a]. Lindgren [Dav60b]. Lindley [Bar66c]. Line [Buc60, Pea63a]. Linear [AD68, BMS67, Bar67a, Bea61, Ber68, BN65, Buc65, Cha61, Cha64, CK68, CK70, CRR60, Dav69, Dow66b, DH67a, Gil60, GPS64, GP65, Har64b, Lea64, MK69, McG62, MB65, PB64, Rao62, Rao63, Rei61, Rei62, Saw64, Seb64, Swi68, TD68, DH67b, LB65, Sea67, dC69, Ait66a, Bar63e, Buc65]. Linear-Loss [AD68]. Lines [Pag62, Saw66a]. Linguistic [Dav64g]. Linkage [Wal63]. Lionel [Sto62]. Lisle [Cor62a]. Literature [Ano64d, Has67, Mor62, Pea66b]. Local [Ric67]. Locally [BN65]. Location [AD68, Fra64, Saw61a]. Log [Dav63f, KOR64]. Log-Normal [Dav63f]. Logarithm [BDO60]. Logarithmic [Cha69, Gow61]. Logarithms [Ano60g, Ano66d]. Logic [Ano64m, Bar66b]. Logistic [Ber60, GG66, Has69b, Hit62, Hit66]. Logit [GZ67, Hit62, Rei61, Rei62]. Lognormal [MM64]. London [Bar63e]. Longest [BC61]. Look [BT62, BT63]. Looking [Hil69]. Loss [AD68, Kal64, LS68]. losses [LB65]. Louis [Kal61b]. Lower [Dem67]. Luce [Mal61b]. Lukacs [Edw61]. Lunar [Mar67a].

M [Ano64d, Arc64, Bar62f, Buc64b, Buc65, Bur60b, Cox67a, Cox67b, Dav60j, Dav60a, Dav60e, Dav62f, Dav64a, Dav64d, Dav64f, Dav65a, Dra64, Ela59, Fin61, Gob67, Joh62c, Men65, Smi63]. Machine [Blo63]. Made [JW65]. Madow [BD61]. Making [LEH60, LEH61]. Management [Mor62]. Many [Fin62]. Marek [Bar64a]. Marginals [IK68]. Maria [Dav60j]. Marked [Cor64b]. Market [Arc61]. Markov [BW63, BW65a, BG60, Con65, Gil60, KM66, KG67, McG62, MB65, Thi67]. Marks [Wil69]. Martin [Arc61]. Matched [Cha68b, Gar69]. Mathematical [Ano60c, Ano63i, Ano63c, Bar64a, Dav61a, Dav61b, Dav63e, Dav63b, Dav63c, Dav63a, Ela59, Ken63b, Wal63, Pea67b, Cox61, Dav62c, Hal63, Haw65a, Llo60]. Mathematics [Ano60c, Dav62a, Ano64e, Ryb66]. Mathematische [Ano60k, Ela59]. Matrices [ANR68, Ait69, Ait71, AG63, Han68, Hil69, IS64, Kel62, Kor69, Mal61c, Mot62, Pil65a, PJ68, AG65, Kes61]. Matrix [CO65, Gan61, GR69, HT68, Kor68, Mar63, Pil64, Pil65b, Pil67, Pol66b, Rad64, Sha69, Whi63]. Matter [Ano60a, Ano60b, Ano60l, Ano60m, Ano61a, Ano61b, Ano61d, Ano61e, Ano62a, Ano62b, Ano62c, Ano62d, Ano63a, Ano63b, Ano63l, Ano63m, Ano64a, Ano64b, Ano64p, Ano64q, Ano65a, Ano65b, Ano65e, Ano65f, Ano66b, Ano66c, Ano66e, Ano66f, Ano67a, Ano67b, Ano67d, Ano67e, Ano68a, Ano68b, Ano68c, Ano68d, Ano68e, Ano68f, Ano69a, Ano69b, Ano69c, Ano69e, Ano69f, Ano69g]. Maurice [Pea66b]. Maxima [Ric67]. Maximum [Bar66a, Ber60, HM66, HM67, HM69, HR67, Kal66, Ken61a, Pat62, RT66, Saw61a, She58, She59, She61, WGH63b, dC69]. Maximum-Likelihood [Bar66a, HM66, HM67, HM69, HR67, Kal66, RT66]. Maxwell [Bar62c, Stu64]. McCarthy [Buc65]. McElgrath [Dav60b]. McKean [Ken66a]. Mean [Ano63h, BP63, BD61, BN64, BGKO66, BGKO71, Bri62, Cam68, Cra62, Dav68, GD68, Har67a, HK68, Kam65, Kam66a, Kam66b, Pea68a, Ram60, Ram61, Roa63, Tik67, Whi61, Wis62, You67a]. Means [Bro67, Fai68, GM62, MGR60, MGR61, MG69b, PT65, Tia66]. Measure [JNAP63, Loy67b, JNAP65]. Measurement [Dav62c, Han63, Joh60c, Moo63]. Measures [Ken62]. Measuring [Mor66]. Mechanisms [Has67]. Médecins [Ano64g]. Media [FMB68]. Median [Gar63]. Medical [Dav62e, Rei60a]. Medicine [Dav62a, Dav62d]. Member [Roa63]. Memoirs [Ano64i]. Mendelian [KM66]. Message [Sto69b]. Messens [Joh60c]. Method [Ati68, Ati69, Bar66f, Bee63, Bra54, Bra64, Cun69, Fra61, Gea66, Gon61a, Gon61b, HG68, Hit62, Hol67, Man69b, Mea66, Mul65, Pea66c, Rao65b, Sch53, Sch69b, Stu67, Pat60, WCV66, Bar60c, Stu64, Wal64]. Methoden [Kal61b]. Methodenlehre [Joh60b, Joh62a]. Méthodes [Ano64g]. Methodology [Wet62]. Methods [Buc65, Cha67b, Cha61, Cox61, Cox67a, Dav60g, Dav61f, Dur61, Dur66, Gho64b, Gow66b, Hol65b, KG62, LM64, Mil61, RW66, Sam62, Wat61a, Wat66, WG68, Ano64c, Dav60c, Dav60a]. Mickey [Rao67]. Micro [FMB68]. Micro-Organisms [FMB68]. Middleton [Bar61b]. Migration [Bai68b]. Mijares [Bar66d]. Miller [Dav63a]. Mills [Rub64, SC65]. Minimal [Rob68]. Minimax [BM65]. minimize [DL65]. Minimum [Hit62, Sha66]. Miriam [Rei60b]. misclassification [MA65]. Mises [Dow66a, DG67, Ste69c]. Missing [Big61]. Mixed [Big61, Han69a, Han69b, HR67, Hus69, Tho69, Tia66]. Mixed-Up [Big61]. Mixture [AS64, AS66, Day69b]. ML [BG62]. Mobile [Man69b]. Modal [Ste67b]. Mode [Fin64]. Model [BW63, BLN62, CK68, CK70, DB69, DB70, Gan61, HR67, Hus69, Jol65, Lar69, LG60, Les62, MK69, Mos67b, PR64, Swi68, Tia66, TB67, Cro61, DL65, Jol63, Sea67]. Models [AS64, AS66, Atk69a, BGL60, Bar62f, CC67, Fin60b, Goo68, KW63, Ken66a, LB63, Les60, Pol66b, RA68, Sev69, TT65, TT66, TD68, Wal61b]. Modern [Bar62e, Dav60c]. Modified [Pag62, SJG68, Sla66]. Molecular [Kas60]. Moment [Saw62b, She58, She59, She61, SW62, Sno62]. Moments [BD61, Ben65, Gho66, Hoa68, KG67, Lea63, Nab61, Pea63b, RT69, SW60, You67a, HPW61]. Monk [Ano60i]. Monographs [Ano64c, Smi60]. Monte [CB66, HK68, Pea66a, Rob67]. Mood [Dav64f]. Moran [Bar61c, Bar66g, Dav64d, Gan61]. Morgenstern [Arc61]. Mortality [Moo63]. Moser [Tay63]. Moses [Ano64j, Mal61a]. Most [BN65, VW60, Ber61]. Motoo [Ken66a]. Mouvement [Ken66a]. Moving [Bri65, Fin60b, Han69a, Han69b, Sha69, Wal61a, Wal61c, Wal61b, Wal62]. Moving-Average [Fin60b, Wal61a, Wal61c, Wal61b, Wal62]. MR [Har79]. Mulhall [Ano64n]. Multi [Pol68, Seb62, Ste69b]. Multi-Sample [Seb62, Ste69b]. Multi-Type [Pol68]. Multinomial [BN68, Fin64, Hea64b, Joh60a, JY60, Mal68, Mos63, OS65, PW66a, Wis63a, Wis63b, Wis64, You67a, Mos62]. Multiparametric [Kal62c]. Multiple [Bar66a, BDF62, Cla66, Goo64, KLTW65, Mar67a, MK69, MT67, Rei61, Rei62, Seb65, Jol63, MT72]. Multiple-Recapture [Seb65]. Multiplier [BN65]. Multiresponse [DH66, DH67b]. Multivariate [Arn64, BKM67, Cha67b, Chi67, DB69, DB70, FKM67, Gab68, Gow66b, Gow68, HT65, Ito62, Ito63, IS64, Joh69, Ksh60, Kud63, Mar67a, Mil65b, Nab61, Pil64, Pil65b, Pil65a, PJ67, Pil67, PB64, PR64, Rad64, RG57, RG61, SO62, Ste62a, Whi63, Joh61, Mik65, Mos62, Yao65, Cha68a]. Murray [Bar62b].

N [Bar63d, Buc62d, Cor64a, Dav60g, Dav62a, Dav65a, Fel63, Gob67, Llo60, Loy66, Stu64, Wal63, Wel62]. Natural [Ano60g, Ano66d, BDF60, Bar65d]. Nearest [Bro64, Hol65a, Rob69]. Nearest-Neighbour [Bro64]. Nearly [HJ63]. Necessary [Kha78, RG66]. Needle [Pra66]. Negative [Cha69, Gon61a, Gon61b, HG68, KG62, Mor67a, Mos63, OS65, Pea66c, SW62, Vid64]. Negro [Rei60b]. Neighbour [Bro64]. Neighbours [Rob69]. Nerve [BLN62]. Nested [Pre67]. Neville [Ano60d]. Neyman [Ano64i, Dav63e, BH60b, BH61b, Dav62d, Hol66, KG62]. Neyman-Shortest [BH60b, BH61b]. Nigeria [Ano64n]. No [Ano66d, Bar66e, EK60, Pea66c]. No.XXVIII [Pea66c]. Nomograms [Ber60]. Non [AS66, BDO60, BN64, BW62, BT65, BW65b, Con65, DH67a, Eri67, Har64b, JP69, Kru67, Ksh61, Ksh66, Low64, Mor67a, Pri64a, Ree69, Rei61, Rei62, San63, Saw66b, SS69, Seb63, Sin67, Sri60, Sub69, Sub70, Suk60, Tik64, Tik65b, Tik66, WL67, Wil69, Wil64, Amo64, Cro61, DH67b, HF69, HPW61, Owe65, SZ60]. Non-Central [BDO60, JP69, Pri64a, San63, Seb63, Tik65b, Tik66, Amo64, HPW61, Owe65, SZ60]. non-existence [HF69]. Non-Homogeneous [Con65, Wil69, Wil64]. Non-Independent [SS69]. Non-Informative [Eri67]. Non-Interactive [AS66]. Non-Isotropic [Ksh61]. Non-Linear [DH67a, Har64b, Rei61, Rei62, DH67b]. Non-Negative [Mor67a]. Non-Normal [Sin67, Sri60, Sub69, Sub70, Tik64]. Non-Normality [BN64, BW62, BW65b]. Non-Null [Ksh66, Cro61]. Non-Orthogonal [Low64, Ree69]. Non-Parametric [Kru67, Saw66b, Suk60]. Non-Sequential [WL67]. Non-Stationary [BT65]. Noncentral [Joh66, JP70]. Nonparametric [Ber69, Dav64b, Dav64e, Noe66, Sch69a]. Normal [AH65, Ait66b, AH68, AC64, Ano60i, Ano60j, Ano60h, BGKO66, BGKO71, Bre63, Cam68, Chi67, DM61, Dav63f, Day69b, Dos62, DG67, Fai68, GPS64, GP65, HJ63, HT65, Har61a, Har61b, HM66, HM67, HM69, Hin69b, Hin70, Imh61, Imh62, KOR64, Kot66, KT69, LL60b, LEH60, LEH61, Mar63, McF60, Nab61, PS64, Pea65b, RGL68, Rub56, Saw61a, Saw61f, Saw62b, SC66, SZ60, Sin67, SH68, Sno62, Sno63, Son61, Sri60, SO62, Ste62a, Sto68, Sub69, Sub70, Swa63, TL67, TF69, Tik64, Tik67, Tur60, AH66, LB65, Rub65, Saw60, Zin61]. Normality [BN64, BW62, BW65b, KT69, PP64, SW65]. Normalizing [Wis60]. Norman [Bar65d]. Nostrand [Bar63e]. Note [Ano61c, Bar67a, BG60, BT64b, Bri62, Cun69, Dow67, Gow61, GR69, Han68, Hit62, Joh65a, JW65, Kal62b, Kal64, Ksh60, Ksh62, Les60, Lit68, Lom66, Mor63a, Mos67b, Pol68, Rad67, Rao62, Rao63, Rao65b, Rao69, Seb65, Sen67, Sha66, SH68, Son61, SO62, Ste68a, Sub65, Tal69, TL67, Tik66, Vid64, You67a, You68, Zin62, GS65, Pat60, Rao68b, Saw60]. Notes [Saw64, Pea68c]. Null [Cox66a, Cox67c, Ksh66, Sto67, Cro61]. Number [Alw62, Cox60, Saw60]. Numbers [BDF60, Clu63, She68]. Numerical [Bur60b, Ewe63, HSD66, Spr69, Bur60c, Ano63f, Mil61]. Numeriche [Dav61e]. Nutritive [FMB68].

O [Ano63e, Dav60c, Dav63c]. Oakford [Ano64j]. Observations [Big61, Eul61, Kal64, Lar69, Mit68, MR68, Oka63, Pha65, Spr65, Swa62, Swa63, Wal61a, Wal61c, Ber61, BC61, Put64, Dav64g]. Obtaining [Joh62d]. Obtuse [Lan69]. Occupancy [BD68, BC60, Nic61, You68]. Off [Pea65b]. Offspring [Pol66a]. Oligopoly [Arc61]. Oliver [Dav65b]. Olkin [Buc62a]. One [Ati64, Ben68, HT65, Har69b, Har70, HS69, Joh64, Kud63, Mar66, Mar63, MT67, Saw66b, Zin62, DG66, MT72]. One- [MT67, MT72]. One-Parameter [Mar66]. One-Sided [Ben68, Joh64, Kud63]. One-Way [Ati64, Har69b, Har70, Zin62, DG66]. Onset [Bat63]. Operations [Bar60b, Mor62]. Operators [Bar63e]. Optimal [Col63, Kin64, LS68]. Optimality [RK69a]. Optimization [Spr69]. Optimum [Zin62]. Order [BD63, Bri69, Dav64c, Ghi68, Har61a, Har61b, Has69a, Hex63, LL60b, RT69, Sha69, Sil64, Sil69, Sub69, Sub70, Thi67, WGH62, WGH63a, WGH63b, You67b, BD65a, DL66, Saw60]. Ordered [Bar61a, GPS64, GP65, SC66]. Ordinary [Dav61j]. Ordinate [KT69]. Organism [RA68]. Organisms [FMB68]. Orientation [Wat66]. Orthant [Ste62a]. Orthogonal [Gon61a, Gon61b, Low64, Ree69, vHG66, HF69]. Oscar [Smi62a]. Osculatory [Joh62f, Joh62e, Joh62c]. Oskar [Arc61]. Other [Ano60n, Ano60o, Ano61f, Ano61g, Ano62e, Ano62f, Ano63i, Ano63n, Ano63o, Ano64r, Ano64s, Ano65g, Ano66h, HG68, Smi60]. Otto [Arc62]. Outer [Pra66]. Outlier [BT68]. Outliers [DP65, QD61, TW63]. outlying [Zin61]. Owen [Ano60i, Ano60j, Ano63j, Dav63d, Joh64, JW65].

P [Ano64n, Bar61c, Bar66g, Cha68a, Dav60c, Dav62e, Dav63a, Dav64d, Haw64, Joh62e, Joh66, Ken66a, Mil61, Sok60]. P.D.F. [You67b]. Packed [Hol65a]. Pair [Bur60a]. Paired [Alw62, Bra54, Bra64, DB69, DB70, RT66, Sla61, SD61a, SD61b, dC69]. Paired-Comparison [SD61a, SD61b]. Pairs [Cha68b, Kot65, Kri67]. Palin [Pea62]. Paper [Sha66]. Paradox [SS65b]. Parallel [Wil60]. Parameter [BMS67, BC60, BH60b, BH61b, DH66, DH67a, Hoe68, Mar66, Pin62, Ver61, DH67b]. Parameters [AD68, AC64, Bat63, BD65b, Cha61, Cha69, Dos62, Fra64, GG66, HT65, HM66, HM67, HM69, Har64b, Has69a, Has69b, HG67, Hit62, Hit66, Mor65, Ohl64, Rao65a, Saw61a, Saw61f, ST66, Swa62, Swa63, Ver61, Wal61b, Wal62, WGH62, WGH63a, WGH63b, Jol63]. Parametric [Kru67, Saw66b, Suk60]. Part [KLTW65]. Partial [Gan65, HK63, Wat61a, Hof63]. Partially [Bar63c, HG61]. Particular [Wil64]. Parzen [Bar62e]. Passmore [Cor64a]. Paths [Ken66a]. Patnaik [GCP66]. Patterns [BDF63]. Paul [Ken66a, Saw61c]. Pearce [Cur66]. Pearson [Ano64i, Hoa68, JNAP63, Kam66b, Pea68c, SC65, Wel62]. Peggy [Joh62d]. Percentage [Amo63, Ano60f, Bar62a, BN68, Cha67a, HT68, Har69a, Har63c, Har64a, IMH68, JNAP63, JP69, JP70, Pea64, PT65, Pea65b, Pea68b, PB61, Pil67, SM68, dVW66, Har65, Har79, JNAP65, Ste66, Bar66d]. Performance [DP65, Far60, GT69, Mar69a]. Periodic [JB67]. Periodogram [Dur69]. Peripheral [BLN62]. Permutation [Arn64, Dan62b, Par62, CB66]. Permutations [Ano64j]. Persistence [BDF62, BD66, Goo64]. Persons [Hai65b]. Perturbation [Bai68a]. Peter [Joh62b]. Petersen [Smi60]. Pfanzagl [Joh60c, Joh60b, Joh62a]. Phages [Gan62]. Phase [DG68a, DG68b, Goo67, Hin69a]. Phase-Dependent [Goo67]. Phenomenes [Mil63]. Photoelectric [Mor68]. Physical [Ano64n]. Physics [Bro62, Dav63e, Ano60c]. Pirie [Cor64a]. Places [Ano60g]. Plackett [Saw61e]. Plan [Rei60a]. Planimeter [Mor68]. Plans [Col63, MSA66, Joh64]. Plant [McG67]. Plot [Big61]. Plots [Big61, Ree69]. Plotting [WG68]. Point [Bar64b, Cro64, Gow68, Hol65a, RW64, RK69b, Bar65b]. Points [Amo63, Ano60f, Bar62a, BDF63, Bar66d, BN68, Cha67a, Efr65, Fai64, HT68, Har69a, Har63c, Har64a, IMH68, JNAP63, JP69, JP70, Nau65, Nic61, Pea63a, Pea64, PT65, Pea65b, Pea68b, PB61, Pil67, Ste65b, SM68, dVW66, Har65, Har79, JNAP65, Ste65a, Ste66]. Poisson [Pea67a, BH61b, Gar64, Gon66, HG67, Hol64, Lac69, Lew65, Mar69b, MY61, PW66b, Rob69, Sub65, Wil64]. Polar [Ano60d]. Pollen [Mos63]. Polykays [BDF60]. Polynomial [Cro64, Dow66b, Spr65, McG60]. Polynomials [Gon61a, Gon61b, vHG66]. Pooled [Oka63]. Population [Bai68a, BMS67, BGL60, BD61, Bar62f, CJ64, Cor64a, Cra62, Dav64i, Dem67, Goo68, GS62a, Hai65b, Ken66a, Pol66b, Saw61f, Sil69, Sno63, Jol63]. Populations [Bai68b, Bar69, BG62, Dos62, Fai68, HM66, HM67, HM69, HS69, Man69b, Rao68a, Sin67, Sri60, Swa63, Wil65b, Zin61]. Populationsgenetik [Kal61b]. Positive [Gon61a, Gon61b]. posterior [Rut67, TT65]. Power [BH61a, Har63c, HT62, HK68, KT69, Pat62, PJ67, PJ68, PB64, Suk60, Whi62, vHG66, BH60a, CB66]. Power-Law [HT62, Whi62]. Powerful [McN68]. Powers [Ito62, Ito63]. pp [Bar63e]. Prabhu [Loy66]. Praktischen [Joh62b]. Precision [Rao67]. Predator [LG60]. Predator-Prey [LG60]. Prediction [AS65, LB63, Bur62]. Predictive [GD68]. Preemptive [Hea61]. Preliminary [GM62]. Presence [MG69b, Spr69]. Prey [LG60]. Primer [Hea64a]. Princeton [Bar63e]. Principal [Ksh61, Ksh66]. Principles [Wal64, Dav61d]. Prior [Cam68, DH67a, SS65b, TZ64, DH67b, RK67]. Priority [Hea61, Kam67]. Probabilistic [Ker64]. Probabilities [All66, All68, Bar61d, Goo67, KM66, Mal68, Mil69, OS65, Pat64, Pol68, Sam67, Sev69, Ste62a, Sto67, Sto68, Wis63a, Wis63b]. Probability [Ano60i, Ano60j, Ano63j, Ano63g, Ano66g, Ano67f, Ano67g, Ano68g, Ano68h, Ano68i, Ano69h, Ano69i, Ano69j, Bar62e, Bar66c, Ben63, Ben65, Ben68, Bro64, Buc62a, DR63, Dav60b, Dav62b, Dav63e, Dav64i, Dem67, Ela59, Fra66, Gil65, Has67, Ken60b, Ken61a, Ken61b, Ken63b, Lan69, Loy67a, Nic61, Pea63b, Phi66, Pin62, Rab69, She68, SW60, Tho69, VW60, WD67, WG68, vHG66, Joh61, Ken68b, Ken68a, Pea65a, Pea67b, Pea68c, Sea67, SZ60, SC65, Bar63b, Dav64d, Lin63, Loy67b, Bar64a, Dav60d, Ken63b, Ken64]. Probability. [Saw61d]. probable [Ber61]. Problem [BD66, BC60, Bro67, DR63, DS68, Kal62b, LL60a, Mar68c, MM64, RS68, Tal69, You68, DS76, FKM67, Rob67, Yao65]. Problems [AS65, Ati62, BW65a, BT68, Dav62d, EEG67, Kru67, LS68, Mor61, Pea63b, PR64, Rao66, Saw66a, Wat66, LB65]. Procedure [BM65, Fai68, WL67]. Procedures [Dav61d, DB69, DB70, Gab68, Rao68a, SA62b, SA69]. Proceedings [Dav63e, Dav62d, Dav63e]. Process [Dav64i, Ewe63, Gan62, Gil60, Kas60, Ker64, KM66, Lom66, Mar67a, Pri63, Pur66, Wal61a, Wal61c, Whi64a, Wil65a, Wil69, Wil64, Wis62]. Processes [Bai68b, Bar64b, Bar65d, Bar65c, Bar66g, Blo60a, Blo60b, Buc62d, Cox67a, Dan62b, EK60, Goo67, Hai65a, Jag67, Kal61a, Kal62a, Kal62c, Ken60a, Ken66a, KG67, Lew65, Par62, Pol68, Pri64b, Wal62, Wau68, Bar65b, Bar62b, Bar65e, Ken66b, Loy66]. Processus [Haw64, Ken66a]. Product [Joh69, Kot65, KZ63, Pol66b, ST66, Ano60f]. Products [Ait63, GO66, Mil65b, Stu62]. Programmes [Har63a]. Programming [Buc65, dC69, Bea61, Hal63, Hea64a]. Programs [Joh66]. Project [Rei60a]. Projective [Arc62]. Prone [Hai65b]. Proneness [Bha67]. Pronterari [Dav61e]. Properties [BDO60, BG62, Bri69, Bro61, Gow61, Gow66b, Joh60a, LG60, Man69b, Par62, Wil62b, KK65, MA65]. Property [Ati61, Bri66, Hea64b, Kam65, Kam66a, Kam66b]. Proportions [Gar69, Mos62]. Publications [Ano66g, Ano67f, Ano67g, Ano68g, Ano68h, Ano68i, Ano69h, Ano69i, Ano69j].

Quadratic [BO59, BO61, Coo63, Goo69, Goo70, Har69b, Har70, Imh61, Imh62, Kha78, LL60a, Sha68, Tur60]. Quadratically [Ati62]. Quadrivariate [McF60, Son61]. Qualitative [Bar62c]. Quality [Bee63, Cam68, Moo61, Wet60]. Quantal [ASB60, AS64, Cox66b, GZ67, Lit68, WCV66]. Quante [Joh62b]. Quantiles [Has69b, Rut67]. Quantitative [Smi62b]. Quantity [Bee63]. Quasi [SS65b]. Que [Hol61b]. Quenouille [Dav60e, Rao65b]. Queue [Bur60b, EF62, Bur60c, Kin62]. Queueing [Fin61, Hai60, Haw65b, Hea61, Mil63, Fin60a, GS65, Cox61]. Queues [Mil63, Wil60, Bar65c, Loy66]. Quick [McN68, Mea66]. Quota [You61]. Qureishi [Yat62].

R [Ano64c, Ano64j, Bar60c, Bar63b, Buc62c, Buc64a, Buc65, Cha68a, Cor64a, Dav60c, Dav61d, Dav61b, Dav64a, Dav65b, Hea64a, Kes61, Mal61b, Men65, Mil63, Mor62, Pea66c, Saw61e, Smi60, Smi63, Sok60, Wat61a]. R. [Pea68c]. R.S.M. [Ano66d, Bar66e]. Radhakrishna [Ait66a]. Raiffa [Saw61b]. Rainville [Ano64l, Dav61g]. Ramakrishna [Ano64h]. Random [Ait63, Ano64j, Bar63d, BDF63, Bat63, Ben63, Ben65, Ben68, Boh66, Bri65, CO65, Cox60, Dav63f, DF64, Efr65, Fai64, Gre67, GPS64, GP65, Hin69b, Hin70, Hol65a, Kin61, Kot65, Kri67, Lan69, Mal61c, MD62, Mos67b, Nau65, Pea66c, Pol68, RG66, Rob68, Spr69, Ste67a, Ste68b, Ste63b, Sti69, Stu62, TR66, TT65, TT66, TD68, You61, LB65, Mil65a, Has67, Ano63g, Bar62b, Bar63a]. Random-Effect [TT65, TT66]. Randomization [CB63, KD69, Sto69a]. randomized [CB66]. Randomness [Hol65b, Pea63a, Ste66]. Range [BN68, BGKO66, BGKO71, Gho63, GCP66, GPS64, LB66, PS64, Sin67]. Ranges [Har63c, KLTW66, Mar67b, PB61, SH68]. Rank [AH68, Bra54, Bra64, FP61, GW61, MT67, MT72, TW63, Hof63, PS62a, AH65]. Ranking [Fai68, Cro61, FKM67, dC69]. Rankings [Bur60a, Lin60, RT66]. Ranks [Ait66b, Ken62, AH66]. Rao [Ait66a]. Rapid [Dav60e]. Ratcliffe [Dav63b]. Rate [Bri62]. Ratio [Cha67a, Gho63, Har63c, Hin69b, Hin70, Joh69, Pea64, PS64, Pea67a, PB61, PB64, Rao67, Rub64, Saw64, Sch66, Sch67, Tik64, VW60, SC65]. Ratios [CB63, GPS64, McG67, Rao65b, RW66]. Ray [Saw62a]. Real [Ric67]. Realization [Con65]. Recapture [Dar61, Jol65, Seb62, Seb65, Jol63]. Received [Ano60n, Ano60o, Ano61f, Ano61g, Ano62e, Ano62f, Ano63n, Ano63o, Ano64r, Ano64s, Ano65c, Ano65g, Ano66h]. Reconsidered [Sok60]. Rectangular [GO66, Roa63, Ano60d]. Rectangular-Polar [Ano60d]. Recurrence [You67b]. Rediscovered [Dav68]. Reduction [Chi67, RW66]. Reference [Cox67a, Saw62b, TD68, TF69]. reflexions [Pea67b]. Regions [Har64b, Pra66, DL65, DL66]. Regression [BW62, BW65b, CK68, CK70, Cro64, Cur60, Dav60a, Dav69, Dur69, Gar64, Gea66, Han63, Han67, Har64b, Hin69a, LL60a, LB63, LS68, LM63, LMH67, LMH71, LM64, MK69, Mit68, MR68, Moh68, Pea66c, Saw66a, Sri60, TZ64, Tia66, Ver61, McG60, Buc61, Saw61e]. Regressions [Hus69]. Regular [BG60]. Reichmann [Bar63g]. Related [Ano60h, Hex63, Rao66, Saw66a]. Relatedness [All67]. Relating [Dan62a]. Relation [Han67]. Relations [Sil64, You67b]. Relationship [Wis62, Dav62f]. Relationships [Bar67a, Dow66a, DG67]. Relative [Ano64h, Gea66, Rao69, Stu67, Rao68b]. Reliability [Buc62d, Buc64b]. Remark [Kha78, Lan66]. Remarks [Bar67b, MR68, Sha68, SS65a]. Removals [Gan62]. Removing [Hai65b]. Renewal [BD66, Hai65a, KG67, Lea63, Lea64, Lom66, Buc64a]. Renewals [Cox60]. Replacement [Sam62, Sam67, Sto68]. Replicate [Pea68a].

Replication [Kin64]. Representation [Bri65]. Representations [Amo64]. Required [DR63]. Research [Ano60c, Ano64i, Mor62, Rei60a, Bar60b, Gru61]. Research/Management [Mor62]. Reseaux [Har63a]. Residual [Ati62, Bar67a, Ber68, Har67b]. Residuals [Atk69b, Dur69, Wal62]. Response [ASB60, AS64, CC67, Lit68, Wil65a, Wis60, WCV66]. Responses [BD65b]. Restricted [Thi67]. Result [Bha67, DG68b, Sha68]. Results [BGL60, Bur60b, DG68a, EK60, Ewe63, Gho64a, Gle68, Jac68, Lew65, Pea66a, Whi64b, Bur60c, CB66]. Review [Ait66a, Ano60i, Ano60j, Ano60h, Ano60g, Ano60f, Ano60d, Ano60c, Ano60e, Ano63j, Ano63i, Ano63d, Ano63g, Ano63e, Ano63f, Ano63c, Ano63k, Ano63h, Ano64n, Ano64m, Ano64l, Ano64k, Ano64i, Ano64d, Ano64c, Ano64f, Ano64j, Ano64g, Ano64e, Ano64h, Ano66d, Ano69d, Arc62, Arc64, Arc61, Bar60b, Bar64a, Bar61b, Bar62b, Bar63a, Bar66b, Bar60c, Bar61c, Bar62e, Bar62f, Bar62c, Bar62d, Bar63d, Bar63b, Bar63c, Bar63g, Bar63f, Bar63e, Bar65d, Bar65e, Bar65c, Bar66c, Bar66g, Bar66f, Bar66d, Bar66e, Bea61, Bee63, Bro62, Buc60, Buc61, Buc62a, Buc62d, Buc62b, Buc62c, Buc64a, Buc64b, Buc65, Bur62, Cha68a, Clu63, Cor62b, Cor62a, Cor64a, Cox61, Cox67a, Cox67b, Cro60, Cur66, Dav60i, Dav60f, Dav60h, Dav60j, Dav60d, Dav60a, Dav60c]. Review [Dav60b, Dav60e, Dav60g, Dav61d, Dav61a, Dav61b, Dav61g, Dav61j, Dav61e, Dav61f, Dav61i, Dav61h, Dav61c, Dav62f, Dav62d, Dav62c, Dav62e, Dav62b, Dav62a, Dav63e, Dav63b, Dav63c, Dav63d, Dav63f, Dav63a, Dav64a, Dav64b, Dav64e, Dav64c, Dav64h, Dav64d, Dav64f, Dav64g, Dav65a, Dav65b, Dra64, Edw61, Ela59, Fel63, Fin62, Gob67, Gru61, Hal63, Har63b, Har63a, Haw64, Haw65a, Hea62, Hea64a, Hol61a, Hol61b, Jen64, Joh60c, Joh60b, Joh62f, Joh62e, Joh62c, Joh62d, Joh62a, Joh62b, Joh64, Joh66, Kal61b, Kal61c, Ken64, Ken66a, Ken66b, Kes61, Lin63, Lin69, Llo60, Loy66, Loy67a, Loy67b, Mal61a, Mal61b, Men65, Mil61, Mil63, Moo61, Moo63, Moo66, Mor62, Noe66, Pea64, Pea66b, Pea66c, Pea67a, Rad66, Rei60a, Rei60b, Ryb66, Saw61b, Saw61e, Saw61d, Saw61c, Saw62a, Smi60, Smi62b]. Review [Smi62a, Smi63, Sok60, Sto62, Stu64, Tay63, Tay64, Wal63, Wal64, Wat61a, Wat66, Wel62, Wet60, Wet62, Yat62]. Reviews [Ano67c]. Rho [DM61]. Rhythms [Wat66]. Richard [Dav64b]. Riordan [Jen64]. Ritchie [Cor64a]. Road [Yeo64]. Roberson [Joh62d]. Robert [Saw61b, BD66]. Robin [Dan69, Dan70]. Robinson [Bar63a, Fin62]. Robustness [Ati62, Ati64, BT62, BW62, BT63, BT64b, BW65b, IS64]. Roentgenologic [Rei60b]. Roessler [Dav62b]. Role [Sto69a, Sto69b]. Rolf [Joh66]. Romig [Ano60e]. Ronald [Ken63a]. Rooms [Ghi68]. Root [Gow66b, Pil65b, Pil67]. Roots [AG63, Bar66a, Bar66d, CO65, HT68, Mot62, Pil64, Pil65a, TF69, AG65]. Rosenblat [Bar62b]. Rosenblatt [JW65]. Rosenhead [Dav63a]. Rosenthal [Mor62]. Rotatable [BD63, BD65a]. rotations [Mil65a]. Round [Dan69, Dan70]. Round-Robin [Dan69, Dan70]. Rounding [Pea65b]. Routines [Yat62]. Row [Lar69]. Roy [Kal61b]. Rudert [Pea67a]. Rules [Eri67]. run [BC61]. Runs [Kri67].

S [Ano64i, Bar62f, Buc60, Cox67a, Cur66, Dav60h, Dav62c, Dav65a, Hal63, Lin69, Loy67b, Moo61, Moo66, Mor62, Pea68c, Saw62a, Smi60, Smi62b, Wel62, Yat62]. Saaty [Mil63]. Sais [Hol61b]. Sais-Je [Hol61b]. Sakti [Joh66]. Salem [Pea67a]. Salzer [Joh62e, Joh62c, Joh62d, Joh62f]. Same [ASB60, DP60, PS64]. Sampford [Dav64a]. Sample [Bar62a, Ber69, Bha60, Cro66, Cro67, Dar61, Dem67, Geh65a, GT69, HR68, Has69b, IMH68, Ken66a, Kru67, MGR60, MGR61, PS64, Rao60, RGL68, Roa63, Saw61a, Saw61f, Seb62, Sil69, SH68, Ste65b, Ste69b, Sto67, Sto68, Suk60, Tik67, Ury66, Wal61a, Wal61c, Wal61b, Wal62, WW64, Wil63, Haj61, Maa66, MG69a, Mur67]. Sampled [Mos63]. Samples [BD61, BGKO66, BGKO71, DM61, Dos62, Geh65b, GT69, HJ63, HM66, HM67, HM69, Mar67b, McN66, Pea65b, Saw66b, Sil64, Sin67, Sno62, Sno63, Sri60, Swa62, Swa63, SW65, Ano64o, Bar63c]. Sampling [BD61, BG62, Cam68, CJ64, Dar61, Dav64a, DG68a, DG68b, EK60, Han62, Joh64, Low64, Mor63a, Pat64, Rao68a, Sam62, Sam67, Sed66, SS65a, SJG68, Sto68, Sto69a, Tik64, You61, LB65, Ano60e, Dra64, Ano60e, Dav61f]. Samuel [Saw61d]. sans [Haw64]. Sarhan [Dav64c]. Särndal [Ano64o]. Satterthwaite [HK68]. Satya [Hol61a]. Saul [Buc65]. Savage [Ano64k, Dav64b]. Saxena [Llo60]. Scale [AD68, BMS67, Fra64, Saw61a, Ver61, WGH63b]. Schedule [Sla61]. Schemes [Cam68, Cha61, Cha64]. Schlaifer [Saw61b]. Schmetterer [Ano63k]. Schreider [Arc62, Hol61b]. Schwartz [Ano64g]. Science [Hea62, Mor62, Wet62]. Sciences [Bar65d]. Scientific [Bee63, Dra64]. Score [Bur60a]. Scores [Dan69, Dan70, Hea64b, Rub56, SC66, Sto68, Rub65]. Scott [Tay63]. Screening [Col63, Kin64, ZF69]. Seckler [Lin63]. Second [BD63, Bri69, Hex63, Joh62d, RT69, BD65a, DL66, MG69a]. Second-Order [Hex63, RT69, DL66]. Secondarily [Bro67]. Selected [Tra69]. Selecting [CJ64, GS62a]. Selection [DS68, Sam67, Sti69, Wil62b, Wil63, DS76]. Sensitivities [SS69]. Separate [Jac68, Wal67, Wal70, WGH63b]. Separation [Kas60]. Sequence [Kri67, Cro61]. Sequences [TR66]. Sequential [All66, All68, BD68, Bha60, BN64, Boh66, Day69a, Ewe61, Gar63, Gho63, Gho64b, Hal62, Hoe68, Kin64, LB65, Man69a, Pha65, Pra66, SA62b, SA69, Sed66, VW60, Wet61, WCV66, Whi64b, WL67, You68, Haj61, MSA66, SA62a, Sis64, Dav62e]. Serial [Cam68, Cox66a, Cox67c, Dur69, Mor67b, Mor63b, Pat68, Whi61, McG60]. Serially [Lar69]. Series [ASB60, Bar67b, BT65, Cha69, Cla66, Gea66, Gow61, Ham68, Han63, JB67, Ken66b, Lea63, McG62, MB65, Mor63b, Ord68, Pat62, Rao69, Tik65b, Wal60, Wal67, Wal70, Wis68, vHG66, Rao68b, Saw60, Wis90]. Server [EF62]. Servers [Ghi68]. Service [Jen64, Mor62]. Set [Bar62a, Bar61d, Efr65, Gle68, pH68]. Sets [Pre66, Pre69]. Seven [Pil64].

Several [BD65b, DP65, GS62b, Oka63, Saw66a, WD67, Wan67, Ber61]. Sexes [Goo68]. Shall [Ken60b]. Shape [Pin62, WGH63b]. Sheppard [Wal64]. Short [KLTW65]. Short-Cut [KLTW65]. Shortest [BH60b, BH61b]. Should [Dav61i]. Shoultz [Joh62e]. Shows [Dav65b]. Shubik [Arc61]. Sided [Ben68, Joh64, Kud63]. sidelight [DG66]. Sighting [Cor64b]. Sign [BJ69, Ury66]. Significance [GCP66, KD69, Rad64, SD61b, Ste65b, Sto69b, Wat60, Ste65a]. Significant [SD61a]. Silvio [Dav61e]. Similar [SS69]. Simple [Ajn68, Bai68a, BDF62, Cox66b, Dav64i, HS69, Ker64, Wil65b, Bai63, CB66, Pat60]. Simplified [HG67]. Simulated [Hof63]. simulation [Mur67]. Simultaneous [Gab68, Gho64b, Lac69, LM63, LMH67, LMH71, Mos67a]. Single [Ano60e, EF62, Ksh61, KLTW65, Seb62, Spr69, Wil63]. Single-Server [EF62]. Singly [AH65, AH68, Dos62, Geh65b, HM66, HM67, HM69, RGL68, AH66]. Singly-Censored [Geh65b]. Situations [DH66, DH67a, WL67, DH67b]. Sixteen [Ano60g]. Size [Bha60, DS68, Dow67, Dow68, HK68, Man69b, Tra69, DS76]. Sizes [Sam62, Sil64]. Skew [Hoa68]. Small [GT69, HS69, MGR60, MGR61, Pin62, Sla66, Hof63]. Smaller [Cra62]. Smallest [GS62a, GS62b]. Smooth [Mar66, Mar67d, Mar68b, Mar68c]. Smoothing [LK69]. Sneddon [Dav62a]. Social [Tay63]. Society [Kal61c]. Solomon [Bur62, Dav62c]. Solution [BDM60, BDM63, Dav61j, Kal61a, Kal62a, Kal62c, Ker64, Sis65, Bai63, Rob67, Yao65]. Some [AS65, Ait69, Ait71, AG63, AG65, Bar60a, Bar65a, BB66, BW65a, BGL60, Bar67b, BDO60, Blo63, BT68, Bro61, Bur60b, Bur60c, CB66, DR63, Day69a, Dow66a, DG67, DG68b, Dur61, Dur66, Fai68, Far60, GT69, Gho64a, Gle68, Gow61, Gow66b, Hal62, HG67, HSD66, Jac68, Joh60a, JY60, KK65, KD69, Kin62, Kru67, Ksh60, Les60, Les62, Lew65, Man69a, Man69b, Mar67c, Mar68a, Mar69a, Mil65b, Mor61, Mur67, Par62, Pea63b, Pea69, Phi66, Pol66a, Pra66, Pre66, Pre69, Pri64a, QD61, Ram60, Ram61, Ree69, RK69b, Saw64, Sev69, Sha68, Sil64, SS65a, Sok60, Ste67a, Ste68b, Ste69a, Sto69b, Suk60, Swi68, Wal60, Wal67, Wal70, Whi64b, Wil62b, You67b, Pea66a, Pea65a, Pea67b, Pea68c, Ano64h]. Somerville [Fai68]. Southern [Ano64n]. Spaced [Spr65]. Spatially [Bai68b]. Spearman [Bro61, DM61]. Special [Cox67a, Rei60a, Saw62b, TD68, TF69, Owe65, Dav61g]. Species [Hol69, LG60, Les62, McG67]. Specification [Bro64]. Specified [LB63, BC61]. Spectra [Pri64b]. Spectral [Bar64b, Bri69, Cla66, Dan62a, Whi63, Bar65b]. Spectrum [Pri63]. Speed [Mor68]. Sperner [Arc62]. Sphere [Bre63, Gil65, MD62, Sel64, Ste67b, Wat60, Wat65]. spherical’ [DL66]. Spider [BD66]. Split [Big61, Ree69]. Split-Plot [Big61]. Splitplot [CB63]. Spread [Mor65, TL67]. Square [Har67a, Har64a, Oka63, Pea64, Pea67a, RG66, San63, Sla66, Tik65a, Har65, Har79, SZ60]. Squared [Goo69, Goo70, Kha78, Seb63]. Squares [Ano63h, Ati62, Dur69, Har67b, HK68, Kel62, Rao65a, Rub56, SC66, Swi68, HF69, Rub65]. Stage [Hal62]. Stahl [Bar63b]. Standard [JNAP63, Mea66, PS64, PT65, Tik67, JNAP65, You62]. State [Con65]. States [Thi67]. Stationary [BT65, Bri65, McG62, MB65, Mor63b, Pri64b, Wis62]. Statistic [Kor68, Ksh66, Oka63, Sel65, Ste63a, Ste64, Ste65b, Ste69a, Sto67, Ver63, Maa66, Ste65a]. Statistical [AS65, Ait66a, Ano64k, Ano64d, Ano64c, Ano64f, BW65a, Bar61b, Bar66b, Bar60c, Bar66g, Coo63, Cro60, Dav60c, Dav60e, Dav63d, Gea66, Hea62, Joh66, Kru67, Moo61, Mor68, Mor69, Pea66b, Saw61b, Smi63, Stu64, Swa62, Tay63, Wel62, Wil62b, Yat62, Ano64i, Bar66f, Buc62d, Buc64b, Dav60f, Dav60g, Joh66, Men65, Rad66, Sto62, Wat66, Wet62]. Statistici [Dav61e]. Statistics [Ano66g, Ano67f, Ano67g, Ano68g, Ano68h, Ano68i, Ano69h, Ano69i, Ano69j, Bar66c, BO59, BO61, Cur66, Dav62f, Dav63e, Dav65b, Ela59, GS62b, Har61a, Har61b, Has67, Has69a, Ken60b, Ken61a, Ken61b, Ken63b, McN66, Noe66, PP64, Rab69, Rao60, She68, Sil64, Sil69, Ste66, Sto69a, Sub69, Sub70, Tik65a, Ver64, WGH62, WGH63a, WGH63b, Wil64, You67b, Ken68b, Ken68a, Pea65a, Pea67b, Pea68c, Saw60, Sea67, Ano63c, Bar64a, Bar63g, Buc62a, Buc62b, Cor62b, Cox67b, Dav60d, Dav60b, Dav61d, Dav61a, Dav61b, Dav62b, Dav63b, Dav63c, Dav64b, Dav64e, Dav64c, Dav64f, Hol61a, Llo60, Moo66, Saw61c, Wet60, Dav60i, Dav60h]. Statistik [Ela59, Ano60k, Bar63f, Joh60b, Joh62a, Joh62b]. Statistiques [Ano64g]. Statistische [Kal61b]. Steel [Dav61d]. Stein [Hal62]. Steinhardt [Bar62d]. Stevens [Pea66c]. Stichproben [Bar63f]. Stimuli [AS66]. Stochastic [Bai68a, Bai68b, BGL60, Bar65d, Bar65c, BLN62, Cox67a, Dow67, Gan61, Goo68, HS69, Jol65, Kas60, Ken66b, Les60, LG60, Les62, Loy66, Mot62, Pol66b, Rao65a, Sis65, Wil65b, Bai63, Bar62f, Jen64]. Stochastiques [Ken66a, Haw64]. Stopping [Eri67]. Storage [Bar61c]. Straight [Buc60]. Straight-Line [Buc60]. Strategy [Arc61]. Stratification [SS65a]. Stratified [Dar61, DG68a, DG68b]. String [Smi60]. Stringent [BN65]. Structural [Bar67a, Fra66]. Structure [Arc61, Dav60j, Fra65, Ham68, JB67, Lin69]. Stuart [Dra64, Rao69, Cox67b, Dav62f, Rao68b]. Student [GJ68, Kot66]. Studentized [Moh68]. Studies [Has67, Ken60b, Ken61a, Ken61b, Ken63b, Ken68b, Ken68a, Pea65a, Pea67b, Pea68c, Rab69, Sea67, She68, Bur62]. Study [Loy66, Ram60, Ram61, Rei60b, Tay63, Mur67, Rei60b]. Subjects [ASB60]. Subset [DS68, GS62a, Tra69, DS76]. Subset-Size [DS68, DS76]. Substitute [PB61]. Successive [EK60, Har67a]. Sufficiency [BG60]. Sufficient [Kha78, Rao60, RG66]. suggested [Sis64]. Sum [Bar66d, Ben65, Ben68, Har67b, MT67, Rub56, SC66, TW63, MT72, Rub65]. Sums [Bar63d, Ben63, GO66]. Superimposed [Wal60]. Superposition [TR66]. Supplement [Ano69d]. Supplementary [Kru67]. Supplemented [Pea60, Pea61]. Supplied [Swa62, Swa63]. Supplies [Cor64a]. Support [Arn64]. Surface [Rao60]. Surfaces [Gon66]. Surveys [HR68, Sed66, Dav61f]. Survival [Cor64b, Pol66a]. Susceptibles [Wil65b]. Swaroop [Hol61a]. Sweep [Ati68, Ati69]. Sweep-Out [Ati68, Ati69]. Symmetric [BN68, Mal61c, Pil65a, Gob67]. Symmetrical [You67a]. Symmetry [Sch69a]. Symposia [Kal61c]. Symposium [Dav62d, Dav63e]. Synthetic [HK68]. Syski [Buc62c]. System [Fin61, Ghi68, Ord67, RA68, Fin60a, GS65]. Systematic [Han62, SJG68]. Systems [Bro64, Han69a, Han69b, Haw64, Les60, Bar63a, Buc62c, Jen64].

T [Ano60i, Bar65d, Bar66d, Clu63, Dav60g, Dav61b, Joh62d, Mil63, Wal63]. Table [BDM63, BH60b, BH61b, Har69a, Har64a, JNAP63, JNAP65, SW60, dVW66, BH60a, BH61a, Har65, Har79, Ano60d, Ano63h]. Tabled [Pea68b]. Tables [Ano60e, Ano66d, Bar62a, BDM60, Bar66e, Bra54, Bra64, Cha67a, Cro68, Dav63c, HT62, IK68, Joh65b, JP69, JP70, Lac69, LEH60, LEH61, MY61, Pea64, Pea65b, TL67, Ver63, Ver64, OI61a, OI61b, Pla64, Dav63d, Dav63a, Gob67, Ano60i, Ano60h, Ano60g, Ano60f, Ano63j, Ano63i, Ano64j, Dav63f, Joh62f, Joh62e, Joh62c, Joh62d, Pea67a]. Tagging [Dar61]. Tail [OS65, Sto67, Sto68]. Takacs [Mil63]. Talbot [Ano64n]. Talmud [Rab69]. Talmudic [Has67]. Tanner [Dav65a, HB60, Hai61]. Tau [Sno62, Sno63]. Tavole [Dav61e]. Taylor [Loy67b]. Technique [Gow66a, LK69, Rad66]. Techniques [Arn64, HG67, RK69b]. Telephone [Buc62c]. Ten [Ano60g]. Tensors [Hex63]. Terminal [Hoa68]. Terms [Mar63, Bai63, Saw60]. Test [Ajn68, Ati62, Atk69a, BN64, BJ69, BW65b, DB69, DB70, Gab68, Gar63, Gar69, Geh65a, Geh65b, Hal62, Har63c, IS64, Kud63, OI61a, OI61b, PB64, PP64, Saw66a, Sel65, Sen67, Sto68, TW63, VW60, Ver63, Ver64, Wat67, WW64, BH60a, BH61a, CB66, GM62, Haj61, McG60, MSA66, Put64, SW65, Sis64, You62]. Testing [Bar62a, Bat63, GW61, Gle68, GM62, pH68, Kor68, Kor69, LB66, Mar68c, McN66, MG69b, Mor67a, Mor67b, Mos67b, PW66a, PW66b, Rao69, Sto69b, Tar79, Tar90, Thi67, Wis68, Rao68b, Ste66, Wis90, Cro60]. Tests [ANR68, Ait69, Ait71, All66, All67, All68, Ano64c, Bar61a, Ber69, BN65, Bha60, BW62, Cha64, Cro68, Cro66, Cro67, DP60, DP65, Dur61, Dur66, Dur69, FP61, Gea66, GT69, Gho63, Gho64b, GCP66, Hex63, Hit66, Hoe68, Hol65b, HK68, Ito62, Ito63, Jac68, KD69, KT69, Ksh60, Lew65, Lin60, Man69a, Mar69a, MGR60, MGR61, McN68, PS62b, PS62a, Pea63a, Pea66a, PJ67, PJ68, QD61, Rad64, Rao69, Saw64, Sch69a, Seb64, Sla66, SD61a, SD61b, Ste62b, Ste62c, Ste67b, Ste69b, Ste69c, Stu67, Suk60, Tho69, Ury66, Wal67, Wal70, Wat60, Wat61b, Wat62, Mik65, MA65, Mur67, Rao68b, SA62a, Bha60]. Tetrachoric [Ham68]. Their [Loy66, vHG66, Ken66a, Tay63]. Theodore [Bar65e]. Theorem [BT62, BT63, Cox67d, Ken61c, Tay64, TZ64]. Theoretical [BGL60, Tan62, Sri60]. Theorie [Joh60c, Mil63]. Theories [Haw65a]. Theory [Arc61, Bar64a, Bar60c, Bar61c, Bar62e, Bar63b, Bar63c, Bar65e, Bar65c, BD66, Blo63, Buc62c, Buc64b, Cox61, Cox67b, Dav62f, Dav63e, Dav64a, Dav64f, Gan65, Gho64a, Har63b, HR68, Joh60c, Ken60a, Ken63b, Kes61, Lin63, Loy67a, Mil63, Mor68, Rao65a, Wal63, Wet62, ZF69, HF69, Blo60a, Bar61b, Bar66g, Buc64a, Mal61a, Saw61b, Sto62, Dav61j]. there [BDF62]. therefrom [Ber61]. Thinking [Dav62c]. Third [Sno62]. Thom [Bro62]. Thomas [Cor64a, Ken68b]. Thompson [Joh62e]. Three [Has69a, Hoa68, Pat69, Rao68a, RS68, TB67]. Three-Component [TB67]. Three-Dimensional [RS68]. Tied [Bur60a]. Time [Bar67b, BT65, Bri65, Cla66, Gea66, Han63, JB67, Ken66b, Les60, Mor63b, Rao69, Wal60, Wal67, Wal70, Rao68b]. Time-Series [Bar67b]. Times [Bur60b, Wil65a, Bur60c]. Titchen [Mor62]. Todhunter [Ken63b]. Tolerance [Joh64, LM63]. Topographic [Whi62]. Toronto [Bar63e]. Torrie [Dav61d]. Tournament [Dan69, Dan70]. Tournaments [Gle60]. Towns [Tay63]. Tracer [Wal64]. Tracts [Pea66c, Pea66c]. Traffic [Haw65a, Haw65b, Yeo64]. Trails [Dav62e]. Transform [Ano64l]. Transformations [HT62, KT69, MY61]. Transformed [Dav60f, PS62a]. Transforms [HM60]. Transition [KM66]. Transport [Har63a]. Transportation [BLN62]. Treatments [Pre66, Pre69]. Trees [Rob68]. Trend [Ken61c, Rao69, SBJ68, Ury66, Rao68b]. Triads [Alw62]. Trials [Cha68b]. Triangle [Lan69]. Tribolium [Les62]. True [Cur60]. Truncated [AH65, Ait66b, AH68, Pat62, RGL68, Sub65, Swa63, AH66]. Truncation [RW64]. Trunks [BLN62]. Tuberculosis [Rei60b, Rei60b]. Tukey [Cox92, MY61]. Two [Bar64b, Ber69, Boh66, Bro67, Cra62, Cro68, Dar61, DG68a, DG68b, Fai64, Geh65a, GT69, Ghi68, Hal62, Har63c, Has69a, HK63, Hin69a, Hin69b, Hin70, Ito62, Ito63, JY60, KW63, KG67, Lar69, LG60, Les62, Low64, MT67, McF60, McG62, MB65, Moh68, Mor65, Mos67a, Nau65, Pil65a, PJ67, PJ68, Pre66, Pre69, Pri64b, RW66, Saw66b, Sla66, Ste65b, Sto67, Sto68, Sto69a, Suk60, TL67, TD68, WW64, Wil60, Yeo64, You62, You67a, Bar65b, Haj61, MT72, Mur67, Rob67, LB65]. Two-Dimensional [Bar64b, Pri64b, Bar65b, Rob67]. Two-Lane [Yeo64]. Two-Means [Bro67]. Two-Phase [DG68a, DG68b, Hin69a]. Two-Sample [Ber69, Dar61, Geh65a, GT69, Ste65b, Sto67, Sto68, Suk60, WW64, Haj61, Mur67]. Two-Stage [Hal62]. Two-Way [KW63, Lar69, Low64, MT67, Mos67a, MT72]. Type [BW63, GCP66, Kam66b, KG62, LG60, Mar67c, Mar68a, Mar69a, Pat63, Pea66c, Pol68, SC65, Ste68a, Ste67a, Ste68b, Hol66, Saw61a, Saw61f, Saw62b].

U [Loy66]. U.S. [Ano60g]. Ulf [Dav60d]. Ultimate [Dow67, Dow68, Wau68]. Unbalanced [Mos67b]. Unbiased [Bar61d, BH60b, BH61b, Har69b, Har70, Rao67]. Uncontrolled [Tan62, Tan67]. Uncorrelated [LL60b]. Unequal [DG66, Han68, Pat64, Rag62, Sam62, Sam67, Mur67]. Uniform [GD68, Ste66]. Uniformity [Ajn68, Ber69, BJ69, Wat67]. Univariate [DB69, DB70, Dem67, Sil69]. Unlimited [LMH67, LMH71]. Unrestricted [You61]. unstructured [Rob67]. Untransformed [HG68]. Upper [Ben65, Dem67, PB61, Pil67]. Usage [Ano64g]. Use [Atk69b, Bar63g, CJ64, CK68, CK70, Cor62b, DH67a, Gon61a, Gon61b, GCP66, HM60, Har69a, Has69a, Hoa68, LB66, Pol66b, Sti69, TZ64, DL66, DH67b]. Used [Gow66b, Kor68, Rad64]. Useful [PR64]. Uses [vHG66]. Using [Hit62, Kal66, Pea63b, PB61, Sis68, WGH62, WGH63a, WGH63b, You61]. utilizing [MG69a].

V [Ano64j, Bar63d, Bar63b, Bar66c, Dav61b, Lin63, Pea66c]. Vaclav [Bar65c]. Vajda [Hal63]. Validity [Sla66]. Value [Cur60, Gea66, GR69, Has69a, Sis68, Stu67]. Values [Ait66b, GW61, HJ63, Har61a, Har61b, KG67, Mil69, Pin62, SC66, Sin67, Ver63, Ver64, AH66, Saw60]. Variability [Spr69]. Variable [Atk69b, HT65, Ric67, RG66, Whi64a]. Variables [Ait63, Ano63g, BKM67, Ben63, Ben65, Ben68, Boh66, Bri65, Dav69, EEG67, GPS64, GP65, Hin69b, Hin70, Imh61, Imh62, Joh64, Kot65, Saw62b, Ste67a, Ste68b, Stu62, Tur60, WD67, You67b, LB65, Bar63d, Dav60f, Fin62]. Variance [Ano63h, Ati62, Bar61a, Cha67a, Cra62, Cun69, DM61, Gab68, GZ67, GS62a, HR67, Har69b, Har70, Ito62, Ito63, IS64, KW63, Lar69, Les60, LB66, LEH60, LEH61, Low64, MGR60, MGR61, Mos67a, Pea66a, PR64, Ree69, Saw64, Sch53, Sch69b, Swi68, TT65, TT66, Tik64, Wal61a, Wal61c, Wan67, Whi61, Wil62a, Wis62, DG66, MG69a, SW65]. Variance-Components [TT65]. Variance-Covariance [IS64]. Variance-Ratio [Saw64, Tik64]. Variances [AH68, Bar62a, Bar67a, BT64a, Gho63, Gla62, GM62, pH68, Low64, RGL68, DG66, Mur67, Zin61]. Variate [Ait66b, AH66]. Variate-Values [Ait66b, AH66]. Variates [Gow66a, GO66, LL60b, Mil65b, Mor67a, Mor67b, Pat69, SC66, Sha66]. Variation [Dav65a, IMH68, JW65, KOR64]. Variations [Ano63e]. Various [GZ67]. Vector [Gow66b, Gow68, Han69b, Ste67b]. Vectors [Mal61c, TF69]. Venn [Ano64m]. Verwandte [Ano63k]. Via [BT62, BT63]. Vianelli [Dav61e]. Viewpoint [Bar66c]. VII [Cox92]. Virology [Gar64]. Visits [Ghi68]. Vogler [Pea64]. Vol [Dav62f, Dav63e, Loy67a]. Vols [Ano69d]. Volume [Ano60p, Ano61h, Ano62g, Ano63p, Ano64t, Ano65h, Ano66i, Ano67h, Ano68j, Ano69k, Dav62d, Cox92].

W [Ano64c, Bar60c, Bar63b, Bar63g, Bea61, Cor64a, Cox92, Dav60c, Dav60b, Dav61c, Dav62c, Kal61c, Men65, Pea66c, Pea67a, Smi60, Wal64]. W. [Pea68c]. Wahrscheinlichkeitsrechnung [Ano60k, Ela59]. Wahrscheinlichkeitstheorie [Ano63k]. Waiting [Bur60b, Ghi68, Bur60c]. Walk [Ste63b]. Walks [Kin61]. Walsh [Dav64e, Noe66]. Walter [Bea61]. Warning [Pag62]. Warranty [MS69]. Watson [Maa66, Pol68, Wat61a]. Wavelets [Bar63a]. Way [Ati64, Har69b, Har70, KW63, Lar69, Low64, MT67, Mos67a, Zin62, DG66, MT72]. Weatherburn [Ano63c]. wedge [MSA66]. Weibull [Lom66, MS69]. Weiner [Dav65a]. Weiss [Sto62]. Where [Roa63, Bar66a, Big61, Cam68, Con65, Ken60b]. Which [LL60b, Saw66b, WL67, DL65, Ord68, Whi64a]. White [Rei60b]. Whole [Big61]. Whole-Plots [Big61]. Whose [Hoa68, Kot65]. Wilcoxon [Ver64, Geh65a, Geh65b, Sto67, Ver63]. Wilks [PG70, Sch67, Mik65, Moo66, PG69, Sch66]. William [Loy67a, Saw62a, Pea62]. Williams [Buc61]. Wishart [HT68, Kab64]. Within [Lar69]. Without [GT69, Sam62, Sam67, Sto68]. Wold [Ken66b]. Wolf [Tay63]. Wolfgang [Wat61a]. Work [Sok60]. works [Cox92]. wrong [DL66].

XI [Ken61a]. XII [Ken61b]. XIII [Ken63b]. XIV [Pea65a]. XIX [Ken68a]. XV [Sea67]. XVI [Has67]. XVII [Pea67b]. XVIII [Ken68b]. XX [Pea68c]. XXI [She68]. XXII [Rab69]. XXVII [Pea66c].

Ya [Bar63b]. Yates [Dav61f]. York [Bar63e, Rei60a]. Young [Ken68b]. Ysidro [Ken68a].

Z [Ano63f]. Zeitschrift [Ano63k]. Zelen [Buc64b]. Zero [BP63, GD68]. Zero-Mean [GD68].

References

Ali:1964:GEP

[AC64] Mir M. Ali and L. K. Chan. On Gupta’s estimates of the parameters of the normal distribution. Biometrika, 51(3/4):498–501, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

          tronic). URL http://www.jstor.org/stable/2334159.

Aitchison:1968:LLI

[AD68] J. Aitchison and I. R. Dunsmore. Linear-loss interval estimation of location and scale parameters. Biometrika, 55(1):141–148, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

          tronic). URL http://www.jstor.org/stable/2334458.

Anderson:1963:SIC

[AG63] T. W. Anderson and S. Das Gupta. Some inequalities on characteristic roots of matrices. Biometrika, 50(3/4):522–524, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

          tronic). URL http://www.jstor.org/stable/2333921. See corri-

genda [AG65].

Anderson:1965:CSI

[AG65] T. W. Anderson and S. Das Gupta. Corrigenda: “Some inequalities on characteristic roots of matrices”. Biometrika, 52(3/4):669, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333729.

See [AG63].

Aitkin:1965:CST

[AH65] M. A. Aitkin and M. W. Hume. Correlation in a singly truncated bivariate normal distribution II. Rank correlation. Biometrika, 52 (3/4):639–643, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333717.

Aitkin:1966:CST

[AH66] M. A. Aitkin and M. W. Hume. Correlation in a singly truncated bivariate normal distribution. III. Correlation between ranks and variate-values. Biometrika, 53(1/2):278–281, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334083.

Aitkin:1968:CST

[AH68] M. A. Aitkin and M. W. Hume. Correlation in a singly truncated bivariate normal distribution IV. Empirical variances of rank correlation coefficients. Biometrika, 55(2):437–438, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334892.

Aitchison:1963:IDI

[Ait63] J. Aitchison. Inverse distributions and independent gammadistributed products of random variables. Biometrika, 50(3/4): 505–508, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333916.

Aitchison:1966:BRB

[Ait66a] J. Aitchison. Book review: Linear Statistical Inference and its Applications, by C. Radhakrishna Rao. Biometrika, 53(3/4):634–635, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333680.

Aitkin:1966:CBV

[Ait66b] M. A. Aitkin. The correlation between variate-values and ranks in a doubly truncated normal distribution. Biometrika, 53(1/2):281–282, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334084.

Aitkin:1969:STC

[Ait69] Murray A. Aitkin. Some tests for correlation matrices. Biometrika, 56(2):443–446, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2334438. See correction [Ait71].

Aitkin:1971:CST

[Ait71] Murray A. Aitkin. Correction: “Some Tests for Correlation Matrices”. Biometrika, 58(1):245, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334343. See [Ait69].

Ajne:1968:STU

[Ajn68] B. Ajne. A simple test for uniformity of a circular distribution. Biometrika, 55(2):343–354, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334875.

Alling:1966:CST

[All66] David W. Alling. Closed sequential tests for binomial probabilities. Biometrika, 53(1/2):73–84, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2334054. See correction [All68].

Alling:1967:TR

[All67] David W. Alling. Tests of relatedness. Biometrika, 54(3/4):459–469, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335037.

Alling:1968:CCS

[All68] David W. Alling. Correction: “Closed Sequential Tests for Binomial Probabilities”. Biometrika, 55(1):268, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334477. See [All66].

Alway:1962:DNC

[Alw62] G. G. Alway. The distribution of the number of circular triads in paired comparisons. Biometrika, 49(1/2):265–269, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333494.

Amos:1963:APP

[Amo63] D. E. Amos. Additional percentage points for the incomplete beta distribution. Biometrika, 50(3/4):449–457, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333913.

Amos:1964:RCN

[Amo64] D. E. Amos. Representations of the central and non-central t distributions. Biometrika, 51(3/4):451–458, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334151.

Anonymous:1960:BMa

[Ano60a] Anonymous. Back matter. Biometrika, 47(1/2):??, June 1960. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332999.

Anonymous:1960:BMb

[Ano60b] Anonymous. Back matter. Biometrika, 47(3/4):??, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333341.

Anonymous:1960:BRBg

[Ano60c] Anonymous. Book review: Journal of Research. B. Mathematics and Mathematical Physics. Biometrika, 47(3/4):490, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333338.

Anonymous:1960:BRBf

[Ano60d] Anonymous. Book review: Rectangular-Polar Conversion Table, by E. H. Neville. Biometrika, 47(1/2):216, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332996.

Anonymous:1960:BRBh

[Ano60e] Anonymous. Book review: Sampling Inspection Tables. Single and Double Sampling. by H. F. Dodge; H. G. Romig. Biometrika, 47(3/4):490, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333339.

Anonymous:1960:BRBe

[Ano60f] Anonymous. Book review: Tables and Percentage Points of the Distribution Function of a Product, by Leo A. Aroian. Biometrika, 47(1/2):215, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332995.

Anonymous:1960:BRBd

[Ano60g] Anonymous. Book review: Tables of Natural Logarithms for Arguments between Five and Ten to Sixteen Decimal Places, by U.S. Department of Commerce. Biometrika, 47(1/2):215, June 1960. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332994.

Anonymous:1960:BRBc

[Ano60h] Anonymous. Book review: Tables of the Bivariate Normal Distribution Function and Related Functions, by Gertrude Blanch. Biometrika, 47(1/2):215, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332993.

Anonymous:1960:BRBa

[Ano60i] Anonymous. Book review: Tables of the Normal Probability Integral, by D. T. Monk; D. B. Owen. Biometrika, 47(1/2):214, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332991.

Anonymous:1960:BRBb

[Ano60j] Anonymous. Book review: The Bivariate Normal Probability Distribution. by D. B. Owen. Biometrika, 47(1/2):214–215, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332992.

Anonymous:1960:CWM

[Ano60k] Anonymous. Corrigenda: Wahrscheinlichkeitsrechnung und Mathematische Statistik. Biometrika, 47(1/2):216, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332997. See [Ela59].

Anonymous:1960:FMa

[Ano60l] Anonymous. Front matter. Biometrika, 47(1/2):??, June 1960. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332951.

Anonymous:1960:FMb

[Ano60m] Anonymous. Front matter. Biometrika, 47(3/4):??, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333293.

Anonymous:1960:OBRa

[Ano60n] Anonymous. Other books received. Biometrika, 47(1/2):217, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332998.

Anonymous:1960:OBRb

[Ano60o] Anonymous. Other books received. Biometrika, 47(3/4):491–492, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333340.

Anonymous:1960:VI

[Ano60p] Anonymous. Volume information. Biometrika, 47(3/4):i–viii, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333292.

Anonymous:1961:BMa

[Ano61a] Anonymous. Back matter. Biometrika, 48(1/2):??, June 1961. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333177.

Anonymous:1961:BMb

[Ano61b] Anonymous. Back matter. Biometrika, 48(3/4):??, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332802.

Anonymous:1961:EN

[Ano61c] Anonymous. Editorial note. Biometrika, 48(1/2):55, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333129.

Anonymous:1961:FMa

[Ano61d] Anonymous. Front matter. Biometrika, 48(1/2):??, June 1961. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333124.

Anonymous:1961:FMb

[Ano61e] Anonymous. Front matter. Biometrika, 48(3/4):??, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332747.

Anonymous:1961:OBRa

[Ano61f] Anonymous. Other books received. Biometrika, 48(1/2):239, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333176.

Anonymous:1961:OBRb

[Ano61g] Anonymous. Other books received. Biometrika, 48(3/4):485, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332801.

Anonymous:1961:VI

[Ano61h] Anonymous. Volume information. Biometrika, 48(3/4):i–viii, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332746.

Anonymous:1962:BMa

[Ano62a] Anonymous. Back matter. Biometrika, 49(1/2):??, June 1962. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333522.

Anonymous:1962:BMb

[Ano62b] Anonymous. Back matter. Biometrika, 49(3/4):??, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334015.

Anonymous:1962:FMa

[Ano62c] Anonymous. Front matter. Biometrika, 49(1/2):i–xii, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333463.

Anonymous:1962:FMb

[Ano62d] Anonymous. Front matter. Biometrika, 49(3/4):i–xii, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333963.

Anonymous:1962:OBRa

[Ano62e] Anonymous. Other books received. Biometrika, 49(1/2):296, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333521.

Anonymous:1962:OBRb

[Ano62f] Anonymous. Other books received. Biometrika, 49(3/4):588, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334014.

Anonymous:1962:VI

[Ano62g] Anonymous. Volume information. Biometrika, 49(3/4):i–viii, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333962.

Anonymous:1963:BMa

[Ano63a] Anonymous. Back matter. Biometrika, 50(1/2):??, June 1963. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333789.

Anonymous:1963:BMb

[Ano63b] Anonymous. Back matter. Biometrika, 50(3/4):??, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333947.

Anonymous:1963:BRBg

[Ano63c] Anonymous. Book review: A First Course in Mathematical Statistics by C. E. Weatherburn. Biometrika, 50(1/2):233, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333787.

Anonymous:1963:BRBc

[Ano63d] Anonymous. Book review: Calculus of Finite Differences by C. Jordan. Biometrika, 50(1/2):232, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333783.

Anonymous:1963:BRBe

[Ano63e] Anonymous. Book review: Lectures on the Calculus of Variations by O. Bolza. Biometrika, 50(1/2):233, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333785.

Anonymous:1963:BRBf

[Ano63f] Anonymous. Book review: Numerical Analysis by Z. Kopal. Biometrika, 50(1/2):233, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333786.

Anonymous:1963:BRBd

[Ano63g] Anonymous. Book review: Random Variables and Probability Distributions by H. Cramér. Biometrika, 50(1/2):232–233, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333784.

Anonymous:1963:BRBi

[Ano63h] Anonymous. Book review: Table of Expectations of Mean Squares in the Analysis of Variance for Crossed Classifications, by K. Abt. Biometrika, 50(3/4):556, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333945.

Anonymous:1963:BRBb

[Ano63i] Anonymous. Book review: Tables of Integrals and Other Mathematical Data by H. B. Dwight. Biometrika, 50(1/2):232, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333782.

Anonymous:1963:BRBa

[Ano63j] Anonymous. Book review: Tables of the Hypergeometric Probability Distribution by G. J. Lieberman; D. B. Owen. Biometrika, 50(1/2): 232, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333781.

Anonymous:1963:BRBh

[Ano63k] Anonymous. Book review: Zeitschrift fur Wahrscheinlichkeitstheorie und Verwandte Gebiete. by L. Schmetterer. Biometrika, 50 (3/4):555–556, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333944.

Anonymous:1963:FMa

[Ano63l] Anonymous. Front matter. Biometrika, 50(1/2):i–xii, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333740.

Anonymous:1963:FMb

[Ano63m] Anonymous. Front matter. Biometrika, 50(3/4):i–xii, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333894.

Anonymous:1963:OBRa

[Ano63n] Anonymous. Other books received. Biometrika, 50(1/2):234, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333788.

Anonymous:1963:OBRb

[Ano63o] Anonymous. Other books received. Biometrika, 50(3/4):557, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333946.

Anonymous:1963:VI

[Ano63p] Anonymous. Volume information. Biometrika, 50(3/4):i–viii, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333893.

Anonymous:1964:BMa

[Ano64a] Anonymous. Back matter. Biometrika, 51(1/2):??, June 1964. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334247.

Anonymous:1964:BMb

[Ano64b] Anonymous. Back matter. Biometrika, 51(3/4):??, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334187.

Anonymous:1964:BRBg

[Ano64c] Anonymous. Book review: Bibliography of Basic Tests and Monographs on Statistical Methods. by W. R. Buckland; R. Fox. Biometrika, 51(3/4):535, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334180.

Anonymous:1964:BRBf

[Ano64d] Anonymous. Book review: Bibliography of Statistical Literature, 1950–58, by M. G. Kendall; Alison G. Doig. Biometrika, 51 (3/4):535, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334179.

Anonymous:1964:BRBk

[Ano64e] Anonymous. Book review: Episodes from the Early History of Mathematics by Asger Aaboe. Biometrika, 51(3/4):536, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334184.

Anonymous:1964:BRBh

[Ano64f] Anonymous. Book review: Introduction to Statistical Inference. by H. Freeman. Biometrika, 51(3/4):535, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334181.

Anonymous:1964:BRBj

[Ano64g] Anonymous. Book review: Méthodes Statistiques à L’Usage des Médecins et des Biologistes by Daniel Schwartz. Biometrika, 51 (3/4):536, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334183.

Anonymous:1964:BRBl

[Ano64h] Anonymous. Book review: Some Aspects of the Relative Efficiencies of Indian Languages. by B. A. Ramakrishna. Biometrika, 51(3/4):536, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334185.

Anonymous:1964:BRBe

[Ano64i] Anonymous. Book review: Statistical Research Memoirs by J. Neyman; E. S. Pearson. Biometrika, 51(1/2):296, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334245.

Anonymous:1964:BRBi

[Ano64j] Anonymous. Book review: Tables of Random Permutations. by L. E. Moses; R. V. Oakford. Biometrika, 51(3/4):535, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334182.

Anonymous:1964:BRBd

[Ano64k] Anonymous. Book review: The Foundations of Statistical Inference by L. J. Savage. Biometrika, 51(1/2):295, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334244.

Anonymous:1964:BRBc

[Ano64l] Anonymous. Book review: The Laplace Transform: An Introduction. by Earl D. Rainville. Biometrika, 51(1/2):295, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334242.

Anonymous:1964:BRBb

[Ano64m] Anonymous. Book review: The Logic of Chance by John Venn. Biometrika, 51(1/2):295, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334241.

Anonymous:1964:BRBa

[Ano64n] Anonymous. Book review: The Physical Anthropology of Southern Nigeria by P. A. Talbot; H. Mulhall. Biometrika, 51(1/2):295, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334240.

Anonymous:1964:BCS

[Ano64o] Anonymous. Information from Censored Samples, by Carl-Erik Särndal. Biometrika, 51(1/2):295, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334243.

Anonymous:1964:FMa

[Ano64p] Anonymous. Front matter. Biometrika, 51(1/2):i–xii, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334188.

Anonymous:1964:FMb

[Ano64q] Anonymous. Front matter. Biometrika, 51(3/4):i–xii, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334135.

Anonymous:1964:OBRa

[Ano64r] Anonymous. Other books received. Biometrika, 51(1/2):297, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334246.

Anonymous:1964:OBRb

[Ano64s] Anonymous. Other books received. Biometrika, 51(3/4):537, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334186.

Anonymous:1964:VI

[Ano64t] Anonymous. Volume information. Biometrika, 51(3/4):i–ix, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334134.

Anonymous:1965:BMa

[Ano65a] Anonymous. Back matter. Biometrika, 52(1/2):??, June 1965. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333847.

Anonymous:1965:BMb

[Ano65b] Anonymous. Back matter. Biometrika, 52(3/4):??, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333739.

Anonymous:1965:BR

[Ano65c] Anonymous. Books received. Biometrika, 52(1/2):307–308, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333846.

Anonymous:1965:EA

[Ano65d] Anonymous. Editorial arrangements. Biometrika, 52(1/2):1, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333807.

Anonymous:1965:FMa

[Ano65e] Anonymous. Front matter. Biometrika, 52(1/2):i–xii, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333806.

Anonymous:1965:FMb

[Ano65f] Anonymous. Front matter. Biometrika, 52(3/4):i–xii, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333684.

Anonymous:1965:OBR

[Ano65g] Anonymous. Other books received. Biometrika, 52(3/4):675, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333738.

Anonymous:1965:VI

[Ano65h] Anonymous. Volume information. Biometrika, 52(3/4):i–viii, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333683.

Anonymous:1966:A

[Ano66a] Anonymous. Announcement. Biometrika, 53(3/4):309, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333637.

Anonymous:1966:BMa

[Ano66b] Anonymous. Back matter. Biometrika, 53(1/2):??, June 1966. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334108.

Anonymous:1966:BMb

[Ano66c] Anonymous. Back matter. Biometrika, 53(3/4):??, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333682.

Anonymous:1966:BRB

[Ano66d] Anonymous. Book review: R.S.M. Tables No. 8: Tables of Natural and Common Logarithms. Biometrika, 53(1/2):306, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334106.

Anonymous:1966:FMa

[Ano66e] Anonymous. Front matter. Biometrika, 53(1/2):i–xii, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334047.

Anonymous:1966:FMb

[Ano66f] Anonymous. Front matter. Biometrika, 53(3/4):i–xii, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333636.

Anonymous:1966:NPS

[Ano66g] Anonymous. New publications in statistics and probability. Biometrika, 53(3/4):636, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333681.

Anonymous:1966:OBR

[Ano66h] Anonymous. Other books received. Biometrika, 53(1/2):307, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334107.

Anonymous:1966:VI

[Ano66i] Anonymous. Volume information. Biometrika, 53(3/4):i–ix, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333635.

Anonymous:1967:BMa

[Ano67a] Anonymous. Back matter. Biometrika, 54(1/2):i–xii, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333892.

Anonymous:1967:BMb

[Ano67b] Anonymous. Back matter. Biometrika, 54(3/4):i–xii, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335076.

Anonymous:1967:BR

[Ano67c] Anonymous. Book reviews. Biometrika, 54(3/4):692, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335074.

Anonymous:1967:FMa

[Ano67d] Anonymous. Front matter. Biometrika, 54(1/2):??, June 1967. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333848.

Anonymous:1967:FMb

[Ano67e] Anonymous. Front matter. Biometrika, 54(3/4):??, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335026.

Anonymous:1967:NPSa

[Ano67f] Anonymous. New publications in statistics and probability. Biometrika, 54(1/2):339, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333891.

Anonymous:1967:NPSb

[Ano67g] Anonymous. New publications in statistics and probability. Biometrika, 54(3/4):693, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335075.

Anonymous:1967:VI

[Ano67h] Anonymous. Volume information. Biometrika, 54(3/4):i–viii, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335025.

Anonymous:1968:BMa

[Ano68a] Anonymous. Back matter. Biometrika, 55(1):ii–viii, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334479.

Anonymous:1968:BMb

[Ano68b] Anonymous. Back matter. Biometrika, 55(2):??, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334898.

Anonymous:1968:BMc

[Ano68c] Anonymous. Back matter. Biometrika, 55(3):599, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334278.

Anonymous:1968:FMa

[Ano68d] Anonymous. Front matter. Biometrika, 55(1):??, March 1968. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334447.

Anonymous:1968:FMb

[Ano68e] Anonymous. Front matter. Biometrika, 55(2):??, July 1968. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334865.

Anonymous:1968:FMc

[Ano68f] Anonymous. Front matter. Biometrika, 55(3):??, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334249.

Anonymous:1968:NPSa

[Ano68g] Anonymous. New publications in statistics and probability. Biometrika, 55(1):i, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334478.

Anonymous:1968:NPSb

[Ano68h] Anonymous. New publications in statistics and probability. Biometrika, 55(2):443–444, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334897.

Anonymous:1968:NPSc

[Ano68i] Anonymous. New publications in statistics and probability. Biometrika, 55(3):598, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334277.

Anonymous:1968:VI

[Ano68j] Anonymous. Volume information. Biometrika, 55(3):i–viii, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334248.

Anonymous:1969:BMa

[Ano69a] Anonymous. Back matter. Biometrika, 56(1):i–vii, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334717.

Anonymous:1969:BMb

[Ano69b] Anonymous. Back matter. Biometrika, 56(2):461, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334446.

Anonymous:1969:BMc

[Ano69c] Anonymous. Back matter. Biometrika, 56(3):718, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334684.

Anonymous:1969:BIS

[Ano69d] Anonymous. Biometrika index, supplement: Author [and review] index: Vols. 49–56 (1962–1969). Biometrika, 56(3):693–717, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334683.

Anonymous:1969:FMa

[Ano69e] Anonymous. Front matter. Biometrika, 56(1):??, March 1969. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334685.

Anonymous:1969:FMb

[Ano69f] Anonymous. Front matter. Biometrika, 56(2):??, August 1969. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334416.

Anonymous:1969:FMc

[Ano69g] Anonymous. Front matter. Biometrika, 56(3):??, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334651.

Anonymous:1969:NPSa

[Ano69h] Anonymous. New publications in statistics and probability. Biometrika, 56(1):230, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334716.

Anonymous:1969:NPSb

[Ano69i] Anonymous. New publications in statistics and probability. Biometrika, 56(2):459, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334445.

Anonymous:1969:NPSc

[Ano69j] Anonymous. New publications in statistics and probability. Biometrika, 56(3):692, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334682.

Anonymous:1969:VI

[Ano69k] Anonymous. Volume information. Biometrika, 56(3):i–viii, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334650.

Aitkin:1968:TCM

[ANR68] M. A. Aitkin, W. C. Nelson, and Karen H. Reinfurt. Tests for correlation matrices. Biometrika, 55(2):327–334, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334873.

Archibald:1961:BRB

[Arc61] G. C. Archibald. Book review: Strategy and Market Structure. Competition, Oligopoly, and the Theory of Games, by Martin Shubik; Oskar Morgenstern. Biometrika, 48(3/4):479–480, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332788.

Archbold:1962:BRB

[Arc62] J. W. Archbold. Book review: Projective Geometry of n Dimensions by Otto Schreider; Emanuel Sperner. Biometrika, 49(3/4):586, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334012.

Archbold:1964:BRB

[Arc64] J. W. Archbold. Book review: A Course in the Geometry of n Dimensions by M. G. Kendall. Biometrika, 51(3/4):532–533, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334173.

Arnold:1964:PSM

[Arn64] H. J. Arnold. Permutation support for multivariate techniques. Biometrika, 51(1/2):65–70, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334196.

Ashford:1964:GMQ

[AS64] J. R. Ashford and C. S. Smith. General models for quantal response to the joint action of a mixture of drugs. Biometrika, 51 (3/4):413–428, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334148.

Aitchison:1965:SPS

[AS65] J. Aitchison and Diane Sculthorpe. Some problems of statistical prediction. Biometrika, 52(3/4):469–483, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333699.

Ashford:1966:MNI

[AS66] J. R. Ashford and C. S. Smith. Models for the non-interactive joint action of a mixture of stimuli in biological assay. Biometrika, 53(1/2):49–59, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334052.

Ashford:1960:QRA

[ASB60] J. R. Ashford, C. S. Smith, and Susannah Brown. The quantal response analysis of a series of biological assays on the same subjects. Biometrika, 47(1/2):23–32, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332954.

Atiqullah:1961:PBD

[Ati61] M. Atiqullah. On a property of balanced designs. Biometrika, 48(1/2):215–218, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333151.

Atiqullah:1962:ERV

[Ati62] M. Atiqullah. The estimation of residual variance in quadratically balanced least-squares problems and the robustness of the F -test. Biometrika, 49(1/2):83–91, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333469.

Atiqullah:1964:RCA

[Ati64] M. Atiqullah. The robustness of the covariance analysis of a oneway classification. Biometrika, 51(3/4):365–372, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334142.

Atiqullah:1968:ESM

[Ati68] M. Atiqullah. On estimation by the sweep-out method. Biometrika, 55(2):305–311, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2334870. See correction [Ati69].

Atiqullah:1969:CES

[Ati69] M. Atiqullah. Corrections: “On Estimation by the Sweep-Out Method”. Biometrika, 56(1):229, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334715. See [Ati68].

Atkinson:1969:TDB

[Atk69a] Anthony C. Atkinson. A test for discriminating between models. Biometrika, 56(2):337–347, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334426.

Atkinson:1969:URC

[Atk69b] Anthony C. Atkinson. The use of residuals as a concomitant variable. Biometrika, 56(1):33–41, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334688.

Bailey:1963:SSE

[Bai63] Norman T. J. Bailey. The simple stochastic epidemic: A complete solution in terms of known functions. Biometrika, 50(3/4):235–240, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333895.

Bailey:1968:PAS

[Bai68a] Norman T. J. Bailey. A perturbation approximation to the simple stochastic epidemic in a large population. Biometrika, 55(1):199– 209, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334464.

Bailey:1968:SBD

[Bai68b] Norman T. J. Bailey. Stochastic birth, death and migration processes for spatially distributed populations. Biometrika, 55(1):189– 198, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334463.

Bardwell:1960:CCS

[Bar60a] George E. Bardwell. On certain characteristics of some discrete distributions. Biometrika, 47(3/4):473–475, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333322.

Barnard:1960:BRB

[Bar60b] G. A. Barnard. Book review: Operations Research, by James H. Batchelor. Biometrika, 47(1/2):211, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332986.

Barton:1960:BRB

[Bar60c] D. E. Barton. Book review: International Journal of Abstracts: Statistical Theory and Method, by W. R. Buckland. Biometrika, 47(1/2):209, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332981.

Bartholomew:1961:OTA

[Bar61a] D. J. Bartholomew. Ordered tests in the analysis of variance. Biometrika, 48(3/4):325–332, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332754.

Bartlett:1961:BRB

[Bar61b] M. S. Bartlett. Book review: An Introduction to Statistical Communication Theory, by David Middleton. Biometrika, 48(3/4):477, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332785.

Barton:1961:BRB

[Bar61c] D. E. Barton. Book review: The Theory of Storage, by P. A. P. Moran. Biometrika, 48(1/2):231, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333160.

Barton:1961:UES

[Bar61d] D. E. Barton. Unbiased estimation of a set of probabilities. Biometrika, 48(1/2):227–229, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333156.

Barnett:1962:LST

[Bar62a] V. D. Barnett. Large sample tables of percentage points for Hartley’s correction to Bartlett’s criterion for testing the homogeneity of a set of variances. Biometrika, 49(3/4):487–494, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333982.

Bartlett:1962:BRB

[Bar62b] M. S. Bartlett. Book review: Random Processes by Murray Rosenblat. Biometrika, 49(3/4):577, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333996.

Barton:1962:BRBc

[Bar62c] D. E. Barton. Book review: Analysing Qualitative Data, by A. E. Maxwell. Biometrika, 49(1/2):291–292, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333512.

Barton:1962:BRBd

[Bar62d] D. E. Barton. Book review: Grundlagen Der Analysis by Edmund Landau; F. Steinhardt. Biometrika, 49(3/4):586–587, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334013.

Barton:1962:BRBa

[Bar62e] D. E. Barton. Book review: Modern Probability Theory and its Applications, by E. Parzen. Biometrika, 49(1/2):287, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333506.

Barton:1962:BRBb

[Bar62f] D. E. Barton. Book review: Stochastic Population Models in Ecology and Epidemiology, by M. S. Bartlett. Biometrika, 49(1/2):288–289, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333508.

Bartlett:1963:BRB

[Bar63a] M. S. Bartlett. Book review: Random Wavelets and Cybernetic Systems, by Enders A. Robinson. Biometrika, 50(3/4):549, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333935.

Barton:1963:BRBb

[Bar63b] D. E. Barton. Book review: An Elementary Introduction to the Theory of Probability by B. V. Gnedenko; A. Ya. Khinchin; W. R.

Stahl. Biometrika, 50(1/2):224–225, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333769.

Barton:1963:BRBc

[Bar63c] D. E. Barton. Book review: Contributions to the Theory of Estimation from Grouped and Partially Grouped Samples by G. Kulldorff. Biometrika, 50(1/2):227–228, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333772.

Barton:1963:BRBa

[Bar63d] D. E. Barton. Book review: Limit Distributions for Sums of Independent Random Variables by B. V. Gmedenko; A. N. Kolmogorov; K. L. Chung. Biometrika, 50(1/2):223–224, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333768.

Barton:1963:BRBf

[Bar63e] D. E. Barton. Book review: Linear Differential Operators, by C. Lanczos. London; New York; Princeton and Toronto: D. Van Nostrand Co. 1961, 564 pp. Biometrika, 50(3/4):555, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333943.

Barton:1963:BRBe

[Bar63f] D. E. Barton. Book review: Stichproben in der Amtlichen Statistik, by Gerhard Furst. Biometrika, 50(3/4):553, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333940.

Barton:1963:BRBd

[Bar63g] D. E. Barton. Book review: The Use and Abuse of Statistics, by W. J. Reichmann. Biometrika, 50(3/4):552, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333938.

Bartholomew:1964:BRB

[Bar64a] D. J. Bartholomew. Book review: Probability Theory and Mathematical Statistics, by Marek Fisz. Biometrika, 51(1/2):289, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334227.

Bartlett:1964:SAT

[Bar64b] M. S. Bartlett. The spectral analysis of two-dimensional point processes. Biometrika, 51(3/4):299–311, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334136. See corrigenda [Bar65b].

Bartholomew:1965:CSB

[Bar65a] D. J. Bartholomew. A comparison of some Bayesian and frequentist inferences. Biometrika, 52(1/2):19–35, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333809.

Bartlett:1965:CSA

[Bar65b] M. S. Bartlett. Corrigenda: “The spectral analysis of twodimensional point processes”. Biometrika, 52(1/2):305, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333845. See [Bar64b].

Barton:1965:BRBc

[Bar65c] D. E. Barton. Book review: General Stochastic Processes in the Theory of Queues, by Vaclav E. Benes. Biometrika, 52(3/4):673, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333735.

Barton:1965:BRBa

[Bar65d] D. E. Barton. Book review: The Elements of Stochastic Processes with Applications to the Natural Sciences, by Norman T. J. Bailey. Biometrika, 52(3/4):671, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333730.

Barton:1965:BRBb

[Bar65e] D. E. Barton. Book review: The Theory of Branching Processes, by Theodore E. Harris. Biometrika, 52(3/4):672, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333733.

Barnett:1966:EML

[Bar66a] V. D. Barnett. Evaluation of the maximum-likelihood estimator where the likelihood equation has multiple roots. Biometrika, 53(1/2):151–165, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334061.

Bartlett:1966:BRB

[Bar66b] M. S. Bartlett. Book review: Logic of Statistical Inference, by Ian Hacking. Biometrika, 53(3/4):631–633, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333679.

Barton:1966:BRBa

[Bar66c] D. E. Barton. Book review: Introduction to Probability and Statistics from a Bayesian Viewpoint, by D. V. Lindley. Biometrika, 53(1/2):296–297, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334092.

Barton:1966:BRBd

[Bar66d] D. E. Barton. Book review: Percentage Points of the Sum V_{1}^{(s)} of s Roots (s = 1–50), by T. A. Mijares. Biometrika, 53(1/2):303–304, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334102.

Barton:1966:BRBe

[Bar66e] D. E. Barton. Book review: R.S.M. Tables No. 10: Bessel Functions. Biometrika, 53(1/2):305, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334104.

Barton:1966:BRBc

[Bar66f] D. E. Barton. Book review: Statistical Method in Biological Assay, by D. J. Finney. Biometrika, 53(1/2):301–302, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334099.

Barton:1966:BRBb

[Bar66g] D. E. Barton. Book review: The Statistical Processes of Evolutionary Theory, by P. A. P. Moran. Biometrika, 53(1/2):299–300, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334096.

Barnett:1967:NLS

[Bar67a] V. D. Barnett. A note on linear structural relationships when both residual variances are known. Biometrika, 54(3/4):670–672, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335062.

Bartlett:1967:SRA

[Bar67b] M. S. Bartlett. Some remarks on the analysis of time-series. Biometrika, 54(1/2):25–38, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333850.

Bartlett:1969:DAC

[Bar69] M. S. Bartlett. Distributions associated with cell populations. Biometrika, 56(2):391–400, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334431.

Batschelet:1963:THE

[Bat63] Edward Batschelet. Testing hypotheses and estimating parameters in human genetics if the age of onset is random. Biometrika, 50 (3/4):265–279, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333898.

Bartholomew:1966:CSB

[BB66] D. J. Bartholomew and E. E. Bassett. A comparison of some Bayesian and frequentist inferences. II. Biometrika, 53(1/2):262– 264, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334077.

Blyth:1960:EPC

[BC60] Colin R. Blyth and Gregory L. Curme. Estimation of a parameter in the classical occupancy problem. Biometrika, 47(1/2):180–185, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332971.

Burr:1961:LRC

[BC61] E. J. Burr and Gwenda Cane. Longest run of consecutive observations having a specified attribute. Biometrika, 48(3/4):461–465, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332771.

Barton:1961:CSM

[BD61] D. E. Barton and F. N. David. The central sampling moments of the mean in samples from a finite population (Aty’s formulae and Madow’s central limit). Biometrika, 48(1/2):199–201, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333146.

Box:1963:CSO

[BD63] George E. P. Box and Norman R. Draper. The choice of a second order rotatable design. Biometrika, 50(3/4):335–352, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333904. See

corrigenda [BD65a].

Box:1965:CCS

[BD65a] G. E. P. Box and N. R. Draper. Corrigenda: “The choice of a second order rotatable design”. Biometrika, 52(1/2):305, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333843. See [BD63].

Box:1965:BEC

[BD65b] George E. P. Box and Norman R. Draper. The Bayesian estimation of common parameters from several responses. Biometrika, 52 (3/4):355–365, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333689.

Barton:1966:PPR

[BD66] D. E. Barton and F. N. David. A persistence problem in renewal theory: Robert the Bruce’s spider. Biometrika, 53(1/2):255–258, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334075.

Barton:1968:SOC

[BD68] D. E. Barton and F. N. David. Sequential occupancy with classification. Biometrika, 55(1):229–241, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334467.

Barton:1960:PNN

[BDF60] D. E. Barton, F. N. David, and Evelyn Fix. The polykays of the natural numbers. Biometrika, 47(1/2):53–59, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332957.

Barton:1962:PCM

[BDF62] D. E. Barton, F. N. David, and Evelyn Fix. Persistence in a chain of multiple events when there is simple dependence. Biometrika, 49(3/4):351–357, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333969.

Barton:1963:RPC

[BDF63] D. E. Barton, F. N. David, and E. Fix. Random points in a circle and the analysis of chromosome patterns. Biometrika, 50(1/2):23–29, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333743.

Barton:1960:TSE

[BDM60] D. E. Barton, F. N. David, and M. Merrington. Tables for the solution of the exponential equation, exp(−a)+ka = 1. Biometrika, 47(3/4):439–445, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333314.

Barton:1963:TSE

[BDM63] D. E. Barton, F. N. David, and M. Merrington. Table for the solution of the exponential equation exp(b) − b/(1 − p) = 1. Biometrika, 50(1/2):169–176, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333757.

Barton:1960:SPD

[BDO60] D. E. Barton, F. N. David, and Anne F. O’Neill. Some properties of the distribution of the logarithm of non-central F. Biometrika, 47(3/4):417–431, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333312.

Beale:1961:BRB

[Bea61] E. M. L. Beale. Book review: Introduction to Linear Programming, by Walter W. Garvin. Biometrika, 48(3/4):477–478, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332786.

Beer:1963:BRB

[Bee63] Stafford Beer. Book review: Quantity and Quality. The Hayden Colloquium on Scientific Method and Concept, by Daniel Lerner. Biometrika, 50(3/4):549–551, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333936.

Bennett:1963:PLD

[Ben63] G. Bennett. On the probability of large deviations from the expectation for sums of bounded, independent random variables. Biometrika, 50(3/4):528–535, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333923.

Bennett:1965:UBM

[Ben65] G. Bennett. Upper bounds on the moments and probability inequalities for the sum of independent, bounded random variables. Biometrika, 52(3/4):559–569, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333705.

Bennett:1968:OSP

[Ben68] G. Bennett. A one-sided probability inequality for the sum of independent, bounded random variables. Biometrika, 55(3):565–569, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334262.

Berkson:1960:NFL

[Ber60] Joseph Berkson. Nomograms for fitting the logistic function by maximum likelihood. Biometrika, 47(1/2):121–141, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332965.

Bernoulli:1961:MPC

[Ber61] Daniel Bernoulli. The most probable choice between several discrepant observations and the formation therefrom of the most likely induction. Biometrika, 48(1/2):3–13, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www. jstor.org/stable/2333125. Translated from Latin by C. G. Allan.

Berenblut:1968:CDB

[Ber68] I. I. Berenblut. Changeover designs balanced for the linear component of first residual effects. Biometrika, 55(2):297–303, July 1968.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334869.

Beran:1969:DNT

[Ber69] R. J. Beran. The derivation of nonparametric two-sample tests from tests for uniformity of a circular distribution. Biometrika, 56(3):561–570, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334664.

Bhat:1960:NSR

[BG60] B. R. Bhat and J. Gani. A note on sufficiency in regular Markov chains. Biometrika, 47(3/4):452–457, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333317.

Bradley:1962:APM

[BG62] Ralph A. Bradley and John J. Gart. The asymptotic properties of ML estimators when sampling from associated populations. Biometrika, 49(1/2):205–214, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333482.

Bland:1966:DRM

[BGKO66] R. P. Bland, R. D. Gilbert, C. H. Kapadia, and D. B. Owen. On the distributions of the range and mean range for samples from a normal distribution. Biometrika, 53(1/2):245–248, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334072. See cor-

rection [BGKO71].

Bland:1971:CDR

[BGKO71] R. P. Bland, R. D. Gilbert, C. H. Kapadia, and D. B. Owen. Correction: “On the Distributions of the Range and Mean Range for Samples from a Normal Distribution”. Biometrika, 58(1):245, April 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334344. See

[BGKO66].

Bartlett:1960:CTE

[BGL60] M. S. Bartlett, J. C. Gower, and P. H. Leslie. A comparison of theoretical and empirical results for some stochastic population models. Biometrika, 47(1/2):1–11, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332952.

Bennett:1960:PFE

[BH60a] B. M. Bennett and P. Hsu. On the power function of the exact test for the 2 × 2 contingency table. Biometrika, 47(3/4):393–398, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333309.

See corrigenda [BH61a].

Blyth:1960:TNS

[BH60b] Colin R. Blyth and David W. Hutchinson. Table of Neymanshortest unbiased confidence intervals for the binomial parameter. Biometrika, 47(3/4):381–391, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333308.

Bennett:1961:CPF

[BH61a] B. M. Bennett and P. Hsu. Corrigenda: “On the Power Function of the Exact test for the 2 × 2 Contingency table”. Biometrika, 48(3/4):475, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332779. See [BH60a].

Blyth:1961:TNS

[BH61b] Colin R. Blyth and David W. Hutchinson. Table of Neymanshortest unbiased confidence intervals for the Poisson parameter. Biometrika, 48(1/2):191–194, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333143.

Bhate:1960:ADS

[Bha60] D. H. Bhate. Approximation to the distribution of the sample size for sequential tests.: II. Tests of composite hypotheses. Biometrika, 47(1/2):190–193, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332973.

Bhattacharya:1967:RAP

[Bha67] S. K. Bhattacharya. A result on accident proneness. Biometrika, 54(1/2):324–325, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333881.

Biggers:1961:EMO

[Big61] J. D. Biggers. Estimation of missing observations in split-plot experiments where whole-plots are missing or mixed-up. Biometrika, 48(3/4):468–472, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332773.

Bhattacharyya:1969:HBS

[BJ69] G. K. Bhattacharyya and Richard A. Johnson. On Hodges’s bivariate sign test and a test for uniformity of a circular distribution. Biometrika, 56(2):446–449, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334439.

Beale:1967:DVM

[BKM67] E. M. L. Beale, M. G. Kendall, and D. W. Mann. The discarding of variables in multivariate analysis. Biometrika, 54(3/4):357–366, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335028.

Bartoszynski:1962:SMA

[BLN62] R. Bartoszyński, L. Lubińska, and S. Niemierko. A stochastic model of AChE transportation in the peripheral nerve trunks. Biometrika, 49(3/4):447–454, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333978.

Blom:1960:HBDa

[Blo60a] Gunnar Blom. Hierarchical birth and death processes: I. Theory. Biometrika, 47(3/4):235–244, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333295.

Blom:1960:HBDb

[Blo60b] Gunnar Blom. Hierarchical birth and death processes: II. Applications. Biometrika, 47(3/4):245–251, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333296.

Blom:1963:SCT

[Blo63] Gunnar Blom. Some contributions to the theory of machine interference. Biometrika, 50(1/2):135–143, June 1963. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333754.

Banerjee:1965:BMC

[BM65] K. S. Banerjee and L. F. Marcus. Bounds in a minimax classification procedure. Biometrika, 52(3/4):653–654, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333722.

Barnett:1967:LEP

[BMS67] F. C. Barnett, K. Mullen, and J. G. Saw. Linear estimates of a population scale parameter. Biometrika, 54(3/4):551–554, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335045.

Bhattacharjee:1964:ENN

[BN64] G. P. Bhattacharjee and Y. Nagendra. Effect of non-normality on a sequential test for mean. Biometrika, 51(1/2):281–287, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334225.

Bhat:1965:LAM

[BN65] B. R. Bhat and B. N. Nagnur. Locally asymptotically most stringent tests and Lagrangian multiplier tests of linear hypotheses. Biometrika, 52(3/4):459–468, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333698.

Bennett:1968:PPR

[BN68] B. M. Bennett and E. Nakamura. Percentage points of the range from a symmetric multinomial distribution. Biometrika, 55(2):377– 379, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334879.

Bush:1959:EQF

[BO59] K. A. Bush and I. Olkin. Extrema of quadratic forms with applications to statistics. Biometrika, 46(3/4):483–486, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333549. See corri-

genda [BO61].

Bush:1961:CEQ

[BO61] K. A. Bush and I. Olkin. Corrigenda: “Extrema of Quadratic Forms with Applications to Statistics”. Biometrika, 48(3/4):474–475, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332777.

See [BO59].

Bohrer:1966:BSD

[Boh66] Robert Bohrer. On Bayes sequential design with two random variables. Biometrika, 53(3/4):469–475, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333652.

Bartlett:1963:DCZ

[BP63] M. S. Bartlett and N. W. Please. Discrimination in the case of zero mean differences. Biometrika, 50(1/2):17–21, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333742.

Bradley:1954:RAI

[Bra54] Ralph Allan Bradley. Rank analysis of incomplete block designs: II. Additional tables for the method of paired comparisons. Biometrika, 41(3/4):502–537, December 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332730. See corrigenda [Bra64].

Bradley:1964:CRA

[Bra64] Ralph A. Bradley. Corrigendum: “The Rank Analysis of Incomplete Block Designs. II. Additional Tables for the Method of Paired Comparisons”. Biometrika, 51(1/2):288, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334226. See [Bra54].

Breitenberger:1963:AND

[Bre63] Ernst Breitenberger. Analogues of the normal distribution on the circle and the sphere. Biometrika, 50(1/2):81–88, June 1963. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333749.

Brillinger:1962:NRC

[Bri62] David R. Brillinger. A note on the rate of convergence of a mean. Biometrika, 49(3/4):574–576, December 1962. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333995.

Brillinger:1965:MAR

[Bri65] David R. Brillinger. A moving average representation for random variables covariance stationary on a finite time interval. Biometrika, 52(1/2):295–297, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333839.

Brillinger:1966:EPC

[Bri66] David R. Brillinger. An extremal property of the conditional expectation. Biometrika, 53(3/4):594–595, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333666.

Brillinger:1969:APS

[Bri69] David R. Brillinger. Asymptotic properties of spectral estimates of second order. Biometrika, 56(2):375–390, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334430.

Brown:1961:SPS

[Bro61] Byron Wm Brown, Jr. Some properties of the Spearman estimator in bioassay. Biometrika, 48(3/4):293–302, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332751.

Brookes:1962:BRB

[Bro62] B. C. Brookes. Book review: Field Computations in Engineering and Physics by A. Thom; C. J. Apelt. Biometrika, 49(3/4):584, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334009.

Brook:1964:DBC

[Bro64] D. Brook. On the distinction between the conditional probability and the joint probability approaches in the specification of nearestneighbour systems. Biometrika, 51(3/4):481–483, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334154.

Brown:1967:TMP

[Bro67] Morton B. Brown. The two-means problem — a secondarily Bayes approach. Biometrika, 54(1/2):85–91, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333853.

Box:1962:FLR

[BT62] G. E. P. Box and G. C. Tiao. A further look at robustness via Bayes’s theorem. Biometrika, 49(3/4):419–432, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333976. See corrigenda [BT63].

Box:1963:CFL

[BT63] G. E. P. Box and G. C. Tiao. Corrigenda: “A Further Look at Robustness via Bayes’s Theorem”. Biometrika, 50(3/4):546, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333929. See

[BT62].

Box:1964:BAI

[BT64a] G. E. P. Box and George C. Tiao. A Bayesian approach to the importance of assumptions applied to the comparison of variances. Biometrika, 51(1/2):153–167, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334203.

Box:1964:NCR

[BT64b] G. E. P. Box and George C. Tiao. A note on criterion robustness and inference robustness. Biometrika, 51(1/2):169–173, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334204.

Box:1965:CLN

[BT65] G. E. P. Box and George C. Tiao. A change in level of a nonstationary time series. Biometrika, 52(1/2):181–192, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333823.

Box:1968:BAS

[BT68] G. E. P. Box and G. C. Tiao. A Bayesian approach to some outlier problems. Biometrika, 55(1):119–129, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334456.

Buckland:1960:BRB

[Buc60] Wm R. Buckland. Book review: Analysis of Straight-Line Data, by Forman S. Acton. Biometrika, 47(3/4):486–487, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333330.

Buckland:1961:BRB

[Buc61] Wm. R. Buckland. Book review: Regression Analysis, by E. J. Williams. Biometrika, 48(1/2):232–233, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333162.

Buckland:1962:BRBa

[Buc62a] W. R. Buckland. Book review: Contributions to Probability and Statistics, by Harold Hotelling; I. Olkin. Biometrika, 49(1/2):286, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333504.

Buckland:1962:BRBc

[Buc62b] W. R. Buckland. Book review: Engineering Statistics by A. H. Bowker; G. J. Lieberman. Biometrika, 49(3/4):583, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334008.

Buckland:1962:BRBd

[Buc62c] W. R. Buckland. Book review: Introduction to Congestion Theory in Telephone Systems by R. Syski. Biometrika, 49(3/4):585, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334011.

Buckland:1962:BRBb

[Buc62d] W. R. Buckland. Book review: Statistical Processes and Reliability Engineering by D. N. Chorafas. Biometrika, 49(3/4):582–583, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334007.

Buckland:1964:BRBa

[Buc64a] W. R. Buckland. Book review: Renewal Theory, by D. R. Cox. Biometrika, 51(1/2):290, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334228.

Buckland:1964:BRBb

[Buc64b] W. R. Buckland. Book review: Statistical Theory of Reliability, by M. Zelen. Biometrika, 51(3/4):531–532, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334171.

Buckland:1965:BRB

[Buc65] W. R. Buckland. Book review: Linear Programming. Methods and Applications, by Saul I. Gass; Elements of Linear Programming with Economic Applications. by R. C. Geary; M. D. McCarthy; M. G. Kendall. Biometrika, 52(3/4):671–672, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333731.

Burr:1960:DKS

[Bur60a] E. J. Burr. The distribution of Kendall’s score S for a pair of tied rankings. Biometrika, 47(1/2):151–171, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332967.

Burrows:1960:CSN

[Bur60b] C. Burrows. Corrigendum: “Some Numerical Results for Waiting Times in the Queue E_{k} /M/1”. Biometrika, 47(3/4):484, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333326. See

[Bur60c].

Burrows:1960:SNR

[Bur60c] C. Burrows. Some numerical results for waiting times in the queue E_{k}/M/1. Biometrika, 47(1/2):202–203, June 1960. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332978. See corrigendum

[Bur60b].

Burt:1962:BRB

[Bur62] Cyril Burt. Book review: Studies in Item Analysis and Prediction by Herbert Solomon. Biometrika, 49(3/4):581–582, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334006.

Box:1962:RNN

[BW62] G. E. P. Box and G. S. Watson. Robustness to non-normality of regression tests. Biometrika, 49(1/2):93–106, June 1962. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333470. See corrigenda

[BW65b].

Bartko:1963:IGM

[BW63] John J. Bartko and Geoffrey A. Watterson. Inference on a genetic model of the Markov chain type. Biometrika, 50(3/4):251–264, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333897.

Bartko:1965:SPS

[BW65a] John J. Bartko and G. A. Watterson. Some problems of statistical inference in absorbing Markov chains. Biometrika, 52(1/2):127–138, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333818.

Box:1965:CRN

[BW65b] G. E. P. Box and G. S. Watson. Corrigenda: “Robustness to Non-Normality of Regression Test”. Biometrika, 52(3/4):669, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333727. See

[BW62].

Campling:1968:SSA

[Cam68] G. E. G. Campling. Serial sampling acceptance schemes for large batches of items where the mean quality has a normal prior distribution. Biometrika, 55(2):393–399, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334882.

Collier:1963:RDR

[CB63] Raymond O. Collier, Jr. and Frank B. Baker. The randomization distribution of F -ratios for the splitplot design — an empirical investigation. Biometrika, 50(3/4):431–438, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333911.

Collier:1966:SMC

[CB66] Raymond O. Collier, Jr. and Frank B. Baker. Some Monte Carlo results on the power of the F -test under permutation in the simple randomized block design. Biometrika, 53(1/2):199–203, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334065.

Chambers:1967:DBA

[CC67] Elizabeth A. Chambers and D. R. Cox. Discrimination between alternative binary response models. Biometrika, 54(3/4):573–578, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335048.

Chanda:1961:CEM

[Cha61] K. C. Chanda. Comparative efficiencies of methods of estimating parameters in linear autoregressive schemes. Biometrika, 48 (3/4):427–432, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332764.

Chanda:1964:AET

[Cha64] K. C. Chanda. Asymptotic expansions for tests of goodness of fit for linear autoregressive schemes. Biometrika, 51(3/4):459–465, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334152.

Chambers:1967:ETP

[Cha67a] C. Chambers. Extension of tables of percentage points of the largest variance ratio, s_{max}^{2}/s_{0}^{2}. Biometrika, 54(1/2):225–227, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333865.

Chambers:1967:MAA

[Cha67b] John M. Chambers. On methods of asymptotic approximation for multivariate distributions. Biometrika, 54(3/4):367–383, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335029.

Chambers:1968:BRB

[Cha68a] J. M. Chambers. Book review: Multivariate Analysis by P. R. Krishnaiah. Biometrika, 55(2):442–443, July 1968. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334895.

Chase:1968:EMP

[Cha68b] G. R. Chase. On the efficiency of matched pairs in Bernoulli trials. Biometrika, 55(2):365–369, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334877.

Chatfield:1969:EPL

[Cha69] C. Chatfield. On estimating the parameters of the logarithmic series and negative binomial distributions. Biometrika, 56(2):411–414, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334434.

Childs:1967:RMN

[Chi67] Donald R. Childs. Reduction of the multivariate normal integral to characteristic form. Biometrika, 54(1/2):293–300, June 1967. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333872.

Chambers:1964:UDS

[CJ64] M. L. Chambers and P. Jarratt. Use of double sampling for selecting best population. Biometrika, 51(1/2):49–64, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334195.

Clemmer:1968:UEB

[CK68] Bennie A. Clemmer and Richard G. Krutchkoff. The use of empirical Bayes estimators in a linear regression model. Biometrika, 55(3):525–534, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2334257. See correction [CK70].

Clemmer:1970:CUE

[CK70] B. A. Clemmer and R. G. Krutchkoff. Correction: “The Use of Empirical Bayes Estimators in a Linear Regression Model”. Biometrika, 57(2):467, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334860. See [CK68].

Claerbout:1966:SFM

[Cla66] Jon F. Claerbout. Spectral factorization of multiple time series. Biometrika, 53(1/2):264–267, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334078.

Clunie:1963:BRB

[Clu63] J. Clunie. Book review: Complex Numbers and Functions, by T. Estermann. Biometrika, 50(3/4):554–555, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333942.

Cacoullos:1965:BFC

[CO65] Theophilos Cacoullos and Ingram Olkin. On the bias of functions of characteristic roots of a random matrix. Biometrika, 52(1/2):87–94, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333814.

Colton:1963:ODS

[Col63] Theodore Colton. Optimal drug screening plans. Biometrika, 50(1/2):31–45, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333744.

Conover:1965:DWR

[Con65] W. J. Conover. The distribution of ∑_{0}^{m} f(Y_{t}), where (Y_{0}, Y_{1}, …) is a realization of a non-homogeneous finite-state Markov chain. Biometrika, 52(1/2):277–279, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333833.

Cooper:1963:SCQ

[Coo63] Paul W. Cooper. Statistical classification with quadratic forms. Biometrika, 50(3/4):439–448, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333912.

Corlett:1962:BRBb

[Cor62a] W. J. Corlett. Book review: L’Anglais Économique. Dictionnaire De Concepts by Edmond Lisle. Biometrika, 49(3/4):580–581, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334004.

Corlett:1962:BRBa

[Cor62b] W. J. Corlett. Book review: The Use of Economic Statistics, by C. A. Blyth. Biometrika, 49(1/2):293–294, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333515.

Corlett:1964:BRB

[Cor64a] W. J. Corlett. Book review: Food Supplies and Population Growth, by Ritchie Calder; B. Benjamin; R. Passmore; N. W. Pirie; Colin Clark; Brinley Thomas. Biometrika, 51(3/4):533, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334174.

Cormack:1964:ESS

[Cor64b] R. M. Cormack. Estimates of survival from the sighting of marked animals. Biometrika, 51(3/4):429–438, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334149.

Cox:1960:NRR

[Cox60] D. R. Cox. On the number of renewals in a random interval. Biometrika, 47(3/4):449–452, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333316.

Cox:1961:BRB

[Cox61] D. R. Cox. Book review: Mathematical Methods in the Theory of Queueing by A. Y. Khinchine. Biometrika, 48(1/2):231–232, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333161.

Cox:1966:NDF

[Cox66a] D. R. Cox. The null distribution of the first serial correlation coefficient. Biometrika, 53(3/4):623–626, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333676. See corrigenda [Cox67c].

Cox:1966:SEC

[Cox66b] D. R. Cox. A simple example of a comparison involving quantal data. Biometrika, 53(1/2):215–220, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334067.

Cox:1967:BRBa

[Cox67a] D. R. Cox. Book review: An Introduction to Stochastic Processes with Special Reference to Methods and Applications, by M. S. Bartlett. Biometrika, 54(1/2):335, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333887.

Cox:1967:BRBb

[Cox67b] D. R. Cox. Book review: The Advanced Theory of Statistics, by M. G. Kendall; A. Stuart. Biometrika, 54(3/4):689–690, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335072.

Cox:1967:CND

[Cox67c] D. R. Cox. Correction: “The Null Distribution of the First Serial Correlation Coefficient”. Biometrika, 54(3/4):688, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335071. See [Cox66a].

Cox:1967:FTG

[Cox67d] D. R. Cox. Fieller’s theorem and a generalization. Biometrika, 54 (3/4):567–572, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335047.

Cox:1992:CWJ

[Cox92] D. R. Cox, editor. The collected works of John W. Tukey. Volume VII. Factorial & ANOVA (1949–1962). Wadsworth & Brooks/Cole Statistics/Probability Series. Wadsworth & Brooks/Cole Advanced Books & Software, Monterey, CA, USA, 1992. ISBN 0-534-05104-9. lxvi + 260 + I6 pp. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Craig:1962:MVS

[Cra62] C. C. Craig. On the mean and variance of the smaller of two drawings from a binomial population. Biometrika, 49(3/4):566–569, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333993.

Crouse:1960:BRB

[Cro60] C. F. Crouse. Book review: Testing Statistical Hypothesis, by E. L. Lehmann. Biometrika, 47(1/2):210, June 1960. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332982.

Crouse:1961:NNR

[Cro61] C. F. Crouse. A non-null ranking model for a sequence of m alternatives. Biometrika, 48(3/4):441–444, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332766.

Crouse:1964:PAP

[Cro64] C. F. Crouse. On a point arising in polynomial regression fitting. Biometrika, 51(3/4):501–503, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334160.

Crouse:1966:DFT

[Cro66] C. F. Crouse. Distribution free tests based on the sample distribution function. Biometrika, 53(1/2):99–108, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334056. See corrigenda [Cro67].

Crouse:1967:CDF

[Cro67] C. F. Crouse. Correction: “Distribution Free Tests Based on the Sample Distribution Function”. Biometrika, 54(1/2):333, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333886. See [Cro66].

Cronholm:1968:TTC

[Cro68] James N. Cronholm. Two tables connected with goodness-of-fit tests for equiprobable alternatives. Biometrika, 55(2):441, July 1968. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334894.

Clunies-Ross:1960:GLD

[CRR60] C. W. Clunies-Ross and R. H. Riffenburgh. Geometry and linear discrimination. Biometrika, 47(1/2):185–189, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332972.

Cunningham:1969:NEV

[Cun69] E. P. Cunningham. A note on the estimation of variance components by the method of fitting constants. Biometrika, 56(3):683–684, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334676.

Curnow:1960:RTV

[Cur60] R. N. Curnow. The regression of true value on estimated value. Biometrika, 47(3/4):457–460, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333318.

Curnow:1966:BRB

[Cur66] R. N. Curnow. Book review: Biological Statistics: An Introduction, by S. C. Pearce. Biometrika, 53(1/2):300–301, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334097.

Daniels:1962:IRS

[Dan62a] H. E. Daniels. An inequality relating the spectral density and autocorrelation function. Biometrika, 49(1/2):262–263, June 1962. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333492.

Daniels:1962:PGP

[Dan62b] H. E. Daniels. Processes generating permutation expansions. Biometrika, 49(1/2):139–149, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333474.

Daniels:1969:RRT

[Dan69] H. E. Daniels. Round-robin tournament scores. Biometrika, 56 (2):295–299, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2334422. See correction [Dan70].

Daniels:1970:CRR

[Dan70] H. E. Daniels. Correction: “Round-Robin Tournament Scores”. Biometrika, 57(2):467, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334862. See [Dan69].

Darroch:1961:TSC

[Dar61] J. N. Darroch. The two-sample capture-recapture census when tagging and sampling are stratified. Biometrika, 48(3/4):241–260, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332748.

David:1960:BRBf

[Dav60a] F. D. David. Book review: Methods of Correlation and Regression Analysis, by M. Ezekiel; K. A. Fox. Biometrika, 47(3/4):487, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333331.

David:1960:BRBh

[Dav60b] F. N. David. Book review: Introduction to Probability and Statistics, by G. W. McElgrath; B. W. Lindgren. Biometrika, 47(3/4):488, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333333.

David:1960:BRBg

[Dav60c] F. N. David. Book review: Modern Statistical Methods, by P. O. Johnson; R. W. B. Jackson. Biometrika, 47(3/4):487–488, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333332.

David:1960:BRBe

[Dav60d] F. N. David. Book review: Probability and Statistics, by Ulf Grenander. Biometrika, 47(3/4):485, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333327.

David:1960:BRBi

[Dav60e] F. N. David. Book review: Rapid Statistical Calculations, by M. H. Quenouille. Biometrika, 47(3/4):488, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333334.

David:1960:BRBb

[Dav60f] F. N. David. Book review: Statistical Estimates and Transformed Beta-Variables, by Gunnar Blom. Biometrika, 47(1/2):210–211, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332984.

David:1960:BRBj

[Dav60g] F. N. David. Book review: Statistical Methods in Biology, by N. T. J. Bailey. Biometrika, 47(3/4):489, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333336.

David:1960:BRBc

[Dav60h] F. N. David. Book review: Statistics: An Introduction, by D. A. S. Fraser. Biometrika, 47(1/2):211, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332985.

David:1960:BRBa

[Dav60i] F. N. David. Book review: Statistics of Extremes, by E. J. Gumbel. Biometrika, 47(1/2):209, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332980.

David:1960:BRBd

[Dav60j] F. N. David. Book review: The Structure of Arithmetic and Algebra, by M. H. Maria. Biometrika, 47(1/2):214, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332990.

David:1961:BRBb

[Dav61a] F. N. David. Book review: An Introduction to Mathematical Statistics, by H. D. Brunk. Biometrika, 48(1/2):235, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333167.

David:1961:BRBc

[Dav61b] F. N. David. Book review: An Introduction to Mathematical Statistics, by R. V. Hogg; A. T. Graig. Biometrika, 48(1/2):235, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333168.

David:1961:BRBj

[Dav61c] F. N. David. Book review: Fundamentals of College Algebra, by W. H. Durfee. Biometrika, 48(3/4):484, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332800.

David:1961:BRBa

[Dav61d] F. N. David. Book review: Principles and Procedures of Statistics, by R. G. D. Steel; J. H. Torrie. Biometrika, 48(1/2):234, June 1961.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333165.

David:1961:BRBf

[Dav61e] F. N. David. Book review: Pronterari Per Calcoli Statistici: Tavole Numeriche e Complementi, by Silvio Vianelli. Biometrika, 48(1/2): 238, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333175.

David:1961:BRBg

[Dav61f] F. N. David. Book review: Sampling Methods for Censuses and Surveys, by F. Yates. Biometrika, 48(3/4):481, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332792.

David:1961:BRBd

[Dav61g] F. N. David. Book review: Special Functions, by Earl D. Rainville. Biometrika, 48(1/2):237, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333172.

David:1961:BRBi

[Dav61h] F. N. David. Book review: The Calculus of Finite Differences, by George Boole. Biometrika, 48(3/4):484, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332799.

David:1961:BRBh

[Dav61i] F. N. David. Book review: The Elliptic Functions as They Should be, by Albert Eagle. Biometrika, 48(3/4):483–484, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332797.

David:1961:BRBe

[Dav61j] F. N. David. Book review: Theory and Solution of Ordinary Differential Equations, by Donald Greenspan. Biometrika, 48(1/2):237– 238, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333174.

David:1962:BRBf

[Dav62a] F. N. David. Book review: An Introduction to the Mathematics of Medicine and Biology by J. G. Defares; I. N. Sneddon. Biometrika, 49(3/4):579–580, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334002.

David:1962:BRBe

[Dav62b] F. N. David. Book review: Introduction to Probability and Statistics by H. L. Alder; E. B. Roessler. Biometrika, 49(3/4):579, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334000.

David:1962:BRBc

[Dav62c] F. N. David. Book review: Mathematical Thinking in the Measurement of Behaviour, by James S. Coleman; Ernest W. Adams; Herbert Solomon. Biometrika, 49(1/2):291, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333511.

David:1962:BRBb

[Dav62d] F. N. David. Book review: Proceedings of the Fourth Berkeley Symposium: Volume IV. Contributions to Biology and Problems of Medicine. by Jerzy Neyman. Biometrika, 49(1/2):286–287, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333505.

David:1962:BRBd

[Dav62e] F. N. David. Book review: Sequential Medical Trails, by P. Armitage. Biometrika, 49(1/2):292, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333513.

David:1962:BRBa

[Dav62f] F. N. David. Book review: The Advanced Theory of Statistics. Vol II. Inference and Relationship by M. G. Kendall; A. Stuart. Biometrika, 49(1/2):285, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333503.

David:1963:BRBf

[Dav63a] F. N. David. Book review: An Index of Mathematical Tables, by A. Fletcher; J. C. P. Miller; L. Rosenhead; L. J. Comrie. Biometrika, 50(3/4):547, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333933.

David:1963:BRBb

[Dav63b] F. N. David. Book review: Elements of Mathematical Statistics by J. F. Ratcliffe. Biometrika, 50(1/2):228, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333773.

David:1963:BRBc

[Dav63c] F. N. David. Book review: Guide to Tables in Mathematical Statistics by J. A. Greenwood; H. O. Hartley. Biometrika, 50(1/2):231, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333778.

David:1963:BRBd

[Dav63d] F. N. David. Book review: Handbook of Statistical Tables by D. B. Owen. Biometrika, 50(1/2):231–232, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333779.

David:1963:BRBa

[Dav63e] F. N. David. Book review: Proceedings of the Fourth Berkeley Symposium on Mathematical Statistics and Probability. Vol. I. Theory of Statistics; Proceedings of the Fourth Berkeley Symposium on Mathematical Statistics and Probability. Vol. II. Probability Theory; Proceedings of the Fourth Berkeley Symposium on Mathematical Statistics and Probability. Vol. III. Astronomy and Physics by Neyman Jerzy. Biometrika, 50(1/2):226–227, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333771.

David:1963:BRBe

[Dav63f] F. N. David. Book review: Tables of Normal and Log-Normal Random Deviates. Biometrika, 50(1/2):232, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333780.

David:1964:BRBa

[Dav64a] F. N. David. Book review: An Introduction to Sampling Theory (with Applications to Agriculture), by M. R. Sampford. Biometrika, 51(1/2):291, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334230.

David:1964:BRBb

[Dav64b] F. N. David. Book review: Bibliography of Nonparametric Statistics, by I. Richard Savage. Biometrika, 51(1/2):292, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334233.

David:1964:BRBd

[Dav64c] F. N. David. Book review: Contributions to Order Statistics, by A. E. Sarhan; B. G. Greenberg. Biometrika, 51(1/2):294, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334236.

David:1964:BRBf

[Dav64d] F. N. David. Book review: Geometrical Probability, by M. G. Kendall; P. A. P. Moran. Biometrika, 51(3/4):532, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334172.

David:1964:BRBc

[Dav64e] F. N. David. Book review: Handbook of Nonparametric Statistics, by John E. Walsh. Biometrika, 51(1/2):293, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334234.

David:1964:BRBg

[Dav64f] F. N. David. Book review: Introduction to the Theory of Statistics, by A. M. Mood; F. A. Graybill. Biometrika, 51(3/4):534, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334176.

David:1964:BRBh

[Dav64g] F. N. David. Book review: The Calculus of Linguistic Observations, by G. Herdan. Biometrika, 51(3/4):534, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334177.

David:1964:BRBe

[Dav64h] F. N. David. Book review: The Case Against Decimalisation, by A. C. Aitken. Biometrika, 51(1/2):295, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334239.

Davis:1964:PGF

[Dav64i] A. W. Davis. On the probability generating functional for the cumulative population in a simple birth-and-death process. Biometrika, 51(1/2):245–249, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334211.

David:1965:BRBa

[Dav65a] F. N. David. Book review: Human Biology: An Introduction to Human Evolution, Variation and Growth, by G. A. Harrison; J. S. Weiner; J. M. Tanner; N. A. Barnicot. Biometrika, 52(3/4):673, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333736.

David:1965:BRBb

[Dav65b] F. N. David. Book review: What Do Statistics Shows? by F. R. Oliver. Biometrika, 52(3/4):674, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333737.

David:1968:GMD

[Dav68] H. A. David. Gini’s mean difference rediscovered. Biometrika, 55(3):573–575, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334264.

Davies:1969:CVD

[Dav69] P. Davies. The choice of variables in the design of experiments for linear regression. Biometrika, 56(1):55–63, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334690.

Day:1969:CSS

[Day69a] N. E. Day. A comparison of some sequential designs. Biometrika, 56(2):301–311, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334423.

Day:1969:ECM

[Day69b] N. E. Day. Estimating the components of a mixture of normal distributions. Biometrika, 56(3):463–474, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334652.

Davidson:1969:MPC

[DB69] Roger R. Davidson and Ralph A. Bradley. Multivariate paired comparisons: The extension of a univariate model and associated estimation and test procedures. Biometrika, 56(1):81–95, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334693. See cor-

rections [DB70].

Davidson:1970:CMP

[DB70] Roger R. Davidson and Ralph A. Bradley. Corrections: “Multivariate Paired Comparisons: The Extension of a Univariate Model and Associated Estimation and Test Procedures”. Biometrika, 57(1): 225, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334961.

See [DB69].

deCani:1969:MLP

[dC69] John S. de Cani. Maximum likelihood paired comparison ranking by linear programming. Biometrika, 56(3):537–545, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334661.

Dempster:1967:ULP

[Dem67] A. P. Dempster. Upper and lower probability inferences based on a sample from a finite univariate population. Biometrika, 54 (3/4):515–528, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335042.

David:1964:IRC

[DF64] F. N. David and Evelyn Fix. Intersections of random chords of a circle. Biometrika, 51(3/4):373–379, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334143.

Draper:1966:UGV

[DG66] Norman R. Draper and Irwin Guttman. Unequal group variances in the fixed-effects one-way analysis of variance: A Bayesian sidelight. Biometrika, 53(1/2):27–35, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334050.

Downs:1967:SRB

[DG67] Thomas D. Downs and A. Lawrence Gould. Some relationships between the normal and von Mises distributions. Biometrika, 54 (3/4):684–687, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335068.

Draper:1968:BST

[DG68a] Norman R. Draper and Irwin Guttman. Bayesian stratified twophase sampling results: k characteristics. Biometrika, 55(3): 587–589, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334270.

Draper:1968:SBS

[DG68b] Norman R. Draper and Irwin Guttman. Some Bayesian stratified two-phase sampling result. Biometrika, 55(1):131–139, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334457.

Draper:1966:DEP

[DH66] Norman R. Draper and William G. Hunter. Design of experiments for parameter estimation in multiresponse situations. Biometrika, 53(3/4):525–533, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333656.

Draper:1967:UPDa

[DH67a] Norman R. Draper and William G. Hunter. The use of prior distributions in the design of experiments for parameter estimation in non-linear situations. Biometrika, 54(1/2):147–153, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333858.

Draper:1967:UPDb

[DH67b] Norman R. Draper and William G. Hunter. The use of prior distributions in the design of experiments for parameter estimation in nonlinear situations: Multiresponse case. Biometrika, 54(3/4):662–665, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335059.

Davis:1969:CCD

[DH69] A. W. Davis and W. B. Hall. Cyclic change-over designs. Biometrika, 56(2):283–293, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334421.

Draper:1965:DWM

[DL65] Norman R. Draper and Willard E. Lawrence. Designs which minimize model inadequacies: Cuboidal regions of interest. Biometrika, 52(1/2):111–118, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333816.

Draper:1966:USO

[DL66] Norman R. Draper and Willard E. Lawrence. The use of secondorder ‘spherical’ and ‘cuboidal’ designs in the wrong regions. Biometrika, 53(3/4):596–599, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333667.

David:1961:VSR

[DM61] F. N. David and C. L. Mallows. The variance of Spearman’s rho in normal samples. Biometrika, 48(1/2):19–28, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333126.

Doss:1962:EBE

[Dos62] S. A. D. C. Doss. On the efficiency of BAN estimates of the parameters of normal populations based on singly censored samples. Biometrika, 49(3/4):570–573, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333994.

Downs:1966:SRA

[Dow66a] Thomas D. Downs. Some relationships among the von Mises distributions of different dimensions. Biometrika, 53(1/2):269–272, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334080.

Downton:1966:LEP

[Dow66b] F. Downton. Linear estimates with polynomial coefficients. Biometrika, 53(1/2):129–141, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334059.

Downton:1967:NUS

[Dow67] F. Downton. A note on the ultimate size of a general stochastic epidemic. Biometrika, 54(1/2):314–316, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333878.

Downton:1968:USC

[Dow68] F. Downton. The ultimate size of carrier-borne epidemics. Biometrika, 55(2):277–289, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334867.

David:1960:CDT

[DP60] H. A. David and Carmen A. Perez. On comparing different tests of the same hypothesis. Biometrika, 47(3/4):297–306, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333301.

David:1965:PST

[DP65] H. A. David and A. S. Paulson. The performance of several tests for outliers. Biometrika, 52(3/4):429–436, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333695.

Daniels:1963:HPS

[DR63] H. E. Daniels and A. S. C. Ross. The Hausa problem and some approximations to the required probability. Biometrika, 50 (3/4):514–517, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333918.

Draper:1964:BRB

[Dra64] J. Draper. Book review: Basic Ideas of Scientific Sampling, by A. Stuart; M. G. Kendall. Biometrika, 51(1/2):290–291, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334229.

Desu:1968:FSS

[DS68] M. Mahamunulu Desu and Milton Sobel. A fixed subset-size approach to the selection problem. Biometrika, 55(2):401–410, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334883. See cor-

rection [DS76].

Desu:1976:CAF

[DS76] M. M. Desu and Milton Sobel. Corrections and amendments: “A fixed subset-size approach to the selection problem” (Biometrika 55 (1968), 401–410). Biometrika, 63(3):685, December 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335756. See [DS68].

Durbin:1961:SMC

[Dur61] J. Durbin. Some methods of constructing exact tests. Biometrika, 48(1/2):41–55, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2333128. See correction [Dur66].

Durbin:1966:CSM

[Dur66] J. Durbin. Correction: “Some Methods of Constructing Exact Tests”. Biometrika, 53(3/4):629, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333678. See [Dur61].

Durbin:1969:TSC

[Dur69] J. Durbin. Tests for serial correlation in regression analysis based on the periodogram of least-squares residuals. Biometrika, 56(1):1–15, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334686.

Weir:1966:TPP

[dVW66] J. B. de V. Weir. Table of 0 · 1 percentage points of Behrens’s d. Biometrika, 53(1/2):267–268, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334079.

Edwards:1961:BRB

[Edw61] D. A. Edwards. Book review: Characteristic Functions, by E. Lukacs. Biometrika, 48(3/4):483, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332796.

Elashoff:1967:CVC

[EEG67] Janet D. Elashoff, R. M. Elashoff, and G. E. Goldman. On the choice of variables in classification problems with dichotomous variables. Biometrika, 54(3/4):668–670, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335061.

Ewens:1962:GSS

[EF62] W. J. Ewens and P. D. Finch. A generalized single-server queue with Erlang input. Biometrika, 49(1/2):242–245, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333486.

Efron:1965:CHR

[Efr65] Bradley Efron. The convex Hull of a random set of points. Biometrika, 52(3/4):331–343, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333687.

Ewan:1960:SIC

[EK60] W. D. Ewan and K. W. Kemp. Sampling inspection of continuous processes with no autocorrelation between successive results. Biometrika, 47(3/4):363–380, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333307.

Elandt:1959:BRB

[Ela59] Regina C. Elandt. Book review: Wahrscheinlichkeitsrechnung und Mathematische Statistik (Probability and Mathematical Statistics), by M. Fisz. Biometrika, 46(3/4):494, December 1959. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333555. See corrigenda

[Ano60k].

Ericson:1967:EDI

[Eri67] W. A. Ericson. An example of discrepancies in inferences under noninformative stopping rules. Biometrika, 54(1/2):329–330, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333883.

Euler:1961:OFD

[Eul61] L. Euler. Observations on the foregoing dissertation of Bernoulli. Biometrika, 48(1/2):13–18, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2333125. Translated from Latin by C. G. Allan.

Ewens:1961:DAS

[Ewe61] W. J. Ewens. Departures from assumption in sequential analysis. Biometrika, 48(1/2):206–211, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333149.

Ewens:1963:NRD

[Ewe63] W. J. Ewens. Numerical results and diffusion approximations in a genetic process. Biometrika, 50(3/4):241–249, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333896.

Fairthorne:1964:DBR

[Fai64] David Fairthorne. The distances between random points in two concentric circles. Biometrika, 51(1/2):275–277, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334223.

Fairweather:1968:SES

[Fai68] William R. Fairweather. Some extensions of Somerville’s procedure for ranking means of normal populations. Biometrika, 55(2):411– 418, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334884.

Farlie:1960:PSC

[Far60] D. J. G. Farlie. The performance of some correlation coefficients for a general bivariate distribution. Biometrika, 47(3/4):307–323, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333302.

Farlie:1963:AED

[Far63] D. J. G. Farlie. The asymptotic efficiency of Daniels’s generalized correlation coefficient. Biometrika, 50(3/4):499–504, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333915.

Feller:1963:BRB

[Fel63] W. Feller. Book review: Combinatorial Chance by F. N. David; D. E. Barton. Biometrika, 50(1/2):225, June 1963. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333770.

Finch:1960:DCI

[Fin60a] P. D. Finch. Deterministic customer impatience in the queueing system GI/M/1. Biometrika, 47(1/2):45–52, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332956. See correction [Fin61].

Finch:1960:CDM

[Fin60b] P. D. Finch. On the covariance determinants of moving-average and autoregressive models. Biometrika, 47(1/2):194–196, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332974.

Finch:1961:DCI

[Fin61] P. D. Finch. Deterministic customer impatience in the queueing system GI/M/1; a correction. Biometrika, 48(3/4):472–473, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332774. See

[Fin60a].

Finch:1962:BRB

[Fin62] P. D. Finch. Book review: An Introduction to Infinitely Many Variables, by Enders A. Robinson. Biometrika, 49(1/2):288, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333507.

Finucan:1964:MMD

[Fin64] H. M. Finucan. The mode of a multinomial distribution. Biometrika, 51(3/4):513–517, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334165.

Folks:1960:EBI

[FK60] John Leroy Folks and Oscar Kempthorne. The efficiency of blocking in incomplete block designs. Biometrika, 47(3/4):273–283, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333299.

Freeman:1967:MRP

[FKM67] Harold Freeman, Arnold Kuzmack, and Rita Maurice. Multivariate t and the ranking problem. Biometrika, 54(1/2):305–308, June 1967.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333874.

Fedorov:1968:EDN

[FMB68] V. D. Fedorov, V. N. Maximov, and V. G. Bogorov. Experimental development of nutritive media for micro-organisms. Biometrika, 55(1):43–51, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334451.

Fieller:1961:TRC

[FP61] E. C. Fieller and E. S. Pearson. Tests for rank correlation coefficients: II. Biometrika, 48(1/2):29–40, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333127.

Fraser:1961:FMI

[Fra61] D. A. S. Fraser. The fiducial method and invariance. Biometrika, 48(3/4):261–280, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332749.

Fraser:1964:FIL

[Fra64] D. A. S. Fraser. Fiducial inference for location and scale parameters. Biometrika, 51(1/2):17–24, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334191.

Fraser:1965:FCG

[Fra65] D. A. S. Fraser. Fiducial consistency and group structure. Biometrika, 52(1/2):55–65, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333811.

Fraser:1966:SPG

[Fra66] D. A. S. Fraser. Structural probability and a generalization. Biometrika, 53(1/2):1–9, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334048.

Gabriel:1968:STP

[Gab68] K. R. Gabriel. Simultaneous test procedures in multivariate analysis of variance. Biometrika, 55(3):489–504, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334253.

Gani:1961:SMG

[Gan61] J. Gani. On the stochastic matrix in a genetic model of Moran. Biometrika, 48(1/2):203–206, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333148.

Gani:1962:EBC

[Gan62] J. Gani. The extinction of a bacterial colony by phages: A branching process with deterministic removals. Biometrika, 49(1/2):272–276, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333496.

Gani:1965:PDE

[Gan65] J. Gani. On a partial differential equation of epidemic theory. I. Biometrika, 52(3/4):617–622, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333711.

Gart:1963:MTS

[Gar63] John J. Gart. A median test with sequential application. Biometrika, 50(1/2):55–62, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333746.

Gart:1964:APR

[Gar64] John J. Gart. The analysis of Poisson regression with an application in virology. Biometrika, 51(3/4):517–521, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334166.

Gart:1969:ETC

[Gar69] John J. Gart. An exact test for comparing matched proportions in crossover designs. Biometrika, 56(1):75–80, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334692.

Grubbs:1966:UPT

[GCP66] Frank E. Grubbs, Helen J. Coon, and E. S. Pearson. On the use of Patnaik type chi approximations to the range in significance tests.

Biometrika, 53(1/2):248–252, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334073.

Geisser:1968:PZM

[GD68] Seymour Geisser and M. M. Desu. Predictive zero-mean uniform discrimination. Biometrika, 55(3):519–524, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334256.

Geary:1966:ACV

[Gea66] R. C. Geary. The average critical value method for adjudging relative efficiency of statistical tests in time series regression analysis. Biometrika, 53(1/2):109–119, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334057.

Gehan:1965:GTS

[Geh65a] Edmund A. Gehan. A generalized two-sample Wilcoxon test for doubly censored data. Biometrika, 52(3/4):650–653, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333721.

Gehan:1965:GWT

[Geh65b] Edmund A. Gehan. A generalized Wilcoxon test for comparing arbitrarily singly-censored samples. Biometrika, 52(1/2):203–223, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333825.

Gupta:1966:EPL

[GG66] Shanti S. Gupta and Mrudulla Gnanadesikan. Estimation of the parameters of the logistic distribution. Biometrika, 53(3/4):565–570, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333660.

Ghirtis:1968:STS

[Ghi68] G. C. Ghirtis. A system of two servers with limited waiting rooms and certain order of visits. Biometrika, 55(1):223–228, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334466.

Ghosh:1963:STR

[Gho63] B. K. Ghosh. On sequential tests of ratio of variances based on range. Biometrika, 50(3/4):419–430, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333910.

Ghosal:1964:SRT

[Gho64a] A. Ghosal. Some results in the theory of inventory. Biometrika, 51(3/4):487–490, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334156.

Ghosh:1964:STS

[Gho64b] B. K. Ghosh. Simultaneous tests by sequential methods in hierarchical classifications. Biometrika, 51(3/4):439–450, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334150.

Ghosh:1966:AEM

[Gho66] B. K. Ghosh. Asymptotic expansions for the moments of the distribution of correlation coefficient. Biometrika, 53(1/2):258–262, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334076.

Gilbert:1960:ELM

[Gil60] N. Gilbert. Estimation from a linear Markov process. Biometrika, 47(3/4):482–483, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333325.

Gilbert:1965:PCS

[Gil65] E. N. Gilbert. The probability of covering a sphere with N circular caps. Biometrika, 52(3/4):323–330, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333686.

Gentleman:1968:ASD

[GJ68] W. M. Gentleman and M. A. Jenkins. An approximation for Student’s t-distribution. Biometrika, 55(3):571–572, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334263.

Glasser:1962:EVC

[Gla62] Gerald J. Glasser. On estimators for variances and covariances. Biometrika, 49(1/2):259–262, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333491.

Glenn:1960:CET

[Gle60] W. A. Glenn. A comparison of the effectiveness of tournaments. Biometrika, 47(3/4):253–262, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333297.

Gleser:1968:TSC

[Gle68] Leon Jay Gleser. On testing a set of correlation coefficients for equality: Some asymptotic results. Biometrika, 55(3):513–517, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334255.

Gurland:1962:TEM

[GM62] John Gurland and Roger S. McCullough. Testing equality of means after a preliminary test of equality of variances. Biometrika, 49 (3/4):403–417, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333975.

Gray:1966:SPR

[GO66] H. L. Gray and P. L. Odell. On sums and products of rectangular variates. Biometrika, 53(3/4):615–617, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333673.

Gobel:1967:BRB

[Gob67] F. Gobel. Book review: Symmetric Function and Allied Tables, by F. N. David; M. G. Kendall; D. E. Barton. Biometrika, 54(1/2):335– 336, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333888.

Gonin:1961:CUO

[Gon61a] H. T. Gonin. Corrigenda: “The Use of Orthogonal Polynomials of Positive and Negative Binomial Frequency Functions in Curve Fitting by Aitken’s Method”. Biometrika, 48(3/4):476, December 1961.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332781. See [Gon61b].

Gonin:1961:UOP

[Gon61b] H. T. Gonin. The use of orthogonal polynomials of the positive and negative binomial frequency functions in curve fitting by Aitken’s method. Biometrika, 48(1/2):115–123, June 1961. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333136. See corrigenda

[Gon61a].

Gonin:1966:PBF

[Gon66] H. T. Gonin. Poisson and binomial frequency surfaces. Biometrika, 53(3/4):617–619, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333674.

Goodman:1964:APC

[Goo64] Leo A. Goodman. The analysis of persistence in a chain of multiple events. Biometrika, 51(3/4):405–411, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334147.

Goodman:1967:PEB

[Goo67] Leo A. Goodman. The probabilities of extinction for birth-and-death processes that are age-dependent or phase-dependent. Biometrika, 54(3/4):579–596, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335049.

Goodman:1968:SMP

[Goo68] Leo A. Goodman. Stochastic models for the population growth of the sexes. Biometrika, 55(3):469–487, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334252.

Good:1969:CQF

[Goo69] I. J. Good. Conditions for a quadratic form to have a chi-squared distribution. Biometrika, 56(1):215–216, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334707. See correction [Goo70]

and remark [Kha78].

Good:1970:CCQ

[Goo70] I. J. Good. Corrections: “Conditions for a Quadratic Form to Have a Chi-Squared Distribution”. Biometrika, 57(1):225, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334962. See [Goo69].

Gower:1961:NSA

[Gow61] J. C. Gower. A note on some asymptotic properties of the logarithmic series distribution. Biometrika, 48(1/2):212–215, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333150.

Gower:1966:TCC

[Gow66a] J. C. Gower. A Q-technique for the calculation of canonical variates. Biometrika, 53(3/4):588–590, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333664.

Gower:1966:SDP

[Gow66b] J. C. Gower. Some distance properties of latent root and vector methods used in multivariate analysis. Biometrika, 53(3/4):325–338, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333639.

Gower:1968:APV

[Gow68] J. C. Gower. Adding a point to vector diagrams in multivariate analysis. Biometrika, 55(3):582–585, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334268.

Gupta:1965:LFO

[GP65] Shanti S. Gupta and K. C. Sreedharan Pillai. On linear functions of ordered correlated normal random variables. Biometrika, 52(3/4):367–379, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333690.

Gupta:1964:DLF

[GPS64] Shanti S. Gupta, K. C. Sreedharan Pillai, and G. P. Steck. On the distribution of linear functions and ratios of linear functions of ordered correlated normal random variables with emphasis on range. Biometrika, 51(1/2):143–151, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334202.

Groves:1969:NEV

[GR69] Theodore Groves and Thomas Rothenberg. A note on the expected value of an inverse matrix. Biometrika, 56(3):690–691, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334680.

Greig:1967:ERA

[Gre67] Margaret Greig. Extremes in a random assembly. Biometrika, 54(1/2):273–282, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333870.

Gruneberg:1961:BRB

[Gru61] H. Gruneberg. Book review: Genetical Research. Biometrika, 48(1/2):236, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333169.

Gupta:1962:SSCa

[GS62a] Shanti S. Gupta and Milton Sobel. On selecting a subset containing the population with the smallest variance. Biometrika, 49 (3/4):495–507, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333983.

Gupta:1962:SSCb

[GS62b] Shanti S. Gupta and Milton Sobel. On the smallest of several correlated F statistics. Biometrika, 49(3/4):509–523, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333984.

Gregory:1965:NQS

[GS65] G. Gregory and P. R. Satyamurty. A note on the queueing system M/M/1 with balking. Biometrika, 52(3/4):643–645, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333718.

Gehan:1969:PST

[GT69] Edmund A. Gehan and Donald G. Thomas. The performance of some two-sample tests in small samples with and without censoring.

Biometrika, 56(1):127–132, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334697.

Glasser:1961:CVC

[GW61] Gerald J. Glasser and Robert F. Winter. Critical values of the coefficient of rank correlation for testing the hypothesis of independence. Biometrika, 48(3/4):444–448, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332767.

Gart:1967:BVE

[GZ67] John J. Gart and James R. Zweifel. On the bias of various estimators of the logit and its variance with application to quantal bioassay. Biometrika, 54(1/2):181–187, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333861.

Haight:1960:QBI

[Hai60] Frank A. Haight. Queueing with balking. II. Biometrika, 47 (3/4):285–296, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333300.

Haight:1961:DAB

[Hai61] Frank A. Haight. A distribution analogous to the Borel–Tanner. Biometrika, 48(1/2):167–173, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333140.

Haight:1965:CDR

[Hai65a] Frank A. Haight. Counting distributions for renewal processes. Biometrika, 52(3/4):395–403, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333692.

Haight:1965:ERP

[Hai65b] Frank A. Haight. On the effect of removing persons with N or more accidents from an accident prone population. Biometrika, 52(1/2):298–300, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333840.

Hajnal:1961:TSS

[Haj61] J. Hajnal. A two-sample sequential t-test. Biometrika, 48(1/2):65– 75, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333131.

Hall:1962:SSA

[Hal62] W. J. Hall. Some sequential analogs of Stein’s two-stage test. Biometrika, 49(3/4):367–378, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333971.

Haley:1963:BRB

[Hal63] K. B. Haley. Book review: Mathematical Programming by S. Vajda. Biometrika, 50(1/2):230–231, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333776.

Hamdan:1968:STS

[Ham68] M. A. Hamdan. On the structure of the tetrachoric series. Biometrika, 55(1):261–262, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334474.

Hannan:1962:SS

[Han62] E. J. Hannan. Systematic sampling. Biometrika, 49(1/2):281–283, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333499.

Hannan:1963:RTS

[Han63] E. J. Hannan. Regression for time series with errors of measurement. Biometrika, 50(3/4):293–302, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333900.

Hannan:1967:ELR

[Han67] E. J. Hannan. The estimation of a lagged regression relation. Biometrika, 54(3/4):409–418, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335033.

Han:1968:NDC

[Han68] Chien-Pai Han. A note on discrimination in the case of unequal covariance matrices. Biometrika, 55(3):586–587, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334269.

Hannan:1969:EMM

[Han69a] E. J. Hannan. The estimation of mixed moving average autoregressive systems. Biometrika, 56(3):579–593, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334666.

Hannan:1969:IVM

[Han69b] E. J. Hannan. The identification of vector mixed autoregressivemoving average systems. Biometrika, 56(1):223–225, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334711.

Harter:1961:CEV

[Har61a] H. Leon Harter. Corrigenda: “Expected Values of Normal Order Statistics”. Biometrika, 48(3/4):476, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332782. See [Har61b].

Harter:1961:EVN

[Har61b] H. Leon Harter. Expected values of normal order statistics. Biometrika, 48(1/2):151–165, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2333139. See corrigenda [Har61a].

Harary:1963:BRBb

[Har63a] F. Harary. Book review: Programmes, Jeux et Reseaux de Transport, by C. Berge; A. Ghouila-Houri. Biometrika, 50(3/4):553–554, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333941.

Harary:1963:BRBa

[Har63b] F. Harary. Book review: The Theory of Graphs and its Applications by Claude Berge. Biometrika, 50(1/2):231, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333777.

Harter:1963:PPR

[Har63c] H. Leon Harter. Percentage points of the ratio of two ranges and power of the associated test. Biometrika, 50(1/2):187–194, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333759.

Harter:1964:NTP

[Har64a] H. Leon Harter. A new table of percentage points of the chisquare distribution. Biometrika, 51(1/2):231–239, June 1964. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334209. See corrigenda

[Har65, Har79].

Hartley:1964:ECR

[Har64b] H. O. Hartley. Exact confidence regions for the parameters in nonlinear regression laws. Biometrika, 51(3/4):347–353, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334140.

Harter:1965:CNT

[Har65] H. Leon Harter. Corrigenda: “A new table of percentage points of the chi-square distribution”. Biometrika, 52(1/2):305, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333844. See [Har64a].

Harper:1967:DMH

[Har67a] W. M. Harper. The distribution of the mean half-square successive difference. Biometrika, 54(3/4):419–433, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335034.

Hartigan:1967:DRS

[Har67b] J. A. Hartigan. Distribution of the residual sum of squares in fitting inequalities. Biometrika, 54(1/2):69–84, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333852.

Harsaae:1969:CUT

[Har69a] Erik Harsaae. On the computation and use of a table of percentage points of Bartlett’s M. Biometrika, 56(2):273–281, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334420.

Harville:1969:QUE

[Har69b] David A. Harville. Quadratic unbiased estimation of variance components for the one-way classification. Biometrika, 56(2):313–326, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334424.

See corrections [Har70].

Harville:1970:CQU

[Har70] David A. Harville. Corrections: “Quadratic Unbiased Estimation of Variance Components for the One-Way Classification”. Biometrika, 57(1):226, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334964. See [Har69b].

Harter:1979:ACN

[Har79] H. L. Harter. Amendments and corrections: “A new table of percentage points of the chi-square distribution” [Biometrika 51 (1964), 231–239; MR 30 #4315]. Biometrika, 66(3):689, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335199. See [Har64a].

Hasofer:1967:SHP

[Has67] A. M. Hasofer. Studies in the history of probability and statistics. XVI. Random mechanisms in Talmudic literature. Biometrika, 54(1/2):316–321, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333879.

Hassanein:1969:EPE

[Has69a] Khatab M. Hassanein. Estimation of the parameters of the extreme value distribution by use of two or three order statistics. Biometrika, 56(2):429–436, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334436.

Hassanein:1969:EPL

[Has69b] Khatab M. Hassanein. Estimation of the parameters of the logistic distribution by sample quantiles. Biometrika, 56(3):684–687, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334677.

Hawkes:1964:BRB

[Haw64] A. G. Hawkes. Book review: Les Systems avec ou sans Attente et les Processus Stochastiques, by P. Le Gall. Biometrika, 51(1/2):292, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334232.

Hawkes:1965:BRB

[Haw65a] A. G. Hawkes. Book review: Mathematical Theories of Traffic Flow, by Frank A. Haight. Biometrika, 52(3/4):672–673, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333734.

Hawkes:1965:QGT

[Haw65b] A. G. Hawkes. Queueing for gaps in traffic. Biometrika, 52(1/2): 79–85, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333813.

Haight:1960:BTD

[HB60] Frank A. Haight and Melvin Allen Breuer. The Borel–Tanner distribution. Biometrika, 47(1/2):143–150, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332966.

Heathcote:1961:PPQ

[Hea61] C. R. Heathcote. Preemptive priority queueing. Biometrika, 48 (1/2):57–63, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333130.

Healy:1962:BRB

[Hea62] M. J. R. Healy. Book review: An Introduction to Statistical Science in Agriculture by D. J. Finney. Biometrika, 49(3/4):579, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334001.

Healy:1964:BRB

[Hea64a] M. J. R. Healy. Book review: Auto-Primer in Computer Programming, by Doris R. Entwistle. Biometrika, 51(1/2):294, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334238.

Healy:1964:PMD

[Hea64b] M. J. R. Healy. A property of the multinomial distribution and the determination of appropriate scores. Biometrika, 51(1/2):265–266, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334219.

Hext:1963:ESO

[Hex63] George R. Hext. The estimation of second-order tensors, with related tests and designs. Biometrika, 50(3/4):353–373, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333905.

Hedayat:1969:AGT

[HF69] A. Hedayat and W. T. Federer. An application of group theory to the existence and non-existence of orthogonal Latin squares. Biometrika, 56(3):547–551, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334662.

Healy:1961:APC

[HG61] M. J. R. Healy and J. C. Gower. Aliasing in partially confounded factorial experiments. Biometrika, 48(1/2):218–220, June 1961. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333152.

Hinz:1967:STE

[HG67] Paul Hinz and John Gurland. Simplified techniques for estimating parameters of some generalized Poisson distributions. Biometrika, 54(3/4):555–566, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335046.

Hinz:1968:MAU

[HG68] Paul Hinz and John Gurland. A method of analysing untransformed data from the negative binomial and other contagious distributions. Biometrika, 55(1):163–170, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334460.

Hills:1969:LLC

[Hil69] M. Hills. On looking at large correlation matrices. Biometrika, 56(2):249–253, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334418.

Hinkley:1969:IAI

[Hin69a] D. V. Hinkley. Inference about the intersection in two-phase regression. Biometrika, 56(3):495–504, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334655.

Hinkley:1969:RTC

[Hin69b] D. V. Hinkley. On the ratio of two correlated normal random variables. Biometrika, 56(3):635–639, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334671. See correction [Hin70].

Hinkley:1970:CRT

[Hin70] D. V. Hinkley. Correction: “On the Ratio of Two Correlated Normal Random Variables”. Biometrika, 57(3):683, December 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334796. See [Hin69b].

Hitchcock:1962:NEP

[Hit62] Shirley E. Hitchcock. A note on the estimation of the parameters of the logistic function, using the minimum logit χ^{2} method. Biometrika, 49(1/2):250–252, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333488.

Hitchcock:1966:THA

[Hit66] Shirley E. Hitchcock. Tests of hypotheses about the parameters of the logistic function. Biometrika, 53(3/4):535–544, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333657.

Haldane:1963:DEN

[HJ63] J. B. S. Haldane and S. D. Jayakar. The distribution of extremal and nearly extremal values in samples from a normal distribution. Biometrika, 50(1/2):89–94, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333750.

Hinkelmann:1963:TCG

[HK63] K. Hinkelmann and O. Kempthorne. Two classes of group divisible partial diallel crosses. Biometrika, 50(3/4):281–291, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333899.

Hudson:1968:MCI

[HK68] John D. Hudson, Jr. and Richard G. Krutchkoff. A Monte Carlo investigation of the size and power of tests employing Satterthwaite’s synthetic mean squares. Biometrika, 55(2):431–433, July 1968. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334890.

Haynam:1965:ACD

[HL65] G. E. Haynam and F. C. Leone. Analysis of categorical data. Biometrika, 52(3/4):654–660, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333723.

Harper:1960:UIT

[HM60] W. M. Harper and J. A. MacDonald. The use of integral transforms to determine expansions of distribution functions. Biometrika, 47 (3/4):460–463, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333319.

Harter:1966:IML

[HM66] H. Leon Harter and Albert H. Moore. Iterative maximum-likelihood estimation of the parameters of normal populations from singly and doubly censored samples. Biometrika, 53(1/2):205–213, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334066. See cor-

rections [HM67, HM69].

Harter:1967:CIM

[HM67] H. Leon Harter and Albert H. Moore. Correction: “Iterative Maximum-Likelihood Estimation of the Parameters of Normal Populations from Singly and Doubly Censored Samples”. Biometrika, 54(3/4):688, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335069. See [HM66].

Harter:1969:CIM

[HM69] H. Leon Harter and Albert H. Moore. Corrections: “Iterative Maximum-Likelihood Estimation of the Parameters of Normal Populations from Singly and Doubly Censored Samples”. Biometrika, 56(1):229, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334714. See [HM66].

Hoadley:1968:UPD

[Hoa68] A. Bruce Hoadley. Use of the Pearson densities for approximating a skew density whose left terminal and first three moments are known. Biometrika, 55(3):559–563, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334261.

Hoel:1968:CST

[Hoe68] David G. Hoel. Closed sequential tests of an exponential parameter. Biometrika, 55(2):387–391, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334881.

Hoflund:1963:SDS

[Hof63] Olle Hoflund. Simulated distributions for small n of Kendall’s partial rank correlation coefficient. Biometrika, 50(3/4):520–522, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333920.

Holland:1961:BRB

[Hol61a] W. W. Holland. Book review: Introduction to Health Statistics, by Satya Swaroop. Biometrika, 48(1/2):236, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333171.

Holt:1961:BRB

[Hol61b] S. B. Holt. Book review: La Biometrie. Que Sais-Je, by Eugene Schreider. Biometrika, 48(1/2):236, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333170.

Holgate:1964:EBP

[Hol64] P. Holgate. Estimation for the bivariate Poisson distribution. Biometrika, 51(1/2):241–245, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334210.

Holgate:1965:DRP

[Hol65a] P. Holgate. The distance from a random point to the nearest point of a closely packed lattice. Biometrika, 52(1/2):261–263, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333828.

Holgate:1965:TRB

[Hol65b] P. Holgate. Tests of randomness based on distance methods. Biometrika, 52(3/4):345–353, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333688.

Holgate:1966:BGN

[Hol66] P. Holgate. Bivariate generalizations of Neyman’s type A distribution. Biometrika, 53(1/2):241–245, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334071.

Holgate:1967:ACM

[Hol67] P. Holgate. The angle-count method. Biometrika, 54(3/4):615–623, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335052.

Holgate:1969:SFD

[Hol69] P. Holgate. Species frequency distributions. Biometrika, 56(3): 651–660, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334673.

Hogben:1961:MNC

[HPW61] D. Hogben, R. S. Pinkham, and M. B. Wilk. The moments of the non-central t-distribution. Biometrika, 48(3/4):465–468, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332772.

Hartley:1967:MLE

[HR67] H. O. Hartley and J. N. K. Rao. Maximum-likelihood estimation for the mixed analysis of variance model. Biometrika, 54(1/2):93–108, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333854.

Hartley:1968:NET

[HR68] H. O. Hartley and J. N. K. Rao. A new estimation theory for sample surveys. Biometrika, 55(3):547–557, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334260.

Hill:1969:SSE

[HS69] Robert T. Hill and Norman C. Severo. The simple stochastic epidemic for small populations with one or more initial infectives. Biometrika, 56(1):183–196, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334703.

Hodson:1966:SEN

[HSD66] F. R. Hodson, P. H. A. Sneath, and J. E. Doran. Some experiments in the numerical analysis of archaeological data. Biometrika, 53 (3/4):311–324, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333638.

Healy:1962:TPL

[HT62] M. J. R. Healy and L. R. Taylor. Tables for power-law transformations. Biometrika, 49(3/4):557–559, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333988.

Hannan:1965:EPM

[HT65] J. F. Hannan and R. F. Tate. Estimation of the parameters for a multivariate normal distribution when one variable is dichotomized. Biometrika, 52(3/4):664–668, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333725.

Hanumara:1968:PPE

[HT68] R. Choudary Hanumara and W. A. Thompson, Jr. Percentage points of the extreme roots of a Wishart matrix. Biometrika, 55(3):505–512, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334254.

Hussain:1969:MMR

[Hus69] Ashiq Hussain. A mixed model for regressions. Biometrika, 56(2):327–336, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334425.

Ireland:1968:CTG

[IK68] C. T. Ireland and S. Kullback. Contingency tables with given marginals. Biometrika, 55(1):179–188, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334462.

Imhof:1961:CDQ

[Imh61] J. P. Imhof. Computing the distribution of quadratic forms in normal variables. Biometrika, 48(3/4):419–426, December 1961. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332763. See corrigenda

[Imh62].

Imhof:1962:CCD

[Imh62] J. P. Imhof. Corrigenda: “Computing the Distribution of Quadratic Forms in Normal Variables”. Biometrika, 49(1/2):284, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333500. See [Imh61].

Iglewicz:1968:PPS

[IMH68] Boris Iglewicz, Raymond H. Myers, and Richard B. Howe. On the percentage points of the sample coefficient of variation. Biometrika, 55(3):580–581, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334267.

Ito:1964:RTM

[IS64] Koichi Ito and William J. Schull. On the robustness of the T_{0}^{2} test in multivariate analysis of variance when variance-covariance matrices are not equal. Biometrika, 51(1/2):71–82, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334197.

Ito:1962:CPT

[Ito62] Koichi Ito. A comparison of the powers of two multivariate analysis of variance tests. Biometrika, 49(3/4):455–462, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333979. See corri-

genda [Ito63].

Ito:1963:CCP

[Ito63] Koichi Ito. Corrigenda: “A Comparison of the Powers of Two Multivariate Analysis of Variance Tests”. Biometrika, 50(3/4):546, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333930.

See [Ito62].

Jackson:1968:SRT

[Jac68] O. A. Y. Jackson. Some results on tests of separate families of hypotheses. Biometrika, 55(2):355–363, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334876.

Jagers:1967:IBP

[Jag67] Peter Jagers. Integrals of branching processes. Biometrika, 54 (1/2):263–271, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333869.

Jones:1967:TSP

[JB67] Richard H. Jones and William M. Brelsford. Time series with periodic structure. Biometrika, 54(3/4):403–408, December 1967. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335032.

Jenkins:1964:BRB

[Jen64] G. M. Jenkins. Book review: Stochastic Service Systems, by J. Riordan. Biometrika, 51(1/2):291–292, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334231.

Johnson:1963:TPP

[JNAP63] N. L. Johnson, Eric Nixon, D. E. Amos, and E. S. Pearson. Table of percentage points of Pearson curves, for given \sqrt{β_{1}} and β_{2}, expressed in standard measure. Biometrika, 50(3/4):459–498, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333914. See

corrigenda [JNAP65].

Johnson:1965:CTP

[JNAP65] N. L. Johnson, Eric Nixon, D. E. Amos, and E. S. Pearson. Corrigenda: “Table of percentage points of Pearson curves, for given √β_{1} and β_{2}, expressed in standard measure”. Biometrika, 52(3/4):669, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333728.

See [JNAP63].

Johnson:1960:AMD

[Joh60a] N. L. Johnson. An approximation to the multinomial distribution: Some properties and applications. Biometrika, 47(1/2):93–102, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332962.

Johnson:1960:BRBb

[Joh60b] N. L. Johnson. Book review: Allgemeine Methodenlehre Der Statistik, by Johann Pfanzagl. Biometrika, 47(3/4):485, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333328.

Johnson:1960:BRBa

[Joh60c] N. L. Johnson. Book review: Die Axiomatischen Grundlagen Einer Allgemeinen Theorie des Messens. (Axiomatic Foundations of the General Theory of Measurement), by J. Pfanzagl. Biometrika, 47(1/2):210, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332983.

John:1961:EPI

[Joh61] S. John. On the evaluation of the probability integral of the multivariate t-distribution. Biometrika, 48(3/4):409–417, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332762.

Johnson:1962:BRBe

[Joh62a] N. L. Johnson. Book review: Allgemeine Methodenlehre Der Statistik by Johann Pfanzagl. Biometrika, 49(3/4):578, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333998.

Johnson:1962:BRBf

[Joh62b] N. L. Johnson. Book review: Lehrbuch Der Praktischen Statistik by Peter Quante. Biometrika, 49(3/4):580, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334003.

Johnson:1962:BRBc

[Joh62c] N. L. Johnson. Book review: Tables of Bivariate Osculatory Interpolation Over a Cartesian Grid, by Herbert E. Salzer; Genevieve M. Kimbro. Biometrika, 49(1/2):295, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333519.

Johnson:1962:BRBd

[Joh62d] N. L. Johnson. Book review: Tables of Coefficients for Obtaining the Second Derivative with Differences by Herbert E. Salzer; Peggy T. Roberson. Biometrika, 49(1/2):295, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333520.

Johnson:1962:BRBb

[Joh62e] N. L. Johnson. Book review: Tables of Osculatory Integration Coefficients, by Herbert E. Salzer; Dexter C. Shoultz; Elizabeth P. Thompson. Biometrika, 49(1/2):294, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333518.

Johnson:1962:BRBa

[Joh62f] N. L. Johnson. Book review: Tables of Osculatory Interpolation Coefficients, by H. E. Salzer. Biometrika, 49(1/2):294, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333517.

Johnson:1964:BRB

[Joh64] N. L. Johnson. Book review: Factors for One-Sided Tolerance Limits and for Variables Sampling Plans, by D. B. Owen. Biometrika, 51(1/2):293–294, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334235.

John:1965:NAI

[Joh65a] J. A. John. A note on the analysis of incomplete block experiments. Biometrika, 52(3/4):633–636, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333715.

Johnson:1965:TFF

[Joh65b] N. L. Johnson. Tables to facilitate fitting S_{U} frequency curves. Biometrika, 52(3/4):547–558, December 1965. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333704.

Johnson:1966:BRB

[Joh66] N. L. Johnson. Book review: Statistical Distribution Programs for a Computer Language and Noncentral Statistical Distribution Programs for a Computer Language, by Rolf E. Bargmann; Sakti P. Ghosh. Biometrika, 53(1/2):306, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334105.

John:1969:MRP

[Joh69] S. John. On multivariate ratio and product estimators. Biometrika, 56(3):533–536, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334660.

Jolly:1963:EPP

[Jol63] G. M. Jolly. Estimates of population parameters from multiple recapture data with both death and dilution—deterministic model. Biometrika, 50(1/2):113–128, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333752.

Jolly:1965:EEC

[Jol65] G. M. Jolly. Explicit estimates from capture-recapture data with both death and immigration-stochastic model. Biometrika, 52 (1/2):225–247, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333826.

Johnson:1969:TPP

[JP69] N. L. Johnson and E. S. Pearson. Tables of percentage points of non-central χ. Biometrika, 56(2):255–272, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334419. See correction [JP70].

Johnson:1970:CTP

[JP70] N. L. Johnson and E. S. Pearson. Correction: “Tables of Percentage Points of Noncentral χ”. Biometrika, 57(2):467, August 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334861. See [JP69].

Johnson:1965:NDC

[JW65] N. L. Johnson and B. L. Welch. Note: The distribution of the coefficient of variation; comment on a criticism made by Koopmans, Owen and Rosenblatt. Biometrika, 52(1/2):303, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333842.

Johnson:1960:SAT

[JY60] N. L. Johnson and D. H. Young. Some applications of two approximations to the multinomial distribution. Biometrika, 47 (3/4):463–469, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333320.

Kabe:1964:DWD

[Kab64] D. G. Kabe. Decomposition of Wishart distribution. Biometrika, 51(1/2):267, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334220.

Kale:1961:SLE

[Kal61a] B. K. Kale. On the solution of the likelihood equation by iteration processes. Biometrika, 48(3/4):452–456, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332769. See corrigenda [Kal62a].

Kalmus:1961:BRBa

[Kal61b] H. Kalmus. Book review: Statistische Methoden der Populationsgenetik, by Henri Louis le Roy. Biometrika, 48(3/4):482, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332793.

Kalmus:1961:BRBb

[Kal61c] H. Kalmus. Book review: Symposia of the Society for Experimental Biology, by J. W. L. Beament. Biometrika, 48(3/4):482, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332794.

Kale:1962:CSL

[Kal62a] B. K. Kale. Corrigenda: “On the Solution of the Likelihood Equation by Iteration Processes”. Biometrika, 49(1/2):284, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333501. See [Kal61a].

Kale:1962:NPE

[Kal62b] B. K. Kale. A note on a problem in estimation. Biometrika, 49 (3/4):553–557, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333987.

Kale:1962:SLE

[Kal62c] B. K. Kale. On the solution of likelihood equations by iteration processes. The multiparametric case. Biometrika, 49(3/4):479–486, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333981.

Kale:1964:NLI

[Kal64] B. K. Kale. A note on the loss of information due to grouping of observations. Biometrika, 51(3/4):495–497, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334158.

Kale:1966:AML

[Kal66] B. K. Kale. Approximations to the maximum-likelihood estimator using grouped data. Biometrika, 53(1/2):282–285, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334085.

Kamat:1965:PMD

[Kam65] A. R. Kamat. A property of the mean deviation for a class of continuous distributions. Biometrika, 52(1/2):288–289, June 1965. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333837.

Kamat:1966:GJP

[Kam66a] A. R. Kamat. A generalization of Johnson’s property of the mean deviation for a class of discrete distributions. Biometrika, 53(1/2):285– 287, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334086.

Kamat:1966:PMD

[Kam66b] A. R. Kamat. A property of the mean deviation for the Pearson type distributions. Biometrika, 53(1/2):287–289, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334087.

Kamat:1967:AP

[Kam67] A. R. Kamat. Acknowledgement of priority. Biometrika, 54(1/2): 333, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333885.

Kastenbaum:1960:SMC

[Kas60] Marvin A. Kastenbaum. The separation of molecular compounds by countercurrent dialysis: A stochastic process. Biometrika, 47(1/2): 69–77, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332959.

Kempthorne:1969:BSS

[KD69] Oscar Kempthorne and T. E. Doerfler. The behaviour of some significance tests under experimental randomization. Biometrika, 56(2):231–248, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334417.

Keller:1962:FML

[Kel62] Joseph B. Keller. Factorization of matrices by least-squares. Biometrika, 49(1/2):239–242, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333485.

Kendall:1960:BDP

[Ken60a] David G. Kendall. Birth-and-death processes, and the theory of carcinogenesis. Biometrika, 47(1/2):13–21, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332953.

Kendall:1960:SHP

[Ken60b] M. G. Kendall. Studies in the history of probability and statistics. Where shall the history of statistics begin? Biometrika, 47 (3/4):447–449, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333315.

Kendall:1961:SHPa

[Ken61a] M. G. Kendall. Studies in the history of probability and statistics: XI. Daniel Bernoulli on maximum likelihood. Biometrika, 48(1/2): 1–2, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333125.

Kendall:1961:SHPb

[Ken61b] M. G. Kendall. Studies in the history of probability and statistics. XII. The book of fate. Biometrika, 48(1/2):220–222, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333153.

Kendall:1961:TTA

[Ken61c] M. G. Kendall. A theorem in trend analysis. Biometrika, 48 (1/2):224–227, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333155.

Kendall:1962:RM

[Ken62] M. G. Kendall. Ranks and measures. Biometrika, 49(1/2):133–137, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333473.

Kendall:1963:RAF

[Ken63a] M. G. Kendall. Ronald Aylmer Fisher, 1890–1962. Biometrika, 50 (1/2):1–15 (1 plate), June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333741.

Kendall:1963:SHP

[Ken63b] M. G. Kendall. Studies in the history of probability and statistics. XIII. Isaac Todhunter’s History of the Mathematical Theory of Probability. Biometrika, 50(1/2):204–205, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333762.

Kendall:1964:BRB

[Ken64] D. G. Kendall. Book review: Journal of Applied Probability. Biometrika, 51(3/4):529, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334169.

Kendall:1966:BRBa

[Ken66a] D. G. Kendall. Book review: Processus Stochastiques et Mouvement Brownien, by Paul Levy; Diffusion Processes and their Sample Paths. by Kiyosi Ito; Henry P. McKean,; Diffusion Models in Population Genetics. by Motoo Kimura. Biometrika, 53(1/2):293–294, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334089.

Kendall:1966:BRBb

[Ken66b] M. G. Kendall. Book review: Bibliography on Time Series and Stochastic Processes, by Herman Wold. Biometrika, 53(1/2):297– 298, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334093.

Kendall:1968:SHPb

[Ken68a] M. G. Kendall. Studies in the history of probability and statistics. XIX. Francis Ysidro Edgeworth (1845–1926). Biometrika, 55 (2):269–275. (1 plate), July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334866.

Kendall:1968:SHPa

[Ken68b] M. G. Kendall. Studies in the history of probability and statistics. XVIII. Thomas Young on coincidences. Biometrika, 55(1):249–250, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334469.

Kerridge:1964:PSS

[Ker64] D. Kerridge. Probabilistic solution of the simple birth process. Biometrika, 51(1/2):258–259, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334215.

Kestelman:1961:BRB

[Kes61] H. Kestelman. Book review: The Theory of Matrices, by F. R. Gantmacher. Biometrika, 48(1/2):237, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333173.

Katti:1962:ECM

[KG62] S. K. Katti and John Gurland. Efficiency of certain methods of estimation for the negative binomial and the Neyman Type A distributions. Biometrika, 49(1/2):215–226, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333483.

Kshirsagar:1967:AVF

[KG67] A. M. Kshirsagar and Y. P. Gupta. Asymptotic values of the first two moments in Markov renewal processes. Biometrika, 54 (3/4):597–603, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335050.

Khatri:1978:RNS

[Kha78] C. G. Khatri. A remark on the necessary and sufficient conditions for a quadratic form to be distributed as chi-squared. Biometrika, 65(1):239–240, April 1978. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335305. See [Goo69].

Kingman:1961:EBR

[Kin61] J. F. C. Kingman. The ergodic behaviour of random walks. Biometrika, 48(3/4):391–396, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332760.

Kingman:1962:SIQ

[Kin62] J. F. C. Kingman. Some inequalities for the queue GI/G/1. Biometrika, 49(3/4):315–324, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333966.

King:1964:ORS

[Kin64] E. P. King. Optimal replication in sequential drug screening. Biometrika, 51(1/2):1–10, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334189.

Kemp:1965:SPH

[KK65] C. D. Kemp and Adrienne W. Kemp. Some properties of the ‘Hermite’ distribution. Biometrika, 52(3/4):381–394, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333691.

Kemp:1966:ADH

[KK66] Adrienne W. Kemp and C. D. Kemp. An alternative derivation of the Hermite distribution. Biometrika, 53(3/4):627–628, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333677.

Kurtz:1965:SCMb

[KLTW65] T. E. Kurtz, R. F. Link, J. W. Tukey, and D. L. Wallace. Shortcut multiple comparisons for balanced single and double classifications: Part 2. Derivations and approximations. Biometrika, 52 (3/4):485–498, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2333700. Reprinted in [Cox92, pp. 431–452].

Kurtz:1966:CRC

[KLTW66] T. E. Kurtz, R. F. Link, J. W. Tukey, and D. L. Wallace. Correlation of ranges of correlated deviates. Biometrika, 53(1/2):191–197, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334064. With

editorial note by Egon S. Pearson.

Khazanie:1966:MMP

[KM66] R. G. Khazanie and H. E. McKean. A Mendelian Markov process with binomial transition probabilities. Biometrika, 53(1/2):37–48, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334051.

Koopmans:1964:CIC

[KOR64] L. H. Koopmans, D. B. Owen, and J. I. Rosenblatt. Confidence intervals for the coefficient of variation for the normal and log normal distributions. Biometrika, 51(1/2):25–32, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334192.

Korin:1968:DSU

[Kor68] Basil P. Korin. On the distribution of a statistic used for testing a covariance matrix. Biometrika, 55(1):171–178, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334461.

Korin:1969:TEC

[Kor69] B. P. Korin. On testing the equality of k covariance matrices. Biometrika, 56(1):216–218, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334708.

Kotlarski:1965:PIR

[Kot65] Ignacy Kotlarski. On pairs of independent random variables whose product follows the gamma distribution. Biometrika, 52(1/2):289– 294, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333838.

Kotlarski:1966:CND

[Kot66] Ignacy Kotlarski. On characterizing the normal distribution by Student’s law. Biometrika, 53(3/4):603–606, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333669.

Krishnaji:1967:JDA

[Kri67] N. Krishnaji. The joint distribution of ascending pairs and ascending runs in a random sequence. Biometrika, 54(1/2):330–332, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333884.

Krutchkoff:1967:SSN

[Kru67] Richard G. Krutchkoff. A supplementary sample non-parametric empirical Bayes approach to some statistical decision problems. Biometrika, 54(3/4):451–458, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335036.

Kshirsagar:1960:NDS

[Ksh60] A. M. Kshirsagar. A note on the derivation of some exact multivariate tests. Biometrika, 47(3/4):480–482, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333324.

Kshirsagar:1961:GFS

[Ksh61] A. M. Kshirsagar. The goodness-of-fit of a single (non-isotropic) hypothetical principal component. Biometrika, 48(3/4):397–407, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332761.

Kshirsagar:1962:NDC

[Ksh62] A. M. Kshirsagar. A note on direction and collinearity factors in canonical analysis. Biometrika, 49(1/2):255–259, June 1962. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333490.

Kshirsagar:1966:NND

[Ksh66] A. M. Kshirsagar. The non-null distribution of a statistic in principal components analysis. Biometrika, 53(3/4):590–594, December 1966.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333665.

Kowalski:1969:COT

[KT69] Charles J. Kowalski and Michael E. Tarter. Co-ordinate transformations to normality and the power of normal tests for independence. Biometrika, 56(1):139–148, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334699.

Kudo:1963:MAO

[Kud63] Akio Kudô. A multivariate analogue of the one-sided test. Biometrika, 50(3/4):403–418, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333909.

Kapadia:1963:VCT

[KW63] C. H. Kapadia and David L. Weeks. Variance components in two-way classification models with interaction. Biometrika, 50 (3/4):327–334, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333903.

Kurkjian:1963:ACF

[KZ63] B. Kurkjian and M. Zelen. Applications of the calculus of factorial arrangements: I. Block and direct product designs. Biometrika, 50(1/2):63–73, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333747.

Lachenbruch:1969:TSC

[Lac69] Peter A. Lachenbruch. Tables of simultaneous confidence limits for the binomial and Poisson distributions. Biometrika, 56(2):452, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334441.

Lancaster:1966:KRH

[Lan66] H. O. Lancaster. Kolmogorov’s remark on the Hotelling canonical correlations. Biometrika, 53(3/4):585–588, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333663.

Langford:1969:PRT

[Lan69] Eric Langford. The probability that a random triangle is obtuse. Biometrika, 56(3):689–690, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334679.

Larsen:1969:AVT

[Lar69] W. A. Larsen. The analysis of variance for the two-way classification fixed-effects model with observations within a row serially correlated. Biometrika, 56(3):509–515, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334657.

Larson:1963:BPR

[LB63] Harold J. Larson and T. A. Bancroft. Biases in prediction by regression for certain incompletely specified models. Biometrika, 50 (3/4):391–402, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333908.

Lindley:1965:SST

[LB65] D. V. Lindley and B. N. Barnett. Sequential sampling: Two decision problems with linear losses for binomial and normal random variables. Biometrika, 52(3/4):507–532, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333702.

Leslie:1966:URT

[LB66] R. T. Leslie and B. M. Brown. Use of range in testing heterogeneity of variance. Biometrika, 53(1/2):221–227, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334068.

Leadbetter:1963:SER

[Lea63] M. R. Leadbetter. On series expansions for the renewal moments. Biometrika, 50(1/2):75–80, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333748.

Leadbetter:1964:BEL

[Lea64] M. R. Leadbetter. Bounds on the error in the linear approximation to the renewal function. Biometrika, 51(3/4):355–364, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334141.

Lindley:1960:TMI

[LEH60] D. V. Lindley, D. A. East, and P. A. Hamilton. Tables for making inferences about the variance of a normal distribution. Biometrika, 47(3/4):433–437, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2333313. See corrigenda [LEH61].

Lindley:1961:CTM

[LEH61] D. V. Lindley, D. A. East, and P. A. Hamilton. Corrigenda: “Tables for Making Inferences about the Variance of a Normal Distribution”. Biometrika, 48(1/2):230, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333159. See [LEH60].

Leslie:1960:NSA

[Les60] P. H. Leslie. A note on some approximations to the variance in discrete-time stochastic models for biological systems. Biometrika, 47(1/2):196–197, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332975.

Leslie:1962:SMT

[Les62] P. H. Leslie. A stochastic model for two competing species of tribolium and its application to some experimental data. Biometrika, 49(1/2):1–25, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333464.

Lewis:1965:SRT

[Lew65] Peter A. W. Lewis. Some results on tests for Poisson processes. Biometrika, 52(1/2):67–77, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333812.

Leslie:1960:PSM

[LG60] P. H. Leslie and J. C. Gower. The properties of a stochastic model for the predator-prey type of interaction between two species. Biometrika, 47(3/4):219–234, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333294.

Linhart:1960:ATR

[Lin60] H. Linhart. Approximate tests for m rankings. Biometrika, 47 (3/4):476–480, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333323.

Lindley:1963:BRB

[Lin63] D. V. Lindley. Book review: The Theory of Probability by B. V. Gnedenko; B. D. Seckler. Biometrika, 50(1/2):223, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333767.

Lindley:1969:BRB

[Lin69] D. V. Lindley. Book review: The Structure of Inference, by D. A. S. Fraser. Biometrika, 56(2):453–456, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334442.

Little:1968:NEQ

[Lit68] R. E. Little. A note on estimation for quantal response data. Biometrika, 55(3):578–579, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334266.

Lemon:1969:EBS

[LK69] Glen H. Lemon and Richard G. Krutchkoff. An empirical Bayes smoothing technique. Biometrika, 56(2):361–365, August 1969. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334428.

Laha:1960:PCQ

[LL60a] R. G. Laha and E. Lukacs. On a problem connected with quadratic regression. Biometrika, 47(3/4):335–343, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333304.

Laha:1960:CFN

[LL60b] R. G. Laha and E. Lukacs. On certain functions of normal variates which are uncorrelated of a higher order. Biometrika, 47(1/2):175– 176, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332969.

Lloyd:1960:BRB

[Llo60] E. H. Lloyd. Book review: Mathematical Statistics, by J. N. Kapur; H. C. Saxena. Biometrika, 47(3/4):486, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333329.

Lieberman:1963:STI

[LM63] Gerald J. Lieberman and Rupert G. Miller, Jr. Simultaneous tolerance intervals in regression. Biometrika, 50(1/2):155–168, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333756.

Lipton:1964:DMF

[LM64] S. Lipton and C. A. McGilchrist. The derivation of methods of fitting exponential regression curves. Biometrika, 51(3/4):504–508, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334162.

Lieberman:1967:USD

[LMH67] G. J. Lieberman, R. G. Miller, Jr., and M. A. Hamilton. Unlimited simultaneous discrimination intervals in regression. Biometrika, 54(1/2):133–145, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2333857. See correction [LMH71].

Lieberman:1971:CAU

[LMH71] G. J. Lieberman, R. G. Miller, Jr., and M. A. Hamilton. Corrections and amendments: “Unlimited Simultaneous Discrimination Intervals in Regression”. Biometrika, 58(3):687, December 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334413. See [LMH67].

Lomnicki:1966:NWR

[Lom66] Z. A. Lomnicki. A note on the Weibull renewal process. Biometrika, 53(3/4):375–381, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333644.

Low:1964:SVE

[Low64] Leone Y. Low. Sampling variances of estimates of components of variance from a non-orthogonal two-way classification. Biometrika, 51(3/4):491–494, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334157.

Loynes:1966:BRB

[Loy66] R. M. Loynes. Book review: Queues and Inventories: A Study of Their Basic Stochastic Processes, by N. U. Prabhu. Biometrika, 53(1/2):299, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334095.

Loynes:1967:BRBa

[Loy67a] R. M. Loynes. Book review: An Introduction to Probability Theory and its Applications, Vol. II, by William Feller. Biometrika, 54(1/2):336–337, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333889.

Loynes:1967:BRBb

[Loy67b] R. M. Loynes. Book review: Introduction to Measure and Probability, by J. F. C. Kingman; S. J. Taylor. Biometrika, 54(1/2):338, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333890.

Laycock:1968:ODR

[LS68] P. J. Laycock and S. D. Silvey. Optimal designs in regression problems with a general convex loss function. Biometrika, 55(1):53–66, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334452.

Mote:1965:IEM

[MA65] V. L. Mote and R. L. Anderson. An investigation of the effect of misclassification on the properties of χ^{2}-tests in the analysis of categorical data. Biometrika, 52(1/2):95–109, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333815.

Maag:1966:SAW

[Maa66] Urs R. Maag. A k-sample analogue of Watson’s U^{2} statistic. Biometrika, 53(3/4):579–583, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333662.

Mallows:1961:BRBa

[Mal61a] C. L. Mallows. Book review: Elementary Decision Theory, by Herman Chernoff; Lincoln E. Moses. Biometrika, 48(1/2):233, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333163.

Mallows:1961:BRBb

[Mal61b] C. L. Mallows. Book review: Individual Choice Behaviour, by R. Duncan Luce. Biometrika, 48(1/2):234, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333164.

Mallows:1961:LVR

[Mal61c] C. L. Mallows. Latent vectors of random symmetric matrices. Biometrika, 48(1/2):133–149, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333138.

Mallows:1968:IIM

[Mal68] C. L. Mallows. An inequality involving multinomial probabilities. Biometrika, 55(2):422–424, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334886.

Manly:1969:ACS

[Man69a] B. F. J. Manly. Approximations to the characteristics of some sequential tests. Biometrika, 56(1):203–206, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334705.

Manly:1969:SPM

[Man69b] B. F. J. Manly. Some properties of a method of estimating the size of mobile animal populations. Biometrika, 56(2):407–410, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334433.

Marsaglia:1963:END

[Mar63] George Marsaglia. Expressing the normal distribution with covariance matrix A + B in terms of one with covariance matrix A. Biometrika, 50(3/4):535–538, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333924.

Maritz:1966:SEB

[Mar66] J. S. Maritz. Smooth empirical Bayes estimation for one-parameter discrete distributions. Biometrika, 53(3/4):417–429, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333648.

Marcus:1967:MIM

[Mar67a] Allan H. Marcus. A multivariate immigration with multiple death process and applications to lunar craters. Biometrika, 54(1/2):251– 261, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333868.

Mardia:1967:CRC

[Mar67b] K. V. Mardia. Correlation of the ranges of correlated samples. Biometrika, 54(3/4):529–539, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335043.

Mardia:1967:SCC

[Mar67c] K. V. Mardia. Some contributions to contingency-type bivariate distributions. Biometrika, 54(1/2):235–249, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333867. See correction [Mar68a].

Maritz:1967:SEB

[Mar67d] J. S. Maritz. Smooth empirical Bayes estimation for continuous distributions. Biometrika, 54(3/4):435–450, December 1967. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335035. See corrections

[Mar68b].

Mardia:1968:CSC

[Mar68a] K. V. Mardia. Corrections: “Some Contributions to Contingency-Type Bivariate Distributions”. Biometrika, 55(3):597, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334276. See

[Mar67c].

Maritz:1968:CSE

[Mar68b] J. S. Maritz. Corrections: “Smooth Empirical Bayes Estimation for Continuous Distribution”. Biometrika, 55(3):597, November 1968.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334275. See [Mar67d].

Maritz:1968:SEB

[Mar68c] J. S. Maritz. On the smooth empirical Bayes approach to testing of hypotheses and the compound decision problem. Biometrika, 55(1):83–100, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334454.

Mardia:1969:PST

[Mar69a] K. V. Mardia. The performance of some tests of independnece for contingency-type bivariate distributions. Biometrika, 56(2):449– 451, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334440.

Maritz:1969:EBE

[Mar69b] J. S. Maritz. Empirical Bayes estimation for the Poisson distribution. Biometrika, 56(2):349–359, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334427.

McGregor:1965:ADC

[MB65] J. R. McGregor and U. M. Bielenstein. The approximate distribution of the correlation between two stationary linear Markov series. II. Biometrika, 52(1/2):301–302, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333841.

McFadden:1960:TEQ

[McF60] J. A. McFadden. Two expansions for the quadrivariate normal integral. Biometrika, 47(3/4):325–333, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333303.

McGregor:1960:ATS

[McG60] J. R. McGregor. An approximate test for serial correlation in polynomial regression. Biometrika, 47(1/2):111–119, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332964.

McGregor:1962:ADC

[McG62] J. R. McGregor. The approximate distribution of the correlation between two stationary linear Markov series. Biometrika, 49 (3/4):379–388, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333972.

McGilchrist:1967:APC

[McG67] C. McGilchrist. Analysis of plant competition experiments for different ratios of species. Biometrika, 54(3/4):471–477, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335038.

McNeil:1966:CSE

[McN66] D. R. McNeil. Consistent statistics for estimating and testing hypotheses from grouped samples. Biometrika, 53(3/4):545–557, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333658.

McNeil:1968:QPT

[McN68] Donald R. McNeil. Quick powerful tests with grouped data. Biometrika, 55(1):264–268, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334476.

Moran:1962:RCS

[MD62] P. A. P. Moran and S. Fazekas De St Groth. Random circles on a sphere. Biometrika, 49(3/4):389–396, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333973.

Mead:1966:QME

[Mea66] R. Mead. A quick method of estimating the standard deviation. Biometrika, 53(3/4):559–564, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333659.

Mendenhall:1965:BRB

[Men65] W. Mendenhall. Book review: Statistical Assessment of the Life Characteristic, by W. R. Buckland; M. G. Kendall. Biometrika, 52(3/4):672, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333732.

Mehta:1969:UIS

[MG69a] J. S. Mehta and John Gurland. On utilizing information from a second sample in estimating variance. Biometrika, 56(3):527–532, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334659.

Mehta:1969:TEM

[MG69b] J. S. Mehta and John Gurland. Testing equality of means in the presence of correlation. Biometrika, 56(1):119–126, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334696.

McCullough:1960:SSB

[MGR60] Roger S. McCullough, John Gurland, and Lloyd Rosenberg. Small sample behaviour of certain tests of the hypothesis of equal means under variance heterogeneity. Biometrika, 47(3/4):345–353, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333305. See

corrigenda [MGR61].

McCullough:1961:CSS

[MGR61] R. S. McCullough, J. Gurland, and L. Rosenberg. Corrigenda: “Small Sample Behaviour of Certain Tests of the Hypothesis of Equal Means under Variance Heterogeneity”. Biometrika, 48(1/2): 230, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333157.

See [MGR60].

Mikhail:1965:CTW

[Mik65] N. N. Mikhail. A comparison of tests of the Wilks–Lawley hypothesis in multivariate analysis. Biometrika, 52(1/2):149–156, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333820.

Miller:1961:BRB

[Mil61] Alan J. Miller. Book review: Numerical Methods of Curve Fitting, by P. G. Guest. Biometrika, 48(3/4):482–483, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332795.

Miller:1963:BRB

[Mil63] A. J. Miller. Book review: Les Phenomenes d’Attente, Theorie et Applications, by A. Kaufmann; R. Cruon; Elements of Queueing Theory with Applications. by T. L. Saaty; Introduction to the Theory of Queues. by L. Takacs. Biometrika, 50(3/4):551–552, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333937.

Miles:1965:RR

[Mil65a] R. E. Miles. On random rotations in R^{3}. Biometrika, 52(3/4): 636–639, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333716.

Miller:1965:SMD

[Mil65b] K. S. Miller. Some multivariate density functions of products of Gaussian variates. Biometrika, 52(3/4):645–646, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333719.

Miles:1969:AVC

[Mil69] R. E. Miles. The asymptotic values of certain coverage probabilities. Biometrika, 56(3):661–680, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334674.

Mitchell:1968:ERC

[Mit68] Ann F. S. Mitchell. Exponential regression with correlated observations. Biometrika, 55(1):149–162, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334459.

Martz:1969:EBE

[MK69] Harry F. Martz, Jr. and Richard G. Krutchkoff. Empirical Bayes estimators in a multiple linear regression model. Biometrika, 56(2):367–374, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334429.

Mostafa:1964:PEB

[MM64] M. D. Mostafa and M. W. Mahmoud. On the problem of estimation for the bivariate lognormal distribution. Biometrika, 51 (3/4):522–527, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334167.

Mohn:1968:JDT

[Moh68] Erik Mohn. The joint distribution of two Studentized regression coefficients. Biometrika, 55(2):424–425, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334887.

Moore:1961:BRB

[Moo61] P. G. Moore. Book review: The Statistical Basis of Quality Control Charts, by S. K. Ekambaram. Biometrika, 48(1/2):234–235, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333166.

Moore:1963:BRB

[Moo63] P. G. Moore. Book review: Measurement of Mortality by Harry Gershenson. Biometrika, 50(1/2):229–230, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333775.

Moore:1966:BRB

[Moo66] P. G. Moore. Book review: Introductory Engineering Statistics, by I. Guttman; S. S. Wilks. Biometrika, 53(1/2):302, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334100.

Morgenthaler:1961:SCC

[Mor61] George W. Morgenthaler. Some circular coverage problems. Biometrika, 48(3/4):313–324, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332753.

Morton:1962:BRB

[Mor62] G. Morton. Book review: Operations Research/Management Science International Literature Digest Service, by R. S. Titchen; A. J. Rosenthal. Biometrika, 49(1/2):294, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333516.

Morris:1963:NDI

[Mor63a] K. W. Morris. A note on direct and inverse binomial sampling. Biometrika, 50(3/4):544–545, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333928.

Morrison:1963:ECS

[Mor63b] Donald F. Morrison. Expectations and covariances of serial and cross-correlation coefficients in a complex stationary time series. Biometrika, 50(1/2):213–216, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333765.

Morgan:1965:EPS

[Mor65] R. W. Morgan. The estimation of parameters from the spread of a disease by considering households of two. Biometrika, 52(1/2):271– 274, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333831.

Moran:1966:MLC

[Mor66] P. A. P. Moran. Measuring the length of a curve. Biometrika, 53 (3/4):359–364, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333642.

Moran:1967:TCB

[Mor67a] P. A. P. Moran. Testing for correlation between non-negative variates. Biometrika, 54(3/4):385–394, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335030.

Moran:1967:TSC

[Mor67b] P. A. P. Moran. Testing for serial correlation with exponentially distributed variates. Biometrika, 54(3/4):395–401, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335031.

Moran:1968:STH

[Mor68] P. A. P. Moran. Statistical theory of a high-speed photoelectric planimeter. Biometrika, 55(2):419–422, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334885.

Moran:1969:SIB

[Mor69] P. A. P. Moran. Statistical inference with bivariate gamma distributions. Biometrika, 56(3):627–634, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334670.

Mosimann:1962:CMD

[Mos62] James E. Mosimann. On the compound multinomial distribution, the multivariate β-distribution, and correlations among proportions. Biometrika, 49(1/2):65–82, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333468.

Mosimann:1963:CNM

[Mos63] James E. Mosimann. On the compound negative multinomial distribution and correlations among inversely sampled pollen counts. Biometrika, 50(1/2):47–54, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333745.

Mostafa:1967:DSE

[Mos67a] M. G. Mostafa. Designs for the simultaneous estimation of functions of variance components from two-way crossed classifications. Biometrika, 54(1/2):127–131, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333856.

Mostafa:1967:NTH

[Mos67b] M. G. Mostafa. Note on testing hypotheses in an unbalanced random effects model. Biometrika, 54(3/4):659–662, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335058.

Mott:1962:LRC

[Mot62] J. L. Mott. The latent roots of certain stochastic matrices. Biometrika, 49(1/2):264–265, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333493.

Mitchell:1968:FRE

[MR68] Ann F. S. Mitchell and Margaret A. Robb. Further remarks on exponential regression with correlated observations. Biometrika, 55(3):575–578, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334265.

Mann:1969:EWA

[MS69] N. R. Mann and S. C. Saunders. On evaluation of warranty assurance when life has a Weibull distribution. Biometrika, 56 (3):615–625, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334669.

Myers:1966:BCW

[MSA66] Max H. Myers, Marvin A. Schneiderman, and Peter Armitage. Boundaries for closed (wedge) sequential t test plans. Biometrika, 53(3/4):431–437, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333649.

McDonald:1967:RSM

[MT67] B. J. McDonald and W. A. Thompson, Jr. Rank sum multiple comparisons in one- and two-way classifications. Biometrika, 54 (3/4):487–497, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2335040. See correction [MT72].

McDonald:1972:CAR

[MT72] B. J. McDonald and W. A. Thompson, Jr. Corrections and amendments: “Rank sum multiple comparisons in one- and two-way classifications” (Biometrika 54 (1967), 487–497). Biometrika, 59(3):699, December 1972. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334827.

See [MT67].

Muller:1965:MCB

[Mul65] E.-R. Muller. A method of constructing balanced incomplete block designs. Biometrika, 52(1/2):285–288, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333836.

Muller:1966:BCF

[Mul66] E.-R. Muller. Balanced confounding of factorial experiments. Biometrika, 53(3/4):507–524, December 1966. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333655.

Murphy:1967:STS

[Mur67] Brian P. Murphy. Some two-sample tests when the variances are unequal: A simulation study. Biometrika, 54(3/4):679–683, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335066.

Mosteller:1961:TFT

[MY61] Frederick Mosteller and Cleo Youtz. Tables of the Freeman– Tukey transformations for the binomial and Poisson distributions. Biometrika, 48(3/4):433–440, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332765.

Nabeya:1961:AIM

[Nab61] S. Nabeya. Absolute and incomplete moments of the multivariate normal distribution. Biometrika, 48(1/2):77–84, June 1961. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333132.

Naus:1965:CRP

[Nau65] J. I. Naus. Clustering of random points in two dimensions. Biometrika, 52(1/2):263–267, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333829.

Nicholson:1961:OPD

[Nic61] W. L. Nicholson. Occupancy probability distribution critical points. Biometrika, 48(1/2):175–180, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333141.

Noether:1966:BRB

[Noe66] G. E. Noether. Book review: Handbook of Nonparametric Statistics, II, by John E. Walsh. Biometrika, 53(1/2):303, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334101.

Ohlsen:1964:EEP

[Ohl64] Sally Ohlsen. On estimating epidemic parameters from household data. Biometrika, 51(3/4):511–512, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334164.

Okamoto:1961:CTI

[OI61a] M. Okamoto and G. Ishii. Corrigenda: “Test of Independence in Intraclass 2 × 2 tables”. Biometrika, 48(3/4):476, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332783. See [OI61b].

Okamoto:1961:TII

[OI61b] Masashi Okamoto and Goro Ishii. Test of independence in intraclass 2 × 2 tables. Biometrika, 48(1/2):181–190, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333142. See corrigenda [OI61a].

Okamoto:1963:CSS

[Oka63] Masashi Okamoto. Chi-square statistic based on the pooled frequencies of several observations. Biometrika, 50(3/4):524–528, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333922.

Ord:1967:SDD

[Ord67] J. K. Ord. On a system of discrete distributions. Biometrika, 54 (3/4):649–656, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335056.

Ord:1968:ADF

[Ord68] J. K. Ord. Approximations to distribution functions which are hypergeometric series. Biometrika, 55(1):243–248, March 1968. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334468.

Olkin:1965:IET

[OS65] Ingram Olkin and Milton Sobel. Integral expressions for tail probabilities of the multinomial and negative multinomial distributions. Biometrika, 52(1/2):167–179, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333822.

Owen:1965:SCB

[Owe65] D. B. Owen. A special case of a bivariate non-central t-distribution. Biometrika, 52(3/4):437–446, December 1965. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333696.

Page:1962:MCC

[Pag62] E. S. Page. A modified control chart with warning lines. Biometrika, 49(1/2):171–176, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333478.

Parry:1962:EPS

[Par62] William Parry. Ergodic properties of some permutation processes. Biometrika, 49(1/2):151–154, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333475.

Patterson:1960:FNS

[Pat60] H. D. Patterson. A further note on a simple method for fitting an exponential curve. Biometrika, 47(1/2):177–180, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332970.

Patil:1962:MLE

[Pat62] G. P. Patil. Maximum likelihood estimation for generalized power series distributions and its application to a truncated binomial distribution. Biometrika, 49(1/2):227–237, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333484.

Patil:1963:CET

[Pat63] G. P. Patil. A characterization of the exponential-type distribution. Biometrika, 50(1/2):205–207, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333763.

Pathak:1964:ISU

[Pat64] P. K. Pathak. On inverse sampling with unequal probabilities. Biometrika, 51(1/2):185–193, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334206.

Patil:1965:ABF

[Pat65] V. H. Patil. Approximation to the Behrens–Fisher distributions. Biometrika, 52(1/2):267–271, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333830.

Patterson:1968:SFD

[Pat68] H. D. Patterson. Serial factorial design. Biometrika, 55(1):67–81, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334453.

Patil:1969:AGB

[Pat69] Venkutai Patil. Approximation to the generalized Behrens–Fisher distribution involving three variates. Biometrika, 56(3):687–689, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334678.

Pillai:1961:UPP

[PB61] K. C. Sreedharan Pillai and Angeles R. Buenaventura. Upper percentage points of a substitute F -ratio using ranges. Biometrika, 48(1/2):195–196, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333144.

Posten:1964:PLR

[PB64] H. O. Posten and R. E. Bargmann. Power of the likelihood-ratio test of the general linear hypothesis in multivariate analysis. Biometrika, 51(3/4):467–480, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334153.

Pearce:1960:SB

[Pea60] S. C. Pearce. Supplemented balance. Biometrika, 47(3/4):263–271, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333298.

See corrigenda [Pea61].

Pearce:1961:CSB

[Pea61] S. C. Pearce. Corrigenda: “Supplemented Balance”. Biometrika, 48(3/4):475, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332778. See [Pea60].

Pearson:1962:WPE

[Pea62] E. S. Pearson. William Palin Elderton, 1877–1962. Biometrika, 49(3/4):297–303, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333964.

Pearson:1963:CTR

[Pea63a] E. S. Pearson. Comparison of tests for randomness of points on a line. Biometrika, 50(3/4):315–325, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333902.

Pearson:1963:SPA

[Pea63b] E. S. Pearson. Some problems arising in approximating to probability distributions, using moments. Biometrika, 50(1/2):95–112, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333751.

Pearson:1964:BRB

[Pea64] E. S. Pearson. Book review: New Tables of the Incomplete Gamma-Function Ratio and of Percentage Points of the Chi-Square and Beta Distributions, by H. Leon Harter; Percentage Points of the Beta Distribution. by L. E. Vogler. Biometrika, 51(3/4):529–531, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334170.

Pearson:1965:SHP

[Pea65a] E. S. Pearson. Studies in the history of probability and statistics. XIV. Some incidents in the early history of biometry and statistics, 1890–94. Biometrika, 52(1/2):3–18, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333808.

Pearson:1965:TPP

[Pea65b] E. S. Pearson. Tables of percentage points of \sqrt{b_{1}} and b_{2} in normal samples; a rounding off. Biometrika, 52(1/2):282–285, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333835.

Pearson:1966:ATH

[Pea66a] E. S. Pearson. Alternative tests for heterogeneity of variance; some Monte Carlo results. Biometrika, 53(1/2):229–234, June 1966. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334069.

Pearson:1966:BRBa

[Pea66b] E. S. Pearson. Book review: Bibliography of Statistical Literature, by Maurice G. Kendall; Alison G. Doig. Biometrika, 53(1/2):298, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334094.

Pearson:1966:BRBb

[Pea66c] E. S. Pearson. Book review: Tracts for Computers, No. XXVII: Random Negative Exponential Deviates, by V. D. Barnett; Tracts for Computers, No.XXVIII: The Fitting of Growth and Allied Curves of the Asymptotic Regression Type by Stevens’s Method. by R. W. Hiorns. Biometrika, 53(1/2):304–305, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334103.

Pearson:1967:BRB

[Pea67a] E. S. Pearson. Book review: Tables of the Incomplete Gamma Function Ratio: Chi-Square Integral, Poisson Distribution, by Salem H. Khamis; W. Rudert. Biometrika, 54(3/4):690–691, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335073.

Pearson:1967:SHP

[Pea67b] E. S. Pearson. Studies in the history of probability and statistics. XVII. Some reflexions on continuity in the development of mathematical statistics, 1885–1920. Biometrika, 54(3/4):341–355, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335027.

Pearce:1968:MEE

[Pea68a] S. C. Pearce. The mean efficiency of equi-replicate designs. Biometrika, 55(1):251–253, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334470.

Pearson:1968:LCI

[Pea68b] E. S. Pearson. Lagrangian coefficients for interpolation between tabled percentage points. Biometrika, 55(1):19–28, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334449.

Pearson:1968:SHP

[Pea68c] E. S. Pearson. Studies in the history of probability and statistics. XX. Some early correspondence between W. S. Gosset, R. A. Fisher and Karl Pearson, with notes and comments. Biometrika, 55(3):445–457, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334250.

Pearson:1969:SCA

[Pea69] E. S. Pearson. Some comments on the accuracy of Box’s approximations to the distribution of M. Biometrika, 56(1):219–220, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334709.

Pillai:1969:EDW

[PG69] K. C. Sreedharan Pillai and Arjun K. Gupta. On the exact distribution of Wilks’s criterion. Biometrika, 56(1):109–118, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334695. See cor-

rection [PG70].

Pillai:1970:CED

[PG70] K. C. Sreedharan Pillai and Arjun K. Gupta. Corrections: “On the Exact Distribution of Wilks’s Criterion”. Biometrika, 57(1):225, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334963. See

[PG69].

Han:1968:THS

[pH68] Chien pai Han. Testing the homogeneity of a set of correlated variances. Biometrika, 55(2):317–326, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334872.

Phatarfod:1965:SAD

[Pha65] R. M. Phatarfod. Sequential analysis of dependent observations. I. Biometrika, 52(1/2):157–165, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333821.

Philip:1966:SIE

[Phi66] J. R. Philip. Some integral equations in geometrical probability. Biometrika, 53(3/4):365–374, December 1966. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333643.

Pillai:1964:DLS

[Pil64] K. C. Sreedharan Pillai. On the distribution of the largest of seven roots of a matrix in multivariate analysis. Biometrika, 51(1/2):270– 275, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334222.

Pillai:1965:ESF

[Pil65a] K. C. Sreedharan Pillai. On elementary symmetric functions of the roots of two matrices in multivariate analysis. Biometrika, 52 (3/4):499–506, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333701.

Pillai:1965:DLC

[Pil65b] K. C. Sreedharan Pillai. On the distribution of the largest characteristic root of a matrix in multivariate analysis. Biometrika, 52 (3/4):405–414, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333693.

Pillai:1967:UPP

[Pil67] K. C. Sreedharan Pillai. Upper percentage points of the largest root of a matrix in multivariate analysis. Biometrika, 54(1/2):189–194, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333862.

Pinkham:1962:API

[Pin62] R. S. Pinkham. An approximation to the probability integral of the gamma distribution for small values of the shape parameter. Biometrika, 49(1/2):276–278, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333497.

Pillai:1967:PCT

[PJ67] K. C. Sreedharan Pillai and Kanta Jayachandran. Power comparisons of tests of two multivariate hypotheses based on four criteria. Biometrika, 54(1/2):195–210, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333863.

Pillai:1968:PCT

[PJ68] K. C. Sreedharan Pillai and Kanta Jayachandran. Power comparisons of tests of equality of two covariance matrices based on four criteria. Biometrika, 55(2):335–342, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334874.

Plackett:1964:CCT

[Pla64] R. L. Plackett. The continuity correction in 2×2 tables. Biometrika, 51(3/4):327–337, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334138.

Pollak:1966:SEF

[Pol66a] Edward Pollak. Some effects of fluctuating offspring distributions on the survival of a gene. Biometrika, 53(3/4):391–396, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333646.

Pollard:1966:UDM

[Pol66b] J. H. Pollard. On the use of the direct matrix product in analysing certain stochastic population models. Biometrika, 53(3/4):397–415, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333647.

Pollard:1968:NMT

[Pol68] J. H. Pollard. A note on multi-type Galton–Watson processes with random branching probabilities. Biometrika, 55(3):589–590, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334271.

Proschan:1964:ANC

[PP64] Frank Proschan and Ronald Pyke. Asymptotic normality of certain test statistics of exponentiality. Biometrika, 51(1/2):253–256, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334213.

Potthoff:1964:GMA

[PR64] Richard F. Potthoff and S. N. Roy. A generalized multivariate analysis of variance model useful especially for growth curve problems. Biometrika, 51(3/4):313–326, December 1964. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334137.

Pratt:1966:ONS

[Pra66] John W. Pratt. The outer needle of some Bayes sequential continuation regions. Biometrika, 53(3/4):455–467, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333651.

Preece:1966:SBI

[Pre66] D. A. Preece. Some balanced incomplete block designs for two sets of treatments. Biometrika, 53(3/4):497–506, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333654. See correction [Pre69].

Preece:1967:NBI

[Pre67] D. A. Preece. Nested balanced incomplete block designs. Biometrika, 54(3/4):479–486, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335039.

Preece:1969:CSB

[Pre69] D. A. Preece. Correction: “Some Balanced Incomplete Block Designs for Two Sets of Treatments”. Biometrika, 56(3):691, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334681. See

[Pre66].

Priestley:1963:SCP

[Pri63] M. B. Priestley. The spectrum of a continuous process derived from a discrete process. Biometrika, 50(3/4):517–520, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333919.

Price:1964:SNC

[Pri64a] Robert Price. Some non-central F -distributions expressed in closed form. Biometrika, 51(1/2):107–122, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334200.

Priestley:1964:ATD

[Pri64b] M. B. Priestley. The analysis of two-dimensional stationary processes with discontinuous spectra. Biometrika, 51(1/2):195–217, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334207.

Pearson:1962:TRC

[PS62a] E. S. Pearson and Barbara A. S. Snow. Tests for rank correlation coefficients. III. Distribution of the transformed Kendall coefficient. Biometrika, 49(1/2):185–191, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333480.

Pearson:1962:GFT

[PS62b] E. S. Pearson and M. A. Stephens. The goodness-of-fit tests based on W_{N}^{2} and U_{N}^{2}. Biometrika, 49(3/4):397–402, December 1962. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333974.

Pearson:1964:RRS

[PS64] E. S. Pearson and M. A. Stephens. The ratio of range to standard deviation in the same normal sample. Biometrika, 51(3/4):484–487, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334155.

Pearson:1965:AMS

[PT65] E. S. Pearson and J. W. Tukey. Approximate means and standard deviations based on distances between percentage points of frequency curves. Biometrika, 52(3/4):533–546, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333703.

Puri:1966:HBD

[Pur66] Prem S. Puri. On the homogeneous birth-and-death process and its integral. Biometrika, 53(1/2):61–71, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334053.

Putter:1964:GFT

[Put64] Joseph Putter. The χ^{2} goodness-of-fit test for a class of cases of dependent observations. Biometrika, 51(1/2):250–252, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334212.

Potthoff:1966:THB

[PW66a] Richard F. Potthoff and Maurice Whittinghill. Testing for homogeneity: I. The binomial and multinomial distributions. Biometrika, 53(1/2):167–182, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334062.

Potthoff:1966:THI

[PW66b] Richard F. Potthoff and Maurice Whittinghill. Testing for homogeneity: II. The Poisson distribution. Biometrika, 53(1/2):183–190, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334063.

Quesenberry:1961:STO

[QD61] C. P. Quesenberry and H. A. David. Some tests for outliers. Biometrika, 48(3/4):379–390, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332759.

Read:1968:SML

[RA68] K. L. Q. Read and J. R. Ashford. A system of models for the life cycle of a biological organism. Biometrika, 55(1):211–221, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334465.

Rabinovitch:1969:SHP

[Rab69] Nachum L. Rabinovitch. Studies in the history of probability and statistics. XXII: Probability in the Talmud. Biometrika, 56(2):437– 441, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334437.

Radcliffe:1964:CMU

[Rad64] J. Radcliffe. The construction of a matrix used in deriving tests of significance in multivariate analysis. Biometrika, 51(3/4):503–504, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334161.

Radcliffe:1966:BRB

[Rad66] J. Radcliffe. Book review: Statistical Estimation in Factor Analysis. A New Technique and its Foundation, by K. G. Joreskog. Biometrika, 53(1/2):301, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334098.

Radcliffe:1967:NAF

[Rad67] J. Radcliffe. A note on an approximate factorization in discriminant analysis. Biometrika, 54(3/4):665–668, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335060.

Raghavarao:1962:BUB

[Rag62] Damaraju Raghavarao. On balanced unequal block designs. Biometrika, 49(3/4):561–562, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333990.

Ramasubban:1960:SDA

[Ram60] T. A. Ramasubban. Some distributions arising in the study of generalized mean differences. Biometrika, 47(3/4):469–473, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333321. See corri-

genda [Ram61].

Ramasubban:1961:CSD

[Ram61] T. A. Ramasubban. Corrigenda: “Some Distributions Arising in the Study of Generalized Mean Differences”. Biometrika, 48(1/2):230, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333158. See

[Ram61].

Rao:1960:FCL

[Rao60] B. Raja Rao. A formula for the curvature of the likelihood surface of a sample drawn from a distribution admitting sufficient statistics. Biometrika, 47(1/2):203–207, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332979.

Rao:1962:NGL

[Rao62] B. Raja Rao. A note on the geometry of linear estimation. Biometrika, 49(3/4):560–561, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333989.

Rao:1963:FNG

[Rao63] B. Raja Rao. A further note on the geometry of linear estimation. Biometrika, 50(3/4):540–541, December 1963. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333926.

Rao:1965:TLS

[Rao65a] C. Radhakrishna Rao. The theory of least squares when the parameters are stochastic and its application to the analysis of growth curves. Biometrika, 52(3/4):447–458, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333697.

Rao:1965:NER

[Rao65b] J. N. K. Rao. A note on estimation of ratios by Quenouille’s method. Biometrika, 52(3/4):647–649, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333720.

Rao:1966:DFB

[Rao66] C. Radhakrishna Rao. Discriminant function between composite hypotheses and related problems. Biometrika, 53(3/4):339–345, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333640.

Rao:1967:PMU

[Rao67] J. N. K. Rao. The precision of Mickey’s unbiased ratio estimator. Biometrika, 54(1/2):321–324, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333880.

Rao:1968:TPS

[Rao68a] Poduri S. R. S. Rao. On three procedures of sampling from finite populations. Biometrika, 55(2):438–441, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334893.

Rao:1968:NAR

[Rao68b] T. Subba Rao. A note on the asymptotic relative efficiencies of Cox and Stuart’s tests for testing trend in dispersion of a p-dependent time series. Biometrika, 55(2):381–385, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334880. See correction [Rao69].

Rao:1969:CNA

[Rao69] T. Subba Rao. Corrections: “A Note on the Asymptotic Relative Efficiencies of Cox and Stuart’s Tests for Testing Trend in Dispersion of a p-Dependent Time Series”. Biometrika, 56(2):457, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334444. See [Rao68b].

Rees:1969:AVS

[Ree69] D. H. Rees. The analysis of variance of some non-orthogonal designs with split plots. Biometrika, 56(1):43–54, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334689.

Reid:1960:BRBa

[Rei60a] D. D. Reid. Book review: Health and Medical Care in New York, by Committee for the Special Research Project in the Health Insurance Plan of Greater New York. Biometrika, 47(1/2):212–213, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332988.

Reid:1960:BRBb

[Rei60b] D. D. Reid. Book review: Tuberculosis in White and Negro Children: 1. The Roentgenologic Aspects of the Harriet Lane Study, by Janet B. Hardy; Tuberculosis in White and Negro Children: 2. The Epidemiological Aspects of the Harriet Lane Study. by Miriam E. Brailey. Biometrika, 47(1/2):213–214, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332989.

Reiersol:1961:LNL

[Rei61] Olav Reiersøl. Linear and non-linear multiple comparisons in logit analysis. Biometrika, 48(3/4):359–365, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332757. See corrigenda [Rei62].

Reiersol:1962:CLN

[Rei62] Olav Reiersøl. Corrigenda: “Linear and Non-Linear Multiple Comparisons in Logit Analysis”. Biometrika, 49(1/2):284, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333502. See [Rei61].

Roy:1957:FCM

[RG57] S. N. Roy and R. Gnanadesikan. Further contributions to multivariate confidence bounds. Biometrika, 44(3/4):399–410, December 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332872. See corri-

genda [RG61].

Roy:1961:CFC

[RG61] S. N. Roy and R. Gnanadesikan. Corrigenda: “Further Contributions to Multivariate Confidence Bounds”. Biometrika, 48(3/4):474, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332775.

See [RG57].

Roberts:1966:NSC

[RG66] Charles Roberts and Seymour Geisser. A necessary and sufficient condition for the square of a random variable to be gamma. Biometrika, 53(1/2):275–278, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334082.

Rao:1968:CBS

[RGL68] B. Raja Rao, Mohan L. Garg, and C. C. Li. Correlation between the sample variances in a singly truncated bivariate normal distribution. Biometrika, 55(2):433–436, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334891.

Richards:1967:FLM

[Ric67] F. S. G. Richards. On finding local maxima of functions of a real variable. Biometrika, 54(1/2):310–311, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333876.

Rutherford:1967:EBAa

[RK67] J. R. Rutherford and R. G. Krutchkoff. The empirical Bayes approach: Estimating the prior distribution. Biometrika, 54(1/2):326– 328, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333882.

Rutherford:1969:AOE

[RK69a] J. R. Rutherford and R. G. Krutchkoff. ϵ asymptotic optimality of empirical Bayes estimators. Biometrika, 56(1):220–223, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334710.

Rutherford:1969:SEB

[RK69b] J. R. Rutherford and R. G. Krutchkoff. Some empirical Bayes techniques in point estimation. Biometrika, 56(1):133–137, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334698.

Roach:1963:FDS

[Roa63] S. A. Roach. The frequency distribution of the sample mean where each member of the sample is drawn from a different rectangular distribution. Biometrika, 50(3/4):508–513, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333917.

Roberts:1967:MCS

[Rob67] F. D. K. Roberts. A Monte Carlo solution of a two-dimensional unstructured cluster problem. Biometrika, 54(3/4):625–628, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335053.

Roberts:1968:RMT

[Rob68] F. D. K. Roberts. Random minimal trees. Biometrika, 55(1):255– 258, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334472.

Roberts:1969:NNP

[Rob69] F. D. K. Roberts. Nearest neighbours in a Poisson ensemble. Biometrika, 56(2):401–406, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334432.

Roberts:1968:TDC

[RS68] F. D. K. Roberts and S. H. Storey. A three-dimensional cluster problem. Biometrika, 55(1):258–260, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334473.

Remage:1966:MLP

[RT66] Russell Remage, Jr. and W. A. Thompson, Jr. Maximum-likelihood paired comparison rankings. Biometrika, 53(1/2):143–149, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334060.

Rohde:1969:EFS

[RT69] C. A. Rohde and G. M. Tallis. Exact first- and second-order moments of estimates of components of covariance. Biometrika, 56(3):517–525, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334658.

Ruben:1956:SSN

[Rub56] H. Ruben. On the sum of squares of normal scores. Biometrika, 43(3/4):456–458, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332924. See corrigenda [Rub65].

Ruben:1964:IFA

[Rub64] Harold Ruben. Irrational fraction approximations to Mills’ ratio. Biometrika, 51(3/4):339–345, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334139.

Ruben:1965:CSS

[Rub65] H. Ruben. Corrigenda: “On the sum of squares of normal scores”. Biometrika, 52(3/4):669, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333726. See [Rub56].

Rutherford:1967:EBAb

[Rut67] J. R. Rutherford. The empirical Bayes approach: Estimating posterior quantiles. Biometrika, 54(3/4):672–675, December 1967. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335063.

Robson:1964:ETP

[RW64] D. S. Robson and J. H. Whitlock. Estimation of a truncation point. Biometrika, 51(1/2):33–39, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334193.

Rao:1966:TMB

[RW66] J. N. K. Rao and J. T. Webster. On two methods of bias reduction in the estimation of ratios. Biometrika, 53(3/4):571–577, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333661.

Rybnikov:1966:BRB

[Ryb66] K. A. Rybnikov. Book review: Applied Combinatorial Mathematics, by E. F. Beckenbach. Biometrika, 53(1/2):294–295, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334090.

Schneiderman:1962:CST

[SA62a] M. A. Schneiderman and P. Armitage. Closed sequential t-tests. Biometrika, 49(3/4):359–366, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333970.

Schneiderman:1962:FCS

[SA62b] M. A. Schneiderman and P. Armitage. A family of closed sequential procedures. Biometrika, 49(1/2):41–56, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333466. See corrections [SA69].

Schneiderman:1969:CFC

[SA69] M. A. Schneiderman and P. Armitage. Corrections: “A Family of Closed Sequential Procedures”. Biometrika, 56(2):457, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334443. See [SA62b].

Sampford:1962:MCS

[Sam62] M. R. Sampford. Methods of cluster sampling with and without replacement for clusters of unequal sizes. Biometrika, 49(1/2):27–40, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333465.

Sampford:1967:SRU

[Sam67] M. R. Sampford. On sampling without replacement with unequal probabilities of selection. Biometrika, 54(3/4):499–513, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335041.

Sankaran:1963:ANC

[San63] Munuswamy Sankaran. Approximations to the non-central chisquare distribution. Biometrika, 50(1/2):199–204, June 1963. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333761.

Sankaran:1964:ABB

[San64] Munuswamy Sankaran. On an analogue of Bhattacharya bound. Biometrika, 51(1/2):268–270, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334221.

Saw:1960:NEA

[Saw60] J. G. Saw. A note on the error after a number of terms of the David– Johnson series for the expected values of normal order statistics. Biometrika, 47(1/2):79–86, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332960.

Saw:1961:BML

[Saw61a] J. G. Saw. The bias of the maximum likelihood estimates of the location and scale parameters given a Type II censored normal sample. Biometrika, 48(3/4):448–451, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332768.

Saw:1961:BRBa

[Saw61b] J. G. Saw. Book review: Applied Statistical Decision Theory, by Howard Raiffa; Robert Schlaifer. Biometrika, 48(3/4):478–479, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332787.

Saw:1961:BRBd

[Saw61c] J. G. Saw. Book review: Elementary Statistics, by Paul G. Hoel. Biometrika, 48(3/4):481, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332791.

Saw:1961:BRBc

[Saw61d] J. G. Saw. Book review: Probability. An Introduction, by Samuel Goldberg. Biometrika, 48(3/4):481, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332790.

Saw:1961:BRBb

[Saw61e] J. G. Saw. Book review: Regression Analysis, by R. L. Plackett. Biometrika, 48(3/4):480, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332789.

Saw:1961:ENP

[Saw61f] J. G. Saw. Estimation of the normal population parameters given a Type I censored sample. Biometrika, 48(3/4):367–377, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332758.

Saw:1962:BRB

[Saw62a] J. G. Saw. Book review: An Introduction to Experimental Design by William S. Ray. Biometrika, 49(3/4):578–579, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333999.

Saw:1962:EME

[Saw62b] J. G. Saw. Efficient moment estimators when the variables are dependent with special reference to Type II censored normal data. Biometrika, 49(1/2):155–161, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333476.

Saw:1964:SNV

[Saw64] John G. Saw. Some notes on variance-ratio tests of the general linear hypothesis. Biometrika, 51(3/4):508–511, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334163.

Saw:1966:CTC

[Saw66a] J. G. Saw. A conservative test for the concurrence of several regression lines and related problems. Biometrika, 53(1/2):272–275, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334081.

Saw:1966:NPC

[Saw66b] J. G. Saw. A non-parametric comparison of two samples one of which is censored. Biometrika, 53(3/4):599–602, December 1966.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333668.

Snell:1968:DBT

[SBJ68] E. J. Snell and J. Bryan-Jones. A design balanced for trend. Biometrika, 55(3):535–539, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334258.

Shenton:1965:MRP

[SC65] L. R. Shenton and J. A. Carpenter. The Mills ratio and the probability integral for a Pearson Type IV distribution. Biometrika, 52(1/2):119–126, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333817.

Saw:1966:CTE

[SC66] J. G. Saw and Bryant Chow. The curve through the expected values of ordered variates and the sum of squares of normal scores. Biometrika, 53(1/2):252–255, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334074.

Scheffe:1953:MJA

[Sch53] Henry Scheffé. A method for judging all contrasts in the analysis of variance. Biometrika, 40(1/2):87–104, June 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333100. See correction [Sch69b].

Schatzoff:1966:EDW

[Sch66] Martin Schatzoff. Exact distributions of Wilks’s likelihood ratio criterion. Biometrika, 53(3/4):347–358, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333641. See correction [Sch67].

Schatzoff:1967:CED

[Sch67] Martin Schatzoff. Correction: “Exact Distributions of Wilks’s Likelihood-Ratio Criterion”. Biometrika, 54(3/4):688, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2335070. See

[Sch66].

Schach:1969:NST

[Sch69a] Siegfried Schach. Nonparametric symmetry tests for circular distributions. Biometrika, 56(3):571–577, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334665.

Scheffe:1969:CMJ

[Sch69b] Henry Scheffe. Corrections: “A Method for Judging All Contrasts in the Analysis of Variance”. Biometrika, 56(1):229, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334713. See [Sch53].

Starks:1961:CST

[SD61a] T. H. Starks and H. A. David. Corrigenda: “Significant Tests for Paired-Comparison Experiments”. Biometrika, 48(3/4):475, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332780. See

[SD61b].

Starks:1961:STP

[SD61b] T. H. Starks and H. A. David. Significance tests for pairedcomparison experiments. Biometrika, 48(1/2):95–108, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333134. See corrigenda

[SD61a].

Seal:1967:SHP

[Sea67] Hilary L. Seal. Studies in the history of probability and statistics. XV. The historical development of the Gauss linear model. Biometrika, 54(1/2):1–24, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333849.

Seber:1962:MSS

[Seb62] G. A. F. Seber. The multi-sample single recapture census. Biometrika, 49(3/4):339–350, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333968.

Seber:1963:NCC

[Seb63] G. A. F. Seber. The non-central chi-squared and beta distributions. Biometrika, 50(3/4):542–544, December 1963. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333927.

Seber:1964:LHI

[Seb64] G. A. F. Seber. Linear hypotheses and induced tests. Biometrika, 51(1/2):41–47, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334194.

Seber:1965:NMR

[Seb65] G. A. F. Seber. A note on the multiple-recapture census. Biometrika, 52(1/2):249–259, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333827.

Sedransk:1966:ASS

[Sed66] J. Sedransk. An application of sequential sampling to analytical surveys. Biometrika, 53(1/2):85–97, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334055.

Selby:1964:GDS

[Sel64] B. Selby. Girdle distributions on a sphere. Biometrika, 51 (3/4):381–392, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334144.

Selby:1965:IDT

[Sel65] B. Selby. The index of dispersion as a test statistic. Biometrika, 52(3/4):627–629, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333713.

Sen:1967:NAE

[Sen67] Pranab Kumar Sen. A note on the asymptotic efficiency of Friedman’s χ_{r}^{2}-test. Biometrika, 54(3/4):677–679, December 1967. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2335065.

Severo:1969:PSE

[Sev69] Norman C. Severo. The probabilities of some epidemic models. Biometrika, 56(1):197–201, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334704.

Smith:1968:NCR

[SH68] W. B. Smith and H. O. Hartley. A note on the correlation of ranges in correlated normal sample. Biometrika, 55(3):595–597, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334274.

Shah:1966:NCP

[Sha66] B. K. Shah. A note on Craig’s paper on the minimum of binomial variates. Biometrika, 53(3/4):614–615, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333672.

Shanbhag:1968:SRC

[Sha68] D. N. Shanbhag. Some remarks concerning Khatri’s result on quadratic forms. Biometrika, 55(3):593–595, November 1968. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334273.

Shaman:1969:ICM

[Sha69] Paul Shaman. On the inverse of the covariance matrix of a first order moving average. Biometrika, 56(3):595–600, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334667.

Shenton:1958:MEM

[She58] L. R. Shenton. Moment estimators and maximum likelihood. Biometrika, 45(3/4):411–420, December 1958. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www. jstor.org/stable/2333188. See corrigenda [She59, She61].

Shenton:1959:CME

[She59] L. R. Shenton. Corrigenda: “Moment Estimators and Maximum Likelihood”. Biometrika, 46(3/4):502, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333569. See [She58].

Shenton:1961:CME

[She61] L. R. Shenton. Corrigenda: “Moment Estimators and Maximum Likelihood”. Biometrika, 48(3/4):474, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332776. See [She58].

Sheynin:1968:SHP

[She68] O. B. Sheynin. Studies in the history of probability and statistics. XXI. On the early history of the law of large numbers. Biometrika, 55(3):459–467, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334251.

Sillitto:1964:SRB

[Sil64] G. P. Sillitto. Some relations between expectations of order statistics in samples of different sizes. Biometrika, 51(1/2):259–262, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334216.

Sillitto:1969:DAI

[Sil69] G. P. Sillitto. Derivation of approximants to the inverse distribution function of a continuous univariate population from the order statistics of a sample. Biometrika, 56(3):641–650, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334672.

Singh:1967:EVR

[Sin67] Chatter Singh. On the extreme values and range of samples from non-normal populations. Biometrika, 54(3/4):541–550, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335044.

Siskind:1964:CSF

[Sis64] V. Siskind. On certain suggested formulae applied to the sequential t-test. Biometrika, 51(1/2):97–106, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334199.

Siskind:1965:SGS

[Sis65] V. Siskind. A solution of the general stochastic epidemic. Biometrika, 52(3/4):613–616, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333710.

Siskind:1968:UIV

[Sis68] V. Siskind. On using an incorrect value of σ_{B}^{2}/σ^{2} in balanced incomplete block designs. Biometrika, 55(1):254–255, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334471.

Singh:1968:MSS

[SJG68] D. Singh, K. K. Jindal, and J. N. Garg. On modified systematic sampling. Biometrika, 55(3):541–546, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334259.

Slater:1961:ISP

[Sla61] Patrick Slater. Inconsistencies in a schedule of paired comparisons. Biometrika, 48(3/4):303–312, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332752.

Slakter:1966:CVC

[Sla66] Malcolm J. Slakter. Comparative validity of the chi-square and two modified chi-square goodness-of-fit tests for small but equal expected frequencies. Biometrika, 53(3/4):619–623, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333675.

Stephens:1968:FPP

[SM68] M. A. Stephens and Urs R. Maag. Further percentage points for W_{N}^{2}. Biometrika, 55(2):428–430, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334889.

Smith:1960:BRB

[Smi60] Cedric A. B. Smith. Book review: String Figures and Other Monographs, by W. W. R. Ball; J. Petersen; H. S. Carslaw; F. Cajori. Biometrika, 47(3/4):490, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333337.

Smith:1962:BRBb

[Smi62a] Cedric A. B. Smith. Book review: Biometrical Genetics by Oscar Kempthorne. Biometrika, 49(3/4):581, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334005.

Smith:1962:BRBa

[Smi62b] Cedric A. B. Smith. Book review: Introduction to Quantitative Genetics, by D. S. Falconer. Biometrika, 49(1/2):292–293, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333514.

Smith:1963:BRB

[Smi63] T. M. F. Smith. Book review: Elements of Statistical Inference, by R. M. Kozelka. Biometrika, 50(3/4):553, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333939.

Snow:1962:TMK

[Sno62] Barbara A. S. Snow. The third moment of Kendall’s Tau in normal samples. Biometrika, 49(1/2):177–183, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333479.

Snow:1963:DKT

[Sno63] Barbara A. S. Snow. The distribution of Kendall’s Tau for samples of four from a normal bivariate population with correlation ρ. Biometrika, 50(3/4):538–540, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333925.

Steck:1962:NEM

[SO62] G. P. Steck and D. B. Owen. A note on the equicorrelated multivariate normal distribution. Biometrika, 49(1/2):269–271, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333495.

Sokal:1960:BRB

[Sok60] Robert R. Sokal. Book review: Darwin’s Biological Work. Some Aspects Reconsidered, by P. R. Bell. Biometrika, 47(3/4):488–489, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333335.

Sondhi:1961:NQN

[Son61] Man Mohan Sondhi. A note on the quadrivariate normal integral. Biometrika, 48(1/2):201–203, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333147.

Sprent:1965:FPC

[Spr65] P. Sprent. Fitting a polynomial to correlated equally spaced observations. Biometrika, 52(1/2):275–276, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333832.

Springer:1969:NOP

[Spr69] Basil G. F. Springer. Numerical optimization in the presence of random variability. The single factor case. Biometrika, 56(1):65–74, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334691.

Srivastava:1960:DRC

[Sri60] A. B. L. Srivastava. The distribution of regression coefficients in samples from bivariate non-normal populations: I. Theoretical investigation. Biometrika, 47(1/2):61–68, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332958.

Singh:1965:SRD

[SS65a] B. D. Singh and D. Singh. Some remarks on double sampling for stratification. Biometrika, 52(3/4):587–590, December 1965. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333708.

Stone:1965:PIQ

[SS65b] M. Stone and B. G. F. Springer. A paradox involving quasi prior distributions. Biometrika, 52(3/4):623–627, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333712.

Schoeman:1969:CSS

[SS69] A. Schoeman and D. E. W. Schumann. Comparison of the sensitivities of similar independent and non-independent experiments. Biometrika, 56(1):17–31, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334687.

Springer:1966:BCL

[ST66] M. D. Springer and W. E. Thompson. Bayesian confidence limits for the product of N binomial parameters. Biometrika, 53 (3/4):611–613, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333671.

Steck:1962:OPE

[Ste62a] G. P. Steck. Orthant probabilities for the equicorrelated multivariate normal distribution. Biometrika, 49(3/4):433–445, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333977.

Stephens:1962:EATa

[Ste62b] M. A. Stephens. Exact and approximate tests for directions. I. Biometrika, 49(3/4):463–477, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333980.

Stephens:1962:EATb

[Ste62c] M. A. Stephens. Exact and approximate tests for directions. II. Biometrika, 49(3/4):547–552, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333986.

Stephens:1963:DGF

[Ste63a] M. A. Stephens. The distribution of the goodness-of-fit statistic U_{N}^{2}. I. Biometrika, 50(3/4):303–313, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333901.

Stephens:1963:RWC

[Ste63b] M. A. Stephens. Random walk on a circle. Biometrika, 50 (3/4):385–390, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333907.

Stephens:1964:DGF

[Ste64] M. A. Stephens. The distribution of the goodness-of-fit statistic, U_{N}^{2}. II. Biometrika, 51(3/4):393–397, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334145.

Stephens:1965:GFS

[Ste65a] M. A. Stephens. The goodness-of-fit statistic V_{N}: Distribution and significance points. Biometrika, 52(3/4):309–321, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333685.

Stephens:1965:SPT

[Ste65b] M. A. Stephens. Significance points for the two-sample statistic U_{M,N}^{2}. Biometrika, 52(3/4):661–663, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333724.

Stephens:1966:SCU

[Ste66] M. A. Stephens. Statistics connected with the uniform distribution: Percentage points and application to testing for randomness of directions. Biometrika, 53(1/2):235–240, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334070.

Steiger:1967:SKT

[Ste67a] W. L. Steiger. Some Kolmogoroff-type inequalities for bounded random variables. Biometrika, 54(3/4):641–647, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2335055. See cor-

rection [Ste68b].

Stephens:1967:TDM

[Ste67b] M. A. Stephens. Tests for the dispersion and for the modal vector of a distribution on a sphere. Biometrika, 54(1/2):211–223, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333864.

Steck:1968:NCT

[Ste68a] G. P. Steck. A note on contingency-type bivariate distributions. Biometrika, 55(1):262–264, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334475.

Steiger:1968:CSK

[Ste68b] W. L. Steiger. Correction: “Some Kolmogoroff-Type Inqualities for Bounded Random Variables”. Biometrika, 55(2):443, July 1968.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334896. See [Ste67a].

Stephens:1969:GFS

[Ste69a] M. A. Stephens. A goodness-of-fit statistic for the circle, with some comparisons. Biometrika, 56(1):161–168, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334701.

Stephens:1969:MST

[Ste69b] M. A. Stephens. Multi-sample tests for the Fisher distribution for directions. Biometrika, 56(1):169–181, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334702.

Stephens:1969:TMD

[Ste69c] M. A. Stephens. Tests for the von Mises distribution. Biometrika, 56(1):149–160, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334700.

Stigler:1969:URA

[Sti69] Stephen M. Stigler. The use of random allocation for the control of selection bias. Biometrika, 56(3):553–560, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334663.

Stone:1962:BRB

[Sto62] M. Stone. Book review: Statistical Decision Theory by Lionel Weiss. Biometrika, 49(3/4):577–578, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333997.

Stone:1967:ETP

[Sto67] M. Stone. Extreme tail probabilities for the null distribution of the two-sample Wilcoxon statistic. Biometrika, 54(3/4):629–640, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2335054.

Stone:1968:ETP

[Sto68] M. Stone. Extreme tail probabilities for sampling without replacement and exact Bahadur efficiency of the two-sample normal scores test. Biometrika, 55(2):371–375, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334878.

Stone:1969:RER

[Sto69a] M. Stone. The role of experimental randomization in Bayesian statistics: Finite sampling and two Bayesians. Biometrika, 56 (3):681–683, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334675.

Stone:1969:RST

[Sto69b] M. Stone. The role of significance testing: Some data with a message. Biometrika, 56(3):485–493, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334654.

Stuart:1962:GDP

[Stu62] Alan Stuart. Gamma-distributed products of independent random variables. Biometrika, 49(3/4):564–565, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333992.

Stuart:1964:BRB

[Stu64] A. Stuart. Book review: Factor Analysis as a Statistical Method, by D. N. Lawley; A. E. Maxwell. Biometrika, 51(3/4):533, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334175.

Stuart:1967:ACV

[Stu67] Alan Stuart. The average critical value method and the asymptotic relative efficiency of tests. Biometrika, 54(1/2):308–310, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333875.

Subrahmaniam:1965:NET

[Sub65] K. Subrahmaniam. A note on estimation in the truncated Poisson. Biometrika, 52(1/2):279–282, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333834.

Subrahmaniam:1969:OSC

[Sub69] K. Subrahmaniam. Order statistics from a class of non-normal distributions. Biometrika, 56(2):415–428, August 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334435. See correction [Sub70].

Subrahmaniam:1970:COS

[Sub70] K. Subrahmaniam. Corrections: “Order Statistics from a Class of Non-Normal Distributions”. Biometrika, 57(1):226, April 1970. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334965. See [Sub69].

Sukhatme:1960:PST

[Suk60] Balkrishna V. Sukhatme. Power of some two-sample non-parametric tests. Biometrika, 47(3/4):355–362, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333306.

Simpson:1960:TBP

[SW60] J. A. Simpson and B. L. Welch. Table of the bounds of the probability integral when the first four moments are given. Biometrika, 47(3/4):399–410, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333310.

Shenton:1962:BME

[SW62] L. R. Shenton and P. A. Wallington. The bias of moment estimators with an application to the negative binomial distribution. Biometrika, 49(1/2):193–204, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333481.

Shapiro:1965:AVT

[SW65] S. S. Shapiro and M. B. Wilk. An analysis of variance test for normality: Complete samples. Biometrika, 52(3/4):591–611, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333709.

Swamy:1962:AIS

[Swa62] P. S. Swamy. On the amount of information supplied by censored samples of grouped observations in the estimation of statistical parameters. Biometrika, 49(1/2):245–249, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333487.

Swamy:1963:AIS

[Swa63] P. S. Swamy. On the amount of information supplied by truncated samples of grouped observations in the estimation of the parameters of normal populations. Biometrika, 50(1/2):207–213, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333764.

Swindel:1968:BSL

[Swi68] Benee F. Swindel. On the bias of some least-squares estimators of variance in a general linear model. Biometrika, 55(2):313–316, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334871.

Severo:1960:NAC

[SZ60] Norman C. Severo and Marvin Zelen. Normal approximation to the chi-square and non-central F probability functions. Biometrika, 47(3/4):411–416, December 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333311.

Tallis:1969:NCP

[Tal69] G. M. Tallis. Note on a calibration problem. Biometrika, 56 (3):505–508, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334656.

Tanner:1961:DBD

[Tan61] J. C. Tanner. A derivation of the Borel distribution. Biometrika, 48(1/2):222–224, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333154.

Tanner:1962:TAD

[Tan62] J. C. Tanner. A theoretical analysis of delays at an uncontrolled intersection. Biometrika, 49(1/2):163–170, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333477.

Tanner:1967:CUI

[Tan67] J. C. Tanner. The capacity of an uncontrolled intersection. Biometrika, 54(3/4):657–658, December 1967. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2335057.

Tarter:1968:ICA

[Tar68] M. E. Tarter. Inverse cumulative approximation and applications. Biometrika, 55(1):29–41, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334450.

Tarone:1979:TGF

[Tar79] R. E. Tarone. Testing the goodness of fit of the binomial distribution. Biometrika, 66(3):585–590, December 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335179. See correction [Tar90].

Tarone:1990:ACT

[Tar90] R. E. Tarone. Amendments and corrections: “Testing the Goodness of fit of the Binomial Distribution”. Biometrika, 77(3):668, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337011. See

[Wis68, Tar79].

Taylor:1963:BRB

[Tay63] Wallis Taylor. Book review: British Towns. A Statistical Study of their Social and Economic Differences by C. A. Moser; Wolf Scott. Biometrika, 50(1/2):228–229, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333774.

Taylor:1964:BRB

[Tay64] S. J. Taylor. Book review: Limit Theorem for Convolutions, by Harold Bergstrom. Biometrika, 51(3/4):534–535, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334178.

Tiao:1967:BAT

[TB67] G. C. Tiao and G. E. P. Box. Bayesian analysis of a three-component hierarchical design model. Biometrika, 54(1/2):109–125, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333855.

Tiao:1968:BAL

[TD68] G. C. Tiao and N. R. Draper. Bayesian analysis of linear models with two random components with special reference to the balanced incomplete block design. Biometrika, 55(1):101–117, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334455.

Tiao:1969:BEL

[TF69] G. C. Tiao and S. Fienberg. Bayesian estimation of latent roots and vectors with special reference to the bivariate normal distribution. Biometrika, 56(1):97–108, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334694.

Thiebaux:1967:TMH

[Thi67] Jean Thiebaux. Testing a Markov hypothesis with independence of intermediate states and restricted order. Biometrika, 54(3/4): 605–614, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335051.

Thomas:1969:DFT

[Tho69] E. A. C. Thomas. Distribution free tests for mixed probability distributions. Biometrika, 56(3):475–484, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334653.

Tiao:1966:BCM

[Tia66] George C. Tiao. Bayesian comparison of means of a mixed model with application to regression analysis. Biometrika, 53(1/2):11–25, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334049.

Tiku:1964:AGN

[Tik64] M. L. Tiku. Approximating the general non-normal variance-ratio sampling distributions. Biometrika, 51(1/2):83–95, June 1964. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334198.

Tiku:1965:CSA

[Tik65a] M. L. Tiku. Chi-square approximations for the distributions of goodness-of-fit statistics U_{n}^{2} and W_{n}^{2}. Biometrika, 52(3/4):630–633, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333714.

Tiku:1965:LSF

[Tik65b] M. L. Tiku. Laguerre series forms of non-central χ^{2} and F distributions. Biometrika, 52(3/4):415–427, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333694.

Tiku:1966:NAN

[Tik66] M. L. Tiku. A note on approximating to the non-central F distribution. Biometrika, 53(3/4):606–610, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333670.

Tiku:1967:EMS

[Tik67] M. L. Tiku. Estimating the mean and standard deviation from a censored normal sample. Biometrika, 54(1/2):155–165, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333859.

Tiao:1967:NTC

[TL67] George C. Tiao and Robert H. Lochner. A note on tables for the comparison of the spread of two normal distributions. Biometrika, 54(3/4):683–684, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335067.

TenHoopen:1966:SRS

[TR66] M. Ten Hoopen and H. A. Reuver. The superposition of random sequences of events. Biometrika, 53(3/4):383–389, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333645.

Trawinski:1969:AAE

[Tra69] B. J. Trawinski. Asymptotic approximation to the expected size of a selected subset. Biometrika, 56(1):207–213, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334706.

Tiao:1965:BAR

[TT65] George C. Tiao and W. Y. Tan. Bayesian analysis of randomeffect models in the analysis of variance. I. Posterior distribution of variance-components. Biometrika, 52(1/2):37–53, June 1965. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333810.

Tiao:1966:BAR

[TT66] George C. Tiao and W. Y. Tan. Bayesian analysis of random-effect models in the analysis of variance: II. Effect of autocorrelated errors. Biometrika, 53(3/4):477–495, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333653.

Turin:1960:CFH

[Tur60] G. L. Turin. The characteristic function of Hermitian quadratic forms in complex normal variables. Biometrika, 47(1/2):199–201, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332977.

Thompson:1963:ERS

[TW63] W. A. Thompson, Jr. and T. A. Willke. On an extreme rank sum test for outliers. Biometrika, 50(3/4):375–383, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333906.

Tiao:1964:BTU

[TZ64] George C. Tiao and Arnold Zellner. Bayes’s theorem and the use of prior knowledge in regression analysis. Biometrika, 51(1/2):219–230, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334208.

Ury:1966:LSS

[Ury66] Hans K. Ury. Large-sample sign tests for trend in dispersion. Biometrika, 53(1/2):289–291, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334088.

Verhagen:1961:ERE

[Ver61] A. M. W. Verhagen. The estimation of regression and error-scale parameters, when the joint distribution of the errors is of any continuous form and known apart from a scale parameter. Biometrika, 48(1/2):125–132, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333137.

Verdooren:1963:ETC

[Ver63] L. R. Verdooren. Extended tables of critical values for Wilcoxon’s test statistic. Biometrika, 50(1/2):177–186, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333758. See corrigenda [Ver64].

Verdooren:1964:CET

[Ver64] L. R. Verdooren. Corrigendum: “Extended Tables of Critical Values for Wilcoxon’s Test Statistics”. Biometrika, 51(3/4):527, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334168. See

[Ver63].

vanHeerden:1966:OPP

[vHG66] D. F. I. van Heerden and H. T. Gonin. The orthogonal polynomials of power series probability distributions and their uses. Biometrika, 53(1/2):121–128, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334058.

Vidwans:1964:NNB

[Vid64] Sudhakar M. Vidwans. A note on the negative binomial distribution. Biometrika, 51(1/2):264–265, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334218.

Vagholkar:1960:MEB

[VW60] M. K. Vagholkar and G. B. Wetherill. The most economical binomial sequential probability ratio test. Biometrika, 47(1/2):103–109, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332963.

Walker:1960:SCS

[Wal60] A. M. Walker. Some consequences of superimposed error in time series analysis. Biometrika, 47(1/2):33–43, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332955.

Walker:1961:CDF

[Wal61a] A. M. Walker. Corrigenda: “On Durbin’s Formula for the Limiting Generalized Variance of a Sample of Consecutive Observations from a Moving-Average Process”. Biometrika, 48(3/4):476, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332784. See

[Wal61c].

Walker:1961:LSE

[Wal61b] A. M. Walker. Large-sample estimation of parameters for movingaverage models. Biometrika, 48(3/4):343–357, December 1961. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332756.

Walker:1961:DFL

[Wal61c] A. M. Walker. On Durbin’s formula for the limiting generalized variance of a sample of consecutive observations from a movingaverage process. Biometrika, 48(1/2):197–199, June 1961. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333145. See corrigenda

[Wal61a].

Walker:1962:LSE

[Wal62] A. M. Walker. Large-sample estimation of parameters for autoregressive processes with moving-average residuals. Biometrika, 49(1/2):117–131, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333472.

Wallace:1963:BRB

[Wal63] Margaret E. Wallace. Book review: Introduction to the Mathematical Theory of Genetic Linkage, by N. T. J. Bailey. Biometrika, 50(3/4):547–549, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333934.

Walley:1964:BRB

[Wal64] D. Walley. Book review: Basic Principles of the Tracer Method, by C. W. Sheppard. Biometrika, 51(1/2):294, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334237.

Walker:1967:STS

[Wal67] A. M. Walker. Some tests of separate families of hypotheses in time series analysis. Biometrika, 54(1/2):39–68, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333851. See correction [Wal70].

Walker:1970:CST

[Wal70] A. M. Walker. Corrections: “Some Tests of Separate Families of Hypotheses in Time Series Analysis”. Biometrika, 57(1):226, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334966. See

[Wal67].

Wang:1967:CSV

[Wan67] Ying Y. Wang. A comparison of several variance component estimators. Biometrika, 54(1/2):301–305, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333873.

Watson:1960:MST

[Wat60] G. S. Watson. More significance tests on the sphere. Biometrika, 47(1/2):87–91, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332961.

Watson:1961:BRB

[Wat61a] G. L. Watson. Book review: Finite-Difference Methods for Partial Differential Equations, by George E. Forsythe; Wolfgang R. Watson. Biometrika, 48(3/4):484, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332798.

Watson:1961:GFT

[Wat61b] G. S. Watson. Goodness-of-fit tests on a circle. Biometrika, 48(1/2):109–114, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333135.

Watson:1962:GFT

[Wat62] G. S. Watson. Goodness-of-fit tests on a circle. II. Biometrika, 49(1/2):57–63, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333467.

Watson:1965:EDS

[Wat65] G. S. Watson. Equatorial distributions on a sphere. Biometrika, 52(1/2):193–201, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333824.

Watson:1966:BRB

[Wat66] G. S. Watson. Book review: Statistical Methods for the Analysis of Problems in Animal Orientation and Certain Biological Rhythms, by Edward Batschelet. Biometrika, 53(1/2):295, June 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334091.

Watson:1967:ATU

[Wat67] Geoffrey S. Watson. Another test for the uniformity of a circular distribution. Biometrika, 54(3/4):675–677, December 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335064.

Waugh:1968:ADB

[Wau68] W. A. O’N. Waugh. Age-dependent branching processes under a condition of ultimate extinction. Biometrika, 55(2):291–296, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334868.

Wetherill:1966:SEQ

[WCV66] G. B. Wetherill, H. Chen, and R. B. Vasudeva. Sequential estimation of quantal response curves: A new method of estimation. Biometrika, 53(3/4):439–454, December 1966. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333650.

Walker:1967:EPE

[WD67] Strother H. Walker and David B. Duncan. Estimation of the probability of an event as a function of several independent variables. Biometrika, 54(1/2):167–179, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333860.

Welch:1962:BRB

[Wel62] B. L. Welch. Book review: Elementary Statistical Exercises, by F. N. David; E. S. Pearson. Biometrika, 49(1/2):291, June 1962.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333510.

Wetherill:1960:BRB

[Wet60] G. B. Wetherill. Book review: Quality Control and Industrial Statistics, by Acheson J. Duncan. Biometrika, 47(1/2):212, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332987.

Wetherill:1961:BSA

[Wet61] G. B. Wetherill. Bayesian sequential analysis. Biometrika, 48 (3/4):281–292, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332750.

Wetherill:1962:BRB

[Wet62] G. B. Wetherill. Book review: Statistical Theory and Methodology in Science and Engineering, by K. A. Brownlee. Biometrika, 49(1/2): 290, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333509.

Wilk:1968:PPM

[WG68] M. B. Wilk and R. Gnanadesikan. Probability plotting methods for the analysis of data. Biometrika, 55(1):1–17, March 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334448.

Wilk:1962:EPG

[WGH62] M. B. Wilk, R. Gnanadesikan, and Marilyn J. Huyett. Estimation of parameters of the gamma distribution using order statistics. Biometrika, 49(3/4):525–545, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333985. See corrigenda [WGH63a].

Wilk:1963:CEP

[WGH63a] M. B. Wilk, R. Gnanadesikan, and Marilyn J. Huyett. Corrigenda: “Estimation of Parameters of the Gamma Distribution Using Order Statistics”. Biometrika, 50(3/4):546, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333931. See [WGH62].

Wilk:1963:SML

[WGH63b] M. B. Wilk, R. Gnanadesikan, and Marilyn J. Huyett. Separate maximum likelihood estimation of scale or shape parameters of the gamma distribution using order statistics. Biometrika, 50(1/2):217– 221, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333766.

White:1961:AEM

[Whi61] John S. White. Asymptotic expansions for the mean and variance of the serial correlation coefficient. Biometrika, 48(1/2):85–94, June 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333133.

Whittle:1962:TCP

[Whi62] P. Whittle. Topographic correlation, power-law covariance functions, and diffusion. Biometrika, 49(3/4):305–314, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333965.

Whittle:1963:FMA

[Whi63] P. Whittle. On the fitting of multivariate autoregressions, and the approximate canonical factorization of a spectral density matrix. Biometrika, 50(1/2):129–134, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333753.

Whittle:1964:BPW

[Whi64a] P. Whittle. A branching process in which individuals have variable lifetimes. Biometrika, 51(1/2):262–264, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334217.

Whittle:1964:SGR

[Whi64b] P. Whittle. Some general results in sequential analysis. Biometrika, 51(1/2):123–141, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334201.

Wilkins:1960:TQP

[Wil60] Coleridge A. Wilkins. On two queues in parallel. Biometrika, 47(1/2):198–199, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332976.

Williams:1962:CIV

[Wil62a] J. S. Williams. A confidence interval for variance components. Biometrika, 49(1/2):278–281, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333498.

Williams:1962:SSP

[Wil62b] J. S. Williams. Some statistical properties of a genetic selection index. Biometrika, 49(3/4):325–337, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333967.

Williams:1963:ISE

[Wil63] J. S. Williams. Index selection and estimation from a single sample. Biometrika, 50(1/2):195–199, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333760.

Willis:1964:SPN

[Wil64] D. M. Willis. The statistics of a particular non-homogeneous Poisson process. Biometrika, 51(3/4):399–404, December 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334146.

Williams:1965:DRT

[Wil65a] Trevor Williams. The distribution of response times in a birth-death process. Biometrika, 52(3/4):581–585, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333707.

Williams:1965:SSE

[Wil65b] Trevor Williams. The simple stochastic epidemic curve for large populations of susceptibles. Biometrika, 52(3/4):571–579, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333706.

Williams:1969:DIM

[Wil69] Trevor Williams. The distribution of inanimate marks over a non-homogeneous birth-death process. Biometrika, 56(1):225–228, March 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334712.

Wise:1960:NIB

[Wis60] M. E. Wise. On normalizing the incomplete beta-function for fitting to dose-response curves. Biometrika, 47(1/2):173–175, June 1960. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332968.

Wise:1962:RBM

[Wis62] J. Wise. The relationship between the mean and variance of a stationary birth-death process, and its economic application. Biometrika, 49(1/2):253–255, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333489.

Wise:1963:CMP

[Wis63a] M. E. Wise. Corrigenda: “Multinomial Probabilities and the χ^{2} and X^{2} Distributions”. Biometrika, 50(3/4):546, December 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333932. See [Wis63b].

Wise:1963:MPD

[Wis63b] M. E. Wise. Multinomial probabilities and the χ^{2} and X^{2} distributions. Biometrika, 50(1/2):145–154, June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333755. See corrigenda [Wis63a].

Wise:1964:CMD

[Wis64] M. E. Wise. A complete multinomial distribution compared with the X^{2} approximation and an improvement to it. Biometrika, 51(1/2):277–281, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334224.

Wisniewski:1968:THB

[Wis68] T. K. M. Wisniewski. Testing for homogeneity of a binomial series. Biometrika, 55(2):426–428, July 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2334888. See corrections [Tar90, Wis90].

Wisniewski:1990:ACT

[Wis90] T. K. M. Wisniewski. Amendments and corrections: Testing for Homogeneity of a Binomial Series. Biometrika, 77(3):668, September 1990. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337012. See

[Wis68].

Watson:1964:HA

[WL64] G. S. Watson and M. R. Leadbetter. Hazard analysis. I. Biometrika, 51(1/2):175–184, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334205.

Whittle:1967:CSW

[WL67] P. Whittle and R. O. D. Lane. A class of situations in which a sequential estimation procedure is non-sequential. Biometrika, 54(1/2):229–234, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333866.

Wheeler:1964:DFT

[WW64] Stanley Wheeler and G. S. Watson. A distribution-free two-sample test on a circle. Biometrika, 51(1/2):256–257, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334214.

Yao:1965:ADF

[Yao65] Ying Yao. An approximate degrees of freedom solution to the multivariate Behrens–Fisher problem. Biometrika, 52(1/2):139–147, June 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333819.

Yates:1962:BRB

[Yat62] F. Yates. Book review: Abstracts of Statistical Computer Routines by Fred C. Leone; J. D. Alanen; Gary Andrew; A. S. Qureishi. Biometrika, 49(3/4):584–585, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334010.

Yeo:1964:TDT

[Yeo64] G. F. Yeo. Traffic delays on a two-lane road. Biometrika, 51(1/2): 11–15, June 1964. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334190.

Young:1961:QFU

[You61] D. H. Young. Quota fulfilment using unrestricted random sampling. Biometrika, 48(3/4):333–342, December 1961. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332755.

Young:1962:TAS

[You62] D. H. Young. Two alternatives to the standard χ^{2}-test of the hypothesis of equal cell frequencies. Biometrika, 49(1/2):107–116, June 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333471.

Young:1967:NFT

[You67a] D. H. Young. A note on the first two moments of the mean deviation of the symmetrical multinomial distribution. Biometrika, 54(1/2):312–314, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333877.

Young:1967:RRB

[You67b] D. H. Young. Recurrence relations between the P.D.F.’s of order statistics of dependent variables, and some applications. Biometrika, 54(1/2):283–292, June 1967. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333871.

Young:1968:NSO

[You68] D. H. Young. A note on a sequential occupancy problem. Biometrika, 55(3):591–593, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334272.

Zelen:1969:TSC

[ZF69] M. Zelen and M. Feinleib. On the theory of screening for chronic diseases. Biometrika, 56(3):601–614, December 1969. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334668.

Zinger:1961:DBO

[Zin61] A. Zinger. Detection of best and outlying normal populations with known variances. Biometrika, 48(3/4):457–461, December 1961. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332770.

Zinger:1962:NOA

[Zin62] A. Zinger. A note on optimum allocation for a one-way layout. Biometrika, 49(3/4):563–564, December 1962. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333991.